### v1

# Diffusion Model in Hyperspectral Image Processing and Analysis: A Review

Abstract
Hyperspectral image processing and analysis has important application value in remote sensing, agriculture and environmental monitoring, but its high dimensionality, data redundancy and noise interference etc. bring great challenges to the analysis. Traditional models have limitations in dealing with these complex data, and it is difficult to meet the increasing demand for analysis. In recent years, Diffusion Model, as an emerging generative model, has shown unique advantages in hyperspectral image processing. By simulating the diffusion process of data in time, the Diffusion Model can effectively process high-dimensional data, generate high-quality samples, and perform well in denoising and data enhancement. In this paper, we review the recent research advances in diffusion modeling for hyperspectral image processing and analysis, and discuss its applications in tasks such as high-dimensional data processing, noise removal, classification, and anomaly detection. The performance of diffusion-based models on image processing is compared and the challenges are summarized. It is shown that the diffusion model can significantly improve the accuracy and efficiency of hyperspectral image analysis, providing a new direction for future research.
Index Terms:
Hyperspectral image, diffusion model, High dimensional data processing, Noise removal, Data enhancement, Anomaly detection.
1School of Optical-Electrical and Computer Engineering, University of Shanghai for Science and Technology, No. 516, Jungong Road, Shanghai 200093, China (e-mail: huxing@usst.edu.cn, 242250440@st.usst.edu.cn, dwzhang@usst.edu.cn)
2School of Electronics and Electrical Engineering, Shanghai University of Engineering Science, Shanghai 201620, China (e-mail: dqq1019@163.com)
3Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing 100094, China (e-mail: hongdf@aircas.ac.cn)
4School of Electronic, Electrical and Communication Engineering, University of Chinese Academy of Sciences, Beijing 100049, China.
I Introduction
Hyperspectral image processing and analysis have important application values in remote sensing, agriculture, environmental monitoring and other fields. With the advancement of technology, the resolution and spectral range of hyperspectral images are continuously improved, which provides the possibility of fine feature identification and classification. However, the high dimensionality, data redundancy, and noise interference etc. of hyperspectral images also bring great challenges to the analysis. Traditional processing methods have certain limitations in dealing with these complex data, so the introduction of new models and methods has become the focus of current research. In recent years, Diffusion Model (DM), as an emerging generative model, has shown unique advantages in hyperspectral image processing and analysis.
Hyperspectral imagery provides a wealth of spectral and spatial data by capturing spectral information of objects in multiple narrow bands. The high-dimensionality and complexity of this data make it widely used in agriculture, environmental monitoring, mineral exploration and other fields. However, hyperspectral image analysis also faces many challenges. First, there is a huge demand for high-dimensional data processing and storage with high computational complexity. Second, the high correlation between bands leads to data redundancy, which increases the difficulty of analysis. In addition, noise interference and mixed image element problems also affect the accuracy of the analysis.
Modeling using neural networks has been a dominant force over the last decade or so, and has had a huge impact in a variety of domains since the emergence of generative models, ranging from images [1, 2] and audio [3, 4] to text [5] and point clouds [6]. From a probabilistic modeling perspective, the key defining feature of a generative model is that it is trained in such a way so that its samples $\widetilde{{x}}\sim p_{\theta}(\widetilde{{x}})$ are from the same distribution as the training data distribution ${x}\sim p_{\theta}({x})$. Energy-based models (EBM) achieve this by defining non-normalized probability densities on the state space; however, these methods require Markov chain Monte Carlo (MCMC) sampling during both training and inference, which is a slow and iterative process [7] In the last few years, there has been a renewed interest in generative modeling due to the advancement of deep learning architectures in general, which have revealed improved visual fidelity and sampling speed. Specifically, Generative Adversarial Networks (GANs) [8], Variational Auto-Encoders (VAEs) [9] and Normalized Streams [10] among others, have emerged. In addition to this, generative models based on diffusion processes provide an alternative to existing VAEs, EBMs, GANs, and normalized streams, which do not require aligning a posteriori distributions, estimating tricky partition functions, introducing additional discriminator networks, or placing network constraints, respectively. To date, diffusion models have been found useful in a variety of domains, ranging from generative modeling tasks (e.g., image generation [11], image super-resolution [12] image restoration [13]) to discriminative tasks (e.g., image segmentation [14], classification [15], and anomaly detection [16]). Recently, the hyperspectral community has witnessed an exponential growth in the number of diffusion-based techniques, and a large number of studies have been dedicated to the application of diffusion models in various hyperspectral scenarios. Since diffusion modeling has recently received a great deal of attention from the research community, the literature has a large number of contributions in this direction. For this reason, this survey aims to provide a comprehensive overview of the latest advances and a comprehensive overview of such models in the field of hyperspectral spectroscopy.
Different from this paper [17], firstly, it is aimed at generalized remote sensing images, covering multimodal data such as multispectral, SAR, and panchromatic images, with a wider scope of research but lack of in-depth analysis of hyperspectral properties, and focuses on sorting out the technological paths of traditional tasks such as image generation and super-resolution reconstruction, with the contribution of systematically establishing a methodological mapping of diffusion modeling in remote sensing for the first time. Secondly, it focuses more on the panoramic depiction of diffusion modeling applications in remote sensing, and although the conditional diffusion model [11] as well as the latent diffusion model [18] are described, the exploration of the emerging architectures stops at early applications before 2023, lacks a systematic analysis of practical engineering constraints, and does not provide an in-depth discussion of the key technological challenges specific to hyperspectral imagery. Instead, this paper focuses on the specific modality of hyperspectral imagery, and designs a multi-perspective classification of diffusion models in hyperspectral domain from the core characteristics of high dimensionality of data and spectral redundancy etc., which provides a systematic taxonomy for the study of diffusion models and their applications. Meanwhile, this paper classifies the existing diffusion models into three categories: denoising diffusion probabilistic models (DDPM) [19] and score-based generative models (SGM) [20] and stochastic differential equation-based models (SDE) [21], and analyzes the use of diffusion models in hyperspectral domain with specific classification. In addition, we will systematically sort out the progress of the application of diffusion models in vertical tasks, which is more comprehensive than the direction of [17], and categorize the tasks into eight categories: data enhancement and synthesis, data generation and complementation, classification, segmentation, image generation, anomaly detection and noise suppression, and data recovery. The significance of this paper lies in the following: first, it fills the gap of systematic review of diffusion model in hyperspectral field; second, it elaborates the methodological innovation of adaptation in hyperspectral field more comprehensively; third, it reveals the unique challenges and optimization paths of diffusion model in hyperspectral scenarios; fourth, it realizes a leap forward breakthrough in the three dimensions of theoretical depth, technological breadth, and application value, and lays down a key theoretical and practical foundation for hyperspectral image processing to move from laboratory research it lays a key theoretical foundation for hyperspectral image processing to move from laboratory research to operational applications.
The overall structure of this paper is organized as follows: Section I will introduce the definition, characteristics, data properties, and main tasks in the analysis of hyperspectral images; Section II will introduce the theory, evolution, and variants of the Diffusion model, the analysis in comparison with the traditional model, and the application in image generation; Section III will introduce the use of Diffusion model in hyperspectral image analysis and its specific applications; Section IV will introduce the challenges and limitations of Diffusion model; Section V will introduce the comparison between Diffusion model and traditional methods, and conduct experiments and analysis; Section VI will provide a future outlook of Diffusion model in hyperspectral image; Section VII will summarize the content of this paper.
II Basics of hyperspectral imagery
II-A Definition and Characteristics of Hyperspectral Imagery
Hyperspectral image is a spectral image with spectral rate in the range of ${10}^{-2}\lambda$ order of magnitude, covering the range from visible to infrared light range of multiple continuous bands, through the hyperspectral sensors (imaging spectrometer) mounted on different platforms, it can be imaged in the ultraviolet, visible, near-infrared and mid-infrared regions of the electromagnetic spectrum by tens to hundreds of consecutively subdivided spectral bands in the target region, and each pixel point includes information of numerous spectral bands, forming a continuous spectral curve, which can obtain spatial information as well as spectral information at the same time, truly combining spectra and images.
Hyperspectral images benefit from the development and maturity of imaging spectroscopy technology, which is a sample that combines image information and spectral information, implying a wealth of features that are useful for feature identification and processing.
Image information is containing external quality information such as size, shape, defects, etc., and spectral information is containing different internal physical structure and chemical composition. Both the remote sensing image with multiple segments and the continuous spectral information of each image element can be obtained, realizing the multi-dimensional expression of feature characteristics, and the combination of spatial and spectral information, which is convenient for accurate identification and classification.
Hyperspectral images have a high spectral resolution, usually at the nanometer level, and are able to finely portray the reflectance spectra of features, differentiate the spectral characteristics of similar substances, and improve the ability to classify and identify features. At the same time, it has a large amount of information, redundancy of information, multi-segmented, data including dozens of hundreds of segments, huge amount of data, high requirements for storage and processing capacity, providing rich spectral information, able to capture subtle spectral differences, and correlate between them spatially and spectrally.
Finally, hyperspectral images are also characterized by fine spectral feature expression, nonlinear characteristics, low signal-to-noise ratio and high dimensionality of data. Moreover, hyperspectral images are widely used in many fields, such as agriculture, environmental monitoring, mineral exploration, medicine and so on, and can recognize and classify different substances.
Figure 1: Hyperspectral image (a) and Illustration of hyperspectral image (b)
II-B High-dimensional characterization of hyperspectral image data
The data of hyperspectral images are characterized by high dimensionality, which is mainly reflected in the multidimensional structure, data redundancy, computational complexity, and information richness.
Hyperspectral image is a three-dimensional data cube, containing spatial, radiometric and spectral dimensions, with a large number of bands, consecutive bands forming high-dimensional vectors, and covering a range of bands with narrow band widths and high resolution at the nanometer level, which finely reflects the physicochemical properties, while adjacent bands are highly similar, resulting in redundancy of information and increasing the complexity of the data. Hyperspectral image is a combination of spatial and spectral, forming a “map in one”, which makes the dimension of the information itself is higher, when analyzing the data, we have to consider the spatial up and down and spectral characteristics, increasing the complexity of the model. High-dimensional data requires a large amount of computational resources when processing, especially when performing tasks such as feature extraction, classification, segmentation, etc., the computational complexity increases significantly. The storage requirements are high due to the huge amount of data, which is in high demand and costly. Traditional machine learning methods are slow to train in high-dimensional space and are susceptible to noise interference, which is especially obvious in high dimensions and increases the difficulty of processing; although it can be automatically extracted under deep learning, a large number of markers are required and the cost is extremely high.
The high dimensional nature of the data creates excessive waste of resources for the processing, which can easily lead to dimensional disaster and inverted performance degradation. In this regard there is a great need for the use of dimensionality reduction methods, such as PCA, LDA, etc. to reduce the dimensionality, retain important information and reduce the complexity of the model. Under the complexity process, it is important to ensure that problems such as overfitting in training occur.
II-C Main tasks in hyperspectral image analysis
The main tasks of hyperspectral image analysis include classification, segmentation, and anomaly detection etc. These tasks have important application value in hyperspectral image processing and can help us to extract useful information from complex high-dimensional data. The following is a detailed description of these tasks:
Classification: Hyperspectral image classification refers to the assignment of each pixel or region in an image to a specific class. These categories can be different types of features, such as vegetation, water bodies, buildings, and so on. It is often applied to crop classification and pest detection in agriculture; land use classification and forest cover monitoring in environmental monitoring; and mineral type identification in mineral exploration. Often exists high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of classification; data labeling: the classification task usually requires a large amount of labeled data, and the labeling cost of hyperspectral images is high; mixed image element problem: due to the limitation of spatial resolution, one pixel may contain multiple features, which increases the difficulty of classification and other problems. The traditional methods are statistical-based methods (e.g., maximum likelihood classification), machine learning-based methods (e.g., support vector machine SVM). Deep learning methods are convolutional neural networks (CNN), recurrent neural networks (RNN), etc.
Segmentation: Hyperspectral image segmentation refers to dividing an image into multiple regions or objects, each with similar spectral characteristics. Segmentation can be pixel-level segmentation or object-level segmentation. It is often applied to the segmentation of tumor regions in medical image analysis, the segmentation of buildings and roads in urban planning, and the segmentation of crop regions in agriculture. Often exist high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of segmentation; noise interference: noise in hyperspectral images affects the accuracy of segmentation; mixed image element problem: a pixel may contain multiple features, which increases the difficulty of segmentation and other problems. The traditional methods are clustering based methods (e.g. Kmeans), edge detection based methods. Deep learning methods are Full Convolutional Networks (FCN), UNet, etc.
Target Detection: Hyperspectral image target detection refers to locating and identifying specific targets or objects in an image. These targets can be military targets, mineral resources, vegetation types and so on. Often used in military target identification and tracking; mineral exploration in the detection of specific minerals; environmental monitoring in the detection of specific vegetation or pollution sources. Often exist high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of target detection; target diversity: the spectral characteristics of different targets may be similar, increasing the difficulty of detection; background interference: complex background will affect the accuracy of target detection and other problems. Traditional methods are based on spectral matching methods, statistical methods. Deep learning methods are Regional Convolutional Neural Networks (RCNN), YOLO (You Only Look Once) and so on.
Anomaly Detection: Hyperspectral image anomaly detection refers to the detection of pixels or regions in an image that are significantly different from their surroundings. These anomalies may be unknown targets or features. It is often applied to the detection of unknown targets in the military, the detection of pollution sources in environmental monitoring, and the early detection of pests and diseases in agriculture. Often exist high-dimensional data processing: the high-dimensional nature of hyperspectral images increases the complexity of anomaly detection; diversity of anomalies: different types of anomalies may have different spectral characteristics, increasing the difficulty of detection; background interference: complex background can affect the accuracy of anomaly detection and other problems. The traditional methods are statistical based methods (e.g. RX algorithm), clustering based methods. Deep learning methods are Autoencoder, Generative Adversarial Networks (GAN), and so on.
Noise suppression: Hyperspectral noise suppression refers to the reduction or elimination of noise in hyperspectral images by means of specific algorithms and techniques, in order to improve the quality of the image and the accuracy of the subsequent analysis. It is mainly used in agriculture, military, food industry, mining exploration and other fields. Commonly used methods are mainly frequency or transform domain frequency reduction, spatial frequency reduction, and deep learning frequency reduction, and so on.
Data recovery: Hyperspectral data recovery refers to the removal of noise and interference in hyperspectral images through specific algorithms and technical means to restore their original spectral and spatial information to improve the quality and usability of data, which belongs to the further implementation of noise suppression. The main application areas are agriculture, environmental monitoring, biomedicine, geological exploration and so on. The commonly used methods are spectral reconstruction, spectral dimension recovery, low-rank representation modeling, and deep learning methods, and so on.
Data Augmentation and Synthesis: Hyperspectral image data augmentation and synthesis refers to increasing the diversity and quantity of training data by generating new hyperspectral image samples, so as to improve the generalization ability of the model. It is often applied in model training to increase the diversity of training data and improve the generalization ability of the model; and in data completion to generate missing hyperspectral image data. Often exist high-dimensional data processing: the high-dimensional nature of hyperspectral images increases the complexity of data enhancement and synthesis; data quality: the generated samples need to be of high quality and diversity to ensure the validity of the model and other issues. The traditional methods are interpolation-based methods, statistics-based methods. Deep learning methods are Generative Adversarial Networks (GAN), Diffusion Model, and so on.
Data generation and completion: Hyperspectral data generation and completion is a core task in data science, aiming to solve the problems of missing data, noise interference and data scarcity, and at the same time extend the application scenarios through synthetic data. Data generation generates synthetic data that conforms to the real data distribution through algorithms, which is used to enhance the data scale or simulate specific scenarios; data completion fills in the missing values or repairs incomplete data to maintain the consistency and completeness of the data distribution. It is usually used in remote sensing, medical imaging, agricultural monitoring and other fields. The traditional methods are based on statistics (interpolation), machine learning (K-nearest neighbor, decision tree) deep learning methods such as GAN, VAE, Diffusion model, etc.
In summary, the main tasks of hyperspectral image analysis include classification, segmentation, target detection, anomaly detection and data enhancement and synthesis. These tasks have important applications in hyperspectral image processing and can help us extract useful information from complex high-dimensional data. Although these tasks face challenges such as high-dimensional data processing, noise interference, and mixed image element problems, the accuracy and efficiency of hyperspectral image analysis can be significantly improved by combining traditional and deep learning methods.
III Diffusion Model
Diffusion Model is an emerging generative model that has demonstrated powerful capabilities in the fields of image generation, denoising, and data enhancement in recent years. It gradually transforms noisy data into target data by modeling the diffusion process of data in time. The core idea of the diffusion model is to gradually transform the data into noise through the forward diffusion process, and then recover the original data from the noise through the reverse diffusion process. This approach performs well in high-dimensional data processing, image generation and denoising tasks, and especially has unique advantages in hyperspectral image processing.
III-A Theoretical Foundations of the Diffusion Model
Diffusion model is divided into two processes: forward and backward diffusion diffusion; and finally training objectives are performed to minimize the difference between forward and backward diffusion processes. The basic diffusion models are usually divided into (1) denoising diffusion probabilistic model (DDPM), (2) score-based generative model (SGM), and stochastic differential equation-based model (SDE/NCSN). The basic principles of these three diffusion models are described below.
III-A1 Denoising Diffusion Probabilistic Model (DDPM)
Denoising Diffusion Probabilistic Model (DDPM) [19, 22] utilize two Markov chains: a forward chain that perturbs the data into noise, and an inverse chain that converts the noise back into data. The former is usually manually designed to convert any data distribution into a simple prior distribution (e.g., a standard Gaussian distribution), while the latter Markov chain inverts the former by learning a transition kernel parameterized by a deep neural network.
New data points are generated by first randomly sampling a vector from the previous distribution and ancestrally sampling it via an inverse Markov chain [23].
Figure 2: Diffusion processes in diffusion models
Formally, given a data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$,, a forward Markov process generates a sequence of random variables $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ replaced by a transition kernel $q(x_{t}\mid{x_{t-1}})$. Using the probability chain rule and the Markov property, the joint distribution of $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ is later decomposed, conditional on $\mathbf{x}_{0}$, and denoted as $q(x_{1},\ldots{,x_{T}}\mid{x_{0}})$, which becomes:
$q(\mathbf{x}_{1},...,\mathbf{x}_{T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$ (1)
In DDPM, the transition kernel $q(x_{t}\mid{x_{t-1}})$ is drawn manually to incrementally transform the data distribution $q(\mathbf{x}_{0})$ into a processable prior distribution. A typical design of a transition kernel is a Gaussian perturbation, and the most common choice of transition kernel is:
$q(\mathrm{x}_{t}\mid\mathrm{x}_{t-1})=N(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\mathrm{I}).$ (2)
where $\beta_{t}\in(0,1)$ is the hyperparameter chosen prior to model training. This kernel is applied to simplify the discussion, although other types of kernels apply in the same vein. As observed by Sohl-Dickstein et al. 2015 [22] this Gaussian transition kernel allows us to marginalize the joint distribution in equation (1) to obtain an analytic form of $q(x_{t}\mid\mathbf{x}_{0})$ for all $t\in{0,1,\cdots,T}$. $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}{:}=\prod_{s=0}^{t}\alpha_{s}$, in particular:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right).$ (3)
Intuitively, this forward process slowly injects noise into the data until all structure is lost. To generate new data samples, the DDPM first generates an unstructured noise vector from the prior distribution (which is usually easy to obtain), and then gradually removes the noise from it by running a learnable Markov chain in the opposite time direction. Specifically, the inverse Markov chain is parameterized by the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ and the learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. We choose the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ because the forward process is constructed such that $q(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$. The learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ employs:
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\right).$ (4)
where $\theta$ denotes the model parameters and the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized by the deep neural network.
The key to the success of the network training goal is to train the reverse process $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$ to match the actual time reversal of the forward process $q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$. This is usually achieved by minimizing the KL scatter (Kullback-Leibler) between the two:
$\mathrm{KL}\left(q(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})|\||p_{\theta}(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})\right)$ (5)
=$\overset{(i)}{\operatorname*{=}}-\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})}[\log p_{\theta}\left(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T}\right)]+\mathrm{const}$ (6)
=$\begin{split}\overset{(ii)}{\operatorname*{=}}\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})}\left[-\mathrm{log}p(\mathbf{x}_{T})-\sum_{t=1}^{T}\log\frac{p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})}{q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})}\right]\\ +\mathrm{const}:=-L_{NLB}(\mathbf{x}_{0}).\end{split}$ (7)
$\overset{(iii)}{\operatorname*{\operatorname*{\geq}}}\mathbb{E}[-\log p_{\theta}\left(\mathbf{x}_{0}\right)]+\mathrm{const}$ (8)
where “const” is a constant that does not depend on the model parameters $\theta$. The goal of DDPM training is to maximize the variational lower bound or equivalently minimize the negative variational lower bound, and thus it can be efficiently estimated by Monte Carlo sampling [24] and efficiently optimized by stochastic optimization [25].
III-A2 Score-Based Generative Models (SGM/NCSN)
At the heart of Score-Based Generative Models (SGM/NCSN) [26, 20, 27] is the notion of a (Stein) score (aka fraction or score function) [28]. Given a probability density function p(x), the score function is defined as the gradient of the logarithmic probability density $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Unlike the Fisher score $\nabla_{\theta}\log p_{\theta}(\mathbf{x})$, which is commonly used in statistics, the Stein score considered here is a function of the data x and not of the model parameters $\theta$. It is a vector field pointing in the direction where the probability density function has the largest growth rate.
The key idea of Score-Based Generative Models (SGM) [20] is to perturb the data with a series of augmented Gaussian noises and jointly estimate the score function for all noisy data distributions by training a deep neural network model (called Noise Conditioned Scoring Network, NCSN, in [20]) conditional on the noise level. The samples are generated by linking the score functions with decreasing noise levels by using score-based sampling methods including Ronzmann Monte Carlo [29, 30, 31, 20, 21], stochastic differential equations [32, 21]], ordinary differential equations [33, 34, 35, 36, 21], and various combinations thereof [21]. In the formulation of the score-based generative model, training and sampling are completely decoupled, so that multiple sampling techniques can be used after estimating the score function.
[…]
$\mathbb{E}_{t\sim\mathcal{U}[[1,T]],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[\lambda(t)\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].$ (9)
$\begin{split}=\mathbb{E}_{t\sim\mathcal{U}\|1,T\|\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\sigma_{t}^{2}\|\nabla_{\mathbf{x}_{t}}\log q\left(\mathbf{x}_{t}\right)-\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].\end{split}$ (10)
=$\begin{split}\overset{(i)}{\operatorname*{=}}\mathbb{E}_{t\sim u[[1,T]]\mathbf{x}_{0}\sim q(\mathbf{x}_{0})\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}[\lambda(t)\sigma_{t}^{2}\\ \prod\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})-\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\parallel^{2}]+\mathrm{const}.\end{split}$ (11)
=$\begin{split}\overset{(ii)}{\operatorname*{=}}\mathbb{E}_{t\sim U[1,T],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\left\|-\frac{\mathbf{x}_{t}-\mathbf{x}_{0}}{\sigma_{t}}-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right\|^{2}\right]+\mathrm{const}.\end{split}$ (12)
$\begin{split}(iii)\\ =&\mathbb{E}_{t\sim\mathcal{U}[[1,T]],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I})}[\lambda(t)\|\epsilon+\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\|^{2}]\\ +\mathrm{const}.\end{split}$ (13)
where $\lambda(t)$ is a positive weighting function, $\mathbf{x_{t}}$ is computed from $\varepsilon$ and $\mathbf{x}_{0}$, $\mathcal{U}[[1,T]]$ is a uniform distribution on the set ${1,2,\cdots,T}$, and $\varepsilon$ is a deep neural network with parameter $\theta$ for predicting the given $\mathbf{x}_{t}$ and t noise vectors. Comparing the equations, the training objectives of the DDPM and SGM are equivalent once $\epsilon_{\theta}(\mathbf{x},t)=-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x},t)$ is set.
III-A3 Diffusion model based on stochastic differential equations (Score-SDE)
When DDPM and SGM can be further generalized to the case of infinite time step or noise level, where the perturbation and denoising process is the solution of a stochastic differential equation (SDE), a formulation known as Score SDE [21], which is utilized for the noise perturbation and sample generation, and the denoising process requires estimation of a score function of the distribution of noisy data.
The SDE was used to score the data using a diffusion process controlled by the following stochastic differential equation (SDE) [21]:
$\mathrm{dx}=\mathrm{f}(\mathrm{x},t)\mathrm{d}t+g(t)\mathrm{dw}.$ (14)
where f(x,t) and g(t) are the diffusion and drift functions of the SDE, and w is a standard Wiener process (aka Brownian motion) The forward processes in both the DDPM and the SGM are discretizations of this SDE. 2020 As demonstrated in the literature [21] by Song et al. the corresponding SDE for the DDPM is: $dx=-\frac{1}{2}\beta(t)xdt+\sqrt{\beta(t)}dw$. For the SGM the corresponding SDE is: $\mathrm{dx}=\sqrt{\frac{\mathrm{d}[\sigma(t)^{2}]}{\mathrm{d}t}}\mathrm{dw}$. where, $e\beta(\frac{t}{T})=T\beta_{t}$, $\sigma(\frac{t}{T})=\sigma_{t}$ when T tends to infinity.
Here $q_{t}({x})$ is used to denote the distribution of $\mathbf{x}_{t}$ in the forward process. For any diffusion process of the form of equation (14), Anderson proposed [42] to show that it can be inverted by solving the following reverse time SDE:
$\mathrm{d}\mathbf{x}=[\mathrm{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})]\mathrm{d}t+g(t)\mathrm{d}\overline{\mathbf{w}}.$ (15)
where $\bar{\mathbf{W}}$ is the standard Wiener process when time flows backwards and dt denotes an infinitesimal negative time step. The solution trajectories of this reverse SDE have the same marginal density as those of the forward SDE, except that they evolve along opposite time directions [21]. Intuitively, the solution of a reverse-time SDE is a diffusion process that gradually converts noise into data. In addition, the literature presented by Song et al. in 2020 [21] proved the existence of an ordinary differential equation (ODE), the probabilistic flow ODE, whose trajectory has the same margin as the reverse time SDE. The probabilistic flow ODE is given by the following equation:
$\mathrm{d}\mathbf{x}=\left[\mathrm{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})\right]\mathrm{d}t.$ (16)
Both the inverse time SDE and the probabilistic flow ODE allow sampling from the same data distribution because their trajectories have the same margins. If one knows the fractional function $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ for each time step t, one can derive its inverse-time equations, which can then be used to generate samples by numerical techniques such as annealed Ronzwan dynamics [20], numerical SDE solvers [32, 21], numerical ODE solvers [33, 34, 21, 36, 43]], and predictor-corrector methods (a combination of MCMC and numerical ODE/SDE solver combination) [21]. As with the SGM, the transient fraction model $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is parameterized and the fraction function is estimated by generalizing the fraction matching objective in Equation (13) to continuous time, resulting in the following objective:
$\begin{split}\mathbb{E}_{t\sim\mathcal{U}[0,T],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{0t}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)\|^{2}\right].\end{split}$ (17)
where $\mathcal{U}[0,T]$ denotes the uniform distribution on $\mathbf{[}0,T]$.
The above is an introduction to the principles and formulas of the basic diffusion model, through their theoretical support, which makes the diffusion model work in the modern image field, solving certain limitations of the traditional model, but also brings defects. The study of Denoising Diffusion Probabilistic Models (DDPM) is the main focus in targeting the hyperspectral field, and it is also the most applied diffusion model in this field, the other two types of applications are relatively less, and they will also be introduced in the following sections.
III-B Evolution of the Diffusion model with different variants
Diffusion model started in 2015, Sohl-Dickstein et al. first proposed the theoretical framework of diffusion model, which combines nonequilibrium thermodynamics with generative model for the first time [22].In 2019, Song Yang et al. proposed Score-Based Generative Models (SGM/NCSN) [20], combining Score-Based Generative Models with Score Matching,which is a score-based generative model that generates by estimating the logarithmic gradient of the data distribution. Immediately following the optimization of the model in 2020, the team proposed NCSN++ (Noise-Conditioned Score Networks++) [27], which improves the network architecture (introducing residual connectivity and adaptive normalization), optimizes the noise scheduling strategy, and proposes a more stable training objective function. In the same year, Ho et al. proposed DDPM (Denoising Diffusion Probabilistic Models) [19], which used fixed noise scheduling and reparameterization techniques and became the basis of modern diffusion models. Later, in order to address the deficiencies of DDPM in terms of generation quality, training efficiency, and log-likelihood optimization, and so on, Improve-DDPM (IDDPM) [44] was proposed by Nichol et al. in 2021, which introduced learnable variance, cosine noise scheduling, and hybrid loss function, and accelerated the sampling strategy by time-step-importance sampling, which made the DDPM effectively improved , but the computation is complicated, the dataset has limitations, the generalization ability is insufficient, and the generation quality and speed belong to this and that.Song et al. proposed DDIM (Denoising Diffusion Implicit Models) [35] in the same year, which accelerates the sampling through non-Markov chains. Meanwhile, Rombach et al. proposed LDM (Latent Diffusion Model) [18], which shifts the diffusion process to latent space instead of operating directly in pixel space, dramatically reducing the computational cost and ensuring the generation quality. The enhancement for NCSN++ was also introduced in the same year with Score-SDE (based on stochastic differential equations) [21] as well as ODE (probabilistic streaming) [21], which optimize the sampling process with stochastic and constant differentiation and optimize the sampling process, respectively. Also in 2021, Dhariwal and Nichol proposed Classifier Guidance (Conditional Diffusion Model/CDM) [11], is to significantly improve the quality of generation by incorporating a classifier to guide the generation.In 2022, Rombach et al. published Stable Diffusion [18], which is based on the LDM model improved to enhance text-to-image generation. There are important variants [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] that have been improved since then, but do not constitute a “landmark” model. Below is a timeline of the evolution of the diffusion model and its main variants:
Figure 3: Evolution timeline of the Diffusion model
III-C Traditional Deep Learning Models
Generative Models (GM) is a class of machine learning models that can learn distributions from data and generate new samples. It includes data generation, data enhancement, image restoration, anomaly detection and so on. It is mainly valuable in computer vision, natural language processing and other fields. In this paper, we will focus on the traditional generative models: the GAN, VAE, autoregressive models.
Generative adversarial network (GAN) is another important network framework for unsupervised anomaly detection techniques.GAN was originally conceptualized in the zero-sum game framework of game theory. It was proposed by Ian Goodfellow et al. in 2014 [8]. The core idea is to generate data by adversarial training of two neural networks: a generator and a discriminator. The generator is responsible for generating data, and the discriminator is responsible for distinguishing the generated data from the real data. Through this adversarial process, the generator gradually learns to generate data similar to the distribution of the real data, which is fast to generate, but the training is unstable and prone to pattern collapse problem [58, 59, 60].
Variational Auto-Encoder (VAE) [9] is a generative model that combines the advantages of Autoencoder and probabilistic graphical models to learn latent representations from data and generate new data samples.The core idea of VAE is to approximate the latent distributions of the data by Variational Inference (VII), and to generate the decoder by generating new data with lower generation quality but stable training [61, 62, 63].
Autoregressive Models (ARMs) [64] are a class of generative models that model sequence data by assuming that the current data point is dependent on previous data points. The core idea of autoregressive models is to utilize the temporal dependence of sequence data to generate new data points step by step, e.g., RNN (Recurrent Neural Network) forms a recursive body in the process of processing the data; LSTM (Long Short-Term Memory Neural Network)) a kind of RNN with a special pattern. Autoregressive models generate the complete data by generating each part of the data step by step, which is slow in generating speed, but generating high quality [65, 66].
III-D Successful Application of Diffusion Modeling in Image Generation
Diffusion model has a wide range of applications, for images on the application began in 2020, mainly based on Denoising Diffusion Probabilistic Models (DDPM), first of all, the initial proposed DDPM [19], IDDM [44], and DDIM [35], etc., through the Markov chain as well as the addition of the noise and denoising process, and finally the prediction of the noise residuals to optimize the quality of the generation. However, the number of sampling steps is numerous and the computation speed is slow. Then ADM-G (Classifier-Guided-Diffusion) [11], introduced by OpenAI team in 2021, made the model surpass GAN in FID metrics for the first time by introducing classifier-guided generation. but dependent on classifiers and high computational cost. Later, in order to optimize text-image alignment and super-resolution generation, Imagen [46] was introduced by the Google Research team, which combines the T5-XXL text encoder with the cascade diffusion architecture, and has outstanding text alignment capability, but with great resource requirements. Later, in order to realize multimodality and efficient generation, Stabel-Diffusion [18] was proposed by Robin Rombach et al. in 2023, which compressed the diffusion process into the latent space, significantly reducing the computational complexity, but losing information in the latent space. Meanwhile, also by the OpenAI team, CLIP-Guded-Diffusion [49] was introduced, which combines the CLIP model to realize cross-modal generation, supports complex semantic control, and generates semantic consistency, but has large training requirements and slow generation speed. Subsequently, in order to unify the framework and extend the multi-tasking domain, OmniGen [53], the first unified framework supporting tasks such as text generation, image editing, and so on, without additional plug-ins, was introduced by the Wisdom Source Institute in 2024. Then, the Byte Jump team introduced D-Dit [55], which combines continuous and discrete diffusion techniques to achieve bidirectional text-image generation, support multimodal tasks, and improve parameter efficiency by about 20%. In 2025 by Yifan Zhou et al. proposed AF-LDM [57], which introduces shift isovariance loss, improves the UNet attention module, reduces potential spatial noise bias, improves video generation consistency, and enhances the model robustness, etc.
At the same time, the diffusion model will continue to be innovated and developed in the hyperspectral field in the future, and these models are improved on the basis of the main diffusion model, aiming at solving the problem of image in classification, anomaly detection and recovery, and so on.
Figure 4: Improved division of the diffusion model according to different variants
TABLE I: Basic Diffusion Models and Variants
Basic Diffusion Model Major Variants Reasons for Improvement Improvement Point Drawbacks Applications
DDPM IDDPM Improve generation quality and training stability. Introduction of cosine noise scheduling and hybrid objective function. to optimize the training process. Generation is still slow. Image generation, audio synthesis.
DDIM Accelerating the generation process. Non-Markov chain inverse process allowing large step size sampling. Sacrificing some generative diversity for speed. Fast image generation, interpolation.
Classifier-Guided Enhanced generation control. Using classifier gradients to guide the direction of generation and improve the quality of conditional generation. Additional training of classifiers is required, increasing computational cost. Conditional image generation, text-to-image.
LDM/Stable Reduced computational costs. Diffusion in latent space to reduce computational overhead in high-resolution images. Latent space compression may lose detail. High-resolution image generation, art creation.
SGM/NSCN NSCN++ Enhancement of model presentation skills. Introducing deeper network structure and improved score matching goals. High training complexity. Complex distribution modeling, scientific simulation.
Score-SDE Harmonization of diffusion frameworks. Modeling Continuous Diffusion Processes by Stochastic Differential Equation (SDE) for More General Theory. The sampling process is computationally intensive. Multimodal generation, cross-domain generation.
ODE Improved generation process stability. Replacing SDE with ordinary differential equations (ODE) for deterministic sampling. No significant increase in generation speed compared to SDE. Controlled generation, inverse problem solving.
IV Applications in Hyperspectral Image Processing and Analysis
Deep learning models are widely used in various fields and show powerful advantages, in recent years, some scholars have also applied them to the field of hyperspectral, while making many achievements. The application in hyperspectral image analysis is embodied in many aspects, mainly including data enhancement and synthesis, generation and complementation of hyperspectral data, image classification, segmentation, target detection, and anomaly detection, and many other aspects, in this section, the hyperspectral image processing is divided into two aspects: one is the image processing under the traditional deep learning model, and the other is the image processing under the diffusion-based model. The following is a description of these methods.
Figure 5: Deep learning model-based hyperspectral image processing
Figure 6: Timeline of the birth of each model
IV-A The main Diffusion model used in the field of hyperspectral imagery
Diffusion Model based Hyperspectral Image Processing, which was first applied in the field of Hyperspectral in 2022, meanwhile, it was obtained through the research that the main use of Diffusion Modeling methods in this field is mainly Denoising Diffusion Probabilistic Models (DDPM), and this paper also presents the technical analysis of DDPM and its variants in the field in this section. The main diffusion models for hyperspectral image processing based on diffusion modeling are DDPM [19], IDDPM [44], DDIM [35], LDM [18], Classifier-Gided [11], Stable-Diffusion [18].
First of all, the basic principles of DDPM have been introduced above, and the basic principles and formulas of IDDPM and DDIM are the same as those of DDPM, so we won’t repeat them here.LDM and Stable-Diffusion belong to the same category, in which the core ideas are the same, and it can be said that Stable-Diffusion is the specific implementation version of LDM, so the following will introduce the following LDM (Latent Diffusion Model), Classifier-Guided (Conditional Diffusion Model) two basic principles, formula.
IV-A1 Conditional Diffusion Model (CDM)
Conditional diffusion models are similar to the development of GANs in that diffusion models begin with unconditional generation, followed immediately by conditional generation [67]. Unconditional generation is often used to explore the upper limits of the model’s capabilities, whereas conditional generation is more beneficial for applications as it allows the output to be controlled according to the human’s wishes.
The first work to introduce this condition in diffusion models was [11], which guided the generation of diffusion models by adding classifiers to trained diffusion models, hence the term guided diffusion models. Although this approach has a low training cost, it increases the inference cost by using the classification results to guide the sampling process of the diffusion model and has poor control over the details to produce images with better results. Therefore, the Google team [68] decided to adopt a simple idea of controlling the generated results by conditionally retraining the DDPM and named it Classifier-Free Guidance.
Formally, given the conditional information $\mathbf{c}$, the distribution of the combined DDPM changes from equations (2)(3), and accordingly, the optimization objective and sampling process are modified to the following form:
$\mathcal{L}_{\mathrm{con}}(\theta)=\mathbb{E}_{x_{0},\epsilon,c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,c,t\right)\right\|^{2}\right].$ (18)
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\tilde{\alpha}_{t}}}\epsilon_{\theta}(y_{t},c,t)\right)+\hat{\beta}_{t}z.$ (19)
This approach is more widely used than guided diffusion models and is the basis for numerous models of important diffusion (e.g., DALL-E2 [45], Imagen [49], and stabilizing diffusion (SD) [18]).
IV-A2 Latent Diffusion Model (LDM)
The computational cost of direct manipulation of the diffusion model in the original image pixel space is high, Rombach et al [60] proposed to train and reason about the diffusion model in a low-dimensional latent space and named it Latent Diffusion Model (LDM).
Figure 7: Architecture of LDM, where E and D denote the encoder and decoder, respectively
As shown in Fig. 7, the LDM consists of three parts. First, the specific diffusion process of the LDM, which shifts the workspace from pixel space to latent space using a pre-trained autoencoder, is shown in the leftmost gray box in Fig. 7, where the noise is added to feature $\mathbf{z}$ to get the lower half of $\mathbf{z_{T}}$ as forward diffusion; the middle orange box is in latent space, where the U-Net with crosscutting attention supports the multimodal inputs in order to restore $\mathbf{z_{T}}$ to $\mathbf{z}$ as reverse diffusion. The right side box is the conditional encoder, which encodes various conditions (e.g., text and image) into feature vectors $\tau_{\theta}$ and feeds them into U-Net to generate an image.LDM is able to effectively reduce computational cost without degrading the quality of image synthesis while speeding up the inference speed, which is widely used in natural scenes and provides inspiration for the application of diffusion models to hyperspectral applications.
IV-B Application of Diffusion model in image processing
IV-B1 Data Enhancement and Synthesis
In data enhancement and synthesis, usually super-resolution techniques have an important role and unique advantages in improving quality. Super-resolution aims at reconstructing high-resolution images from low-resolution images [69], but the images can be disturbed, due to imaging equipment, weather conditions or downsampling. The loss of detail in hyperspectral images can be significant when compared to natural images, making inverted reconstruction of high-resolution images more challenging. Before the advent of diffusion models, deep neural network based approaches, e.g., Convolutional Neural Networks (CNNs) and Transformers, were finding suitable mapping functions to integrate feature information into a single step. To achieve this goal, methods including self-supervised learning [70], adaptive marker selection [71] and frequency information assistance [72] were used. However, the diffusion model-based approach fuses low-resolution images and bootstrap images in each step of the diffusion process, which helps to better integrate different image information and, at the same time, has gained wide attention in the hyperspectral field.
Although HSI has a high spectral resolution, its spatial resolution is relatively low [73], which may limit the performance of various HSI-based applications. To obtain high-resolution hyperspectral maps, there are two types of methods: panchromatic sharpening [74] and multispectral and HSI fusion [75]. The former enhances HSI by injecting detail information in the panchromatic (PAN) image, while the latter utilizes multispectral images to help hyperspectral maps learn spatial details. For example, Shi et al [76] used multispectral and hyperspectral images in tandem as conditional inputs to a diffusion model, and then generated hyperspectral maps with high resolution from the information found in the two image modalities.Qu et al [77] devised a conditional recurrent diffusion framework that allows the spatial SR process and the spectral SR process to be guided by the multispectral image and the hyperspectral map, respectively, to generate high-resolution hyperspectral maps with complementary spatial and spectral information to generate high-resolution hyperspectral maps.
More contemporary research focuses on the use of PAN image HSI super-resolution [78, 79, 80, 81, 82, 83, 84].Meng et al [79] proposed a modal calibration module to enhance and extract features in two images, where the enhanced features are used as conditional inputs to the diffusion model.Cao et al [78] argued that the different image modal unique information should not be mixed for processing and designed two conditional modulation modules to extract coarse-grained style information and fine-grained frequency information as conditional inputs.
Figure 8: Converged network fusion mode [82]
In order to take full advantage of the unique information of different modalities, Li et al [82] proposed a panchromatic sharpening network based on a biconditional diffusion model, where HSI and PAN are used as independent conditional inputs to learn the spectral features and spatial texture, respectively. The network structure is designed based on LDM and achieves excellent SR performance while significantly reducing the computational cost.Rui et al [84] also perform the sampling process in low-dimensional space. Based on the fact that the high-resolution hyperspectral map can be decomposed into the product of two low-rank tensors, one of the low-rank tensors is first computed using the low-resolution hyperspectral map. This tensor, together with the LRHSI and PAN, is then conditioned and fed into a pre-trained diffusion model [85] for the other low-rank tensor. Unlike the above methods, this method is a completely unsupervised deep learning method that does not require the involvement of high-resolution hyperspectral maps in any of the processes, providing feasibility for its application in practice.
IV-B2 Data Generation and Completion
Generation and Complementation in general can be categorized as text-to-image and image-to-image, but in the hyperspectral domain only image-to-image exists. For diffusion models, diffusion models have image bootstrapping to generate new images. Guided images can be in various forms such as maps [86], semantic layouts [87, 88, 89, 90]. Even though these images contain only some specific information, they still show excellent performance in image generation. For example, Espinosa and Crowley [86] proposed to use maps and historical maps to train conditional diffusion models, which generated more realistic satellite images.Yuan et al. [87] generated high-quality hyperspectral images under the guidance of semantic masks, and also solved the intrinsic problem that diffusion models require a long time of training for model convergence. However, these things alone are not enough, so it is important to generate not only high-quality images, but also image-annotation pairs that are more practical. For example, Zhao et al [89] implemented a two-stage fine-tuning of the SD model to enable the generation from noise to annotations and then from annotations to images.Toker et al [90] extended the standard diffusion model to a joint probabilistic model for images and their corresponding labels, which enabled the simultaneous generation of labels and images.
Hyperspectral images (HSI) can provide richer spectral spatial information. Meanwhile, the acquisition cost of HSI is much higher than that of multispectral images. Therefore, subsequent scientists then used easily available multispectral remote sensing images to generate hyperspectral images [91, 92]. However, the hyperspectral maps generated by the diffusion model used need to match the input noise dimensions with the spectral bands, which leads to too much noise sampling space, thus making the model convergence difficult. To solve this problem, Zhang et al [91] proposed a spectral folding technique to convert the input hyperspectral map into a pseudo-color image before training the diffusion model.Inspired by the LDM, Liu et al [92] used Conditional Vector Quantization GAN (VQGAN) [93] to obtain the Latent Diffusion Model latententent space for the hyperspectral map, and performed the diffusion model training and sampling in this space process. Afterwards, in order to address the status quo of high cost and low spatial resolution, a DM-based hyperspectral spectral decomposition method, DiffUn [94], was proposed by Deng et al. The spectral prior distributions learned by unconditional DM and the likelihoods derived from the LMM were combined to facilitate the deblurring performance and synthesize the hyperspectral maps.A novel ISPDiff [95] model was proposed by Dong et al. A transparent unified scale-flexible model was designed to solve such problems using an iterative optimization algorithm with MSRN for scale-matching recovery and CMUN for cross-scale upsampling.
Overall, image-to-image generation methods based on conditional diffusion models [67, 68], which use bootstrap images as conditional inputs to conditional DDPM, are essentially generating images from noise rather than directly. In order to realize a true image-to-image conversion, Yu et al. proposed UnmixDiff [96] to address the feasibility of depth generation model in HSI synthesis task, which synthesizes HSI from the perspective of material distribution and introduces a clear physical interpretation into the hyperspectral image synthesis task. By introducing an unmixing autoencoder in the diffusion generation model, the hyperspectral image synthesis is projected from the high-dimensional spectral domain to the low-dimensional abundance domain, which greatly reduces the computational complexity and effectively improves the spectral confidence of the generated hyperspectral images. By constructing a diffusion generation model in the abundance space, at the same time, hsi with real spatial details can be generated to maintain spectral consistency. Also at the same time this type of approach, which first employs self-supervised learning to extract the feature embeddings of the input patches, and then uses these embeddings as conditional inputs to guide the learning of the diffusion model. Notably, this preserves the spatial arrangement of each color patch in the original image so that the generated patches can be assembled into a large and coherent image.
Figure 9: UnmixDiff generates a simple diagram of the HIS [96]
IV-B3 Classification
Among the applications in the hyperspectral field, classification task is very important in image analysis as it can accurately identify and characterize different structures and anomalies in an image. It helps professionals in their ability to interpret large amounts of complex data Despite this potential, image classification aims at obtaining more valuable information about detailed features, such as buildings or grass, by assigning each pixel to a specific class [97, 98]. However, hyperspectral images contain many diverse and complex scenes and information, which increases the difficulty of accurate classification. Therefore, the current use of diffusion model to enhance classification is still a major challenge that needs to be further strengthened and solved. Meanwhile, because diffusion models can learn and simulate more complex data distributions than other deep learning models, more and more people are committed to applying diffusion models to hyperspectral image classification, which leads to better and better classification results [99, 100, 101, 102, 103, 104].
For image classification, hyperspectral images exhibit more complex data distributions, posing greater challenges to the application of traditional deep learning models in classification. The good thing, however, is that diffusion models can better capture the spectral-spatial joint features of hyperspectral images [105, 106, 107], improve the classification accuracy [108, 109], and improve the generalization ability [110]. Based on the diffusion model to realize the above properties, Zhou et al [111] used the temporal information of the diffusion model to build a time-stepping feature library, and also proposed a dynamic fusion module to combine spectral-spatial features with temporal features, and these improvements can make it possible to obtain sufficient image information before classification.Li et al [112] designed a two-branched diffusion model, which extracts features from HSI and LiDAR images, which achieves complementarity of information between different modal images, enhances the distinguishability between pixels, and also shows excellent performance in more complex hyperspectral tasks [37].Qu et al [113] go deeper by setting the encoder of the two-branch diffusion model to operate in parameter-sharing mode, which ensures that the shared features are extracted from the multimodal images. In addition, Chen et al [114] utilized the diffusion model to assist deep subspace construction and achieved excellent HSI classification performance in an unsupervised manner. Different from the above approaches, Qu et al [115] used a diffusion model to combine the SR (super-resolution) network and classification network to progressively reconstruct high-quality images by iteratively iterating the image super-resolution and classification tasks to achieve mutual guidance of the two tasks. The multi-scale classification results are utilized to guide the target-scale image to learn new knowledge and further reconstruct high-quality images. The super-resolution performance and classification performance are significantly improved.Jiang et al. proposed an image fusion [107] classification network based on Denoising Diffusion Probabilistic Models (DDPM), and after fusion, the classification accuracy is significantly improved.Zhu et al. proposed a Spatial-Spectral Diffusion Comparison Representation Network (DiffCRN) based on Denoising Diffusion Probabilistic Models (DDPM) [102], which is structured spatially self-attentive denoising module (SSAD) and spectral group self-attentive denoising module (SGSAD), and improves the efficiency of spectral spatial feature learning by designing new DDPM models loss and CL with log absolute error (LAE), as well as designing an adaptive weighted addition module (AWAM) and cross-time step spectral spatial fusion module (CTSSFM), which can effectively mitigate the interference due to spatial spectral heterogeneity and noise effects, etc., to effectively mitigate the interference due to spatial spectral heterogeneity and better classification.
Figure 10: Diffusion model based image classification process diagram
IV-B4 Segmentation
Image segmentation is an important task in computer vision, like semantic segmentation [116], image segmentation [14, 117], and in the field of deep learning, semantic segmentation is a subtask of image segmentation [112, 118]. Hyperspectral image segmentation, which investigates simplifying the complexity of an image by decomposing it into multiple meaningful image segments. Specifically, it facilitates image analysis by providing useful information about hyperspectral related fields. However, deep learning models typically require a large number of different pixel-annotated training data to produce generalizable results. Nonetheless, the number of images and labels that can be used for hyperspectral image segmentation is limited due to the time, cost, and expertise required. For this reason, diffusion modeling has emerged as a promising approach in image segmentation research that synthesizes labeled data and eliminates the need for pixel annotation data [102, 104].
Dong et al [95] are using Denoising Diffusion Probabilistic Models combined with Scale Matching Recovery Networks (P2MSRN) and cross-scale model-driven up-sampling networks.Chen et al [119] design image conditional mixing guides and combine them with Denoising Diffusion Probabilistic Models.Zhang et al [110] utilize Diffusion Probabilistic Models for labeling information.Zhu et al [120] utilize Variational Auto-Encoder as well as semi-supervised diffusion model for labeling annotation.Li et al. proposed a multidomain diffusion-driven feature learning network (MDFL) [112] to improve the pixel-level classification (i.e., segmentation) performance of high-dimensional remote sensing images through a multidomain diffusion model and feature reuse mechanism.Zhu et al. proposed a diffusion-based diffusion model coupled with contrast learning for diffusion contrast representation networks (DiffCRNs) [102] to extract the spatial-spectral features and finally generates pixel-level classification circles. Xu et al. proposed a data-enhanced semantic segmentation framework based on the diffusion model [104], which generates pseudo-labeled data through the diffusion model, and uses ControlNet to achieve cross-modal generation of “label→multispectral image”, and multi-stage training, which improves the segmentation accuracy.
IV-B5 Target Detection
Target detection differs from pixel-level ground classification in that target detection for hyperspectral images uses bounding boxes to localize instances like an airplane, a vehicle, or a ship in the image [121, 122]. One of the most challenging problems in target detection is that the training data is ambiguous, disturbed by external conditions [123, 118], and lacks realism and diversity. This problem is due to the intrinsic nature of hyperspectral images, which usually results in objects that need to be detected being complex, difficult to separate, and distributed in different regions of the image. Therefore, augmenting the objects to be detected with diffusion models has become an effective solution.
In order to solve all the above mentioned problems, Yu et al. proposed Diff-Mosaic (Enhanced Target Detection via Diffusion Prior), a diffusion model based approach [118], which introduces an enhancement network called Pixel-Prior, which coordinates pixels to generate highly coordinated and lifelike mosaic images, using a powerful diffusion model prior that resamples the pixels with samples to introduce real-world information. This resampling makes the generated samples realistic and diverse. Similarly, Qi et al. proposed a novel joint framework for underwater hyperspectral image restoration and target detection based on conditional diffusion framework [123], which uses a variable spectral group extraction module and a joint underwater hyperspectral image restoration and target detection (JURTD) module to simulate different ranges of underwater target spectra, achieve high-quality restored images and superior detection performance, and effectively overcome the challenges of underwater hyperspectral imaging. spectral imaging challenges.Zhu et al. proposed ODGEN (Domain-specific Object Detection Data Generation with Diffusion Models), a novel approach [124] to generate high-quality images conditioned on bounding boxes, thereby facilitating data synthesis for object detection.The use of synthetic visual cues with spatial constraints and textual descriptions by object to control the diffusion model allows ODGEN to perform well at demonstrate robustness when dealing with complex scenes and specific domains. Meanwhile, in order to improve the precision and accuracy, Wang et al. proposed OrientedDiffDet (pairwise object-oriented diffusion model) [125], which combines oriented frame representation, feature alignment and accelerated diffusion, a process that detects objects directly from a set of random frames. In the training phase, the horizontal detection boxes are first converted to oriented detection boxes. The model then learns to reverse this conversion process by diffusing from ground truth oriented boxes to a random distribution. In the inference phase, the model progressively refines a set of randomly generated boxes to generate the final output.
Figure 11: OrientedDiffDet target detection diagram with targets boxed in different colors in the detection diagram [125]
IV-B6 Anomaly Detection
Image anomaly detection is an important topic in computer vision, aiming to highlight anomalous regions of an image [115, 126, 127, 128]. In recent years, generative modeling has greatly influenced the queries for anomaly detection and shown promising results. Therefore, it will be explored in this section how the diffusion model, as the main generative model in anomaly detection, can be applied in hyperspectral collar anomaly detection. In fact, there are relatively few studies on hyperspectral image anomaly detection based on diffusion modeling, however, there are relatively numerous applications in fields such as medical impact areas [129, 130], industrial images [131, 132], etc., so with the development of diffusion modeling in this hyperspectral image anomaly detection will be increasingly used.
Meanwhile, in the field of anomaly detection, it is often faced that on the one hand, the complexity of the background of the input hyperspectral image leads to the confusion of anomalies in the background. On the other hand, the lack of labeled samples for hyperspectral images leads to poor generalization of existing anomaly detection methods. To solve this problem, there Ma et al. proposed a novel solution for anomaly detection, BSDM (Background Suppression Diffusion Model) [133], which simultaneously learns the Latent Background distribution and generalizes it to different datasets to suppress the complex background. It is characterized by three aspects:(1) Pseudo background noise is designed for the complex background of HSIs and Diffusion Model (DM) is used to learn the latent background distribution therein. (2) For the generalizability problem, a statistical bias module is applied to adapt the BSDM to datasets with different domains without labeling the samples. (3) To achieve background suppression, the inference process of DM was innovatively improved by feeding the original hyperspectral image into a denoising network, which removes the background as noise. They were also the first to apply diffusion modeling to hyperspectral image anomaly detection. Wu et al. proposed the diffusion background dictionary method (DBD) [134], which combines the advantages of model-driven and data-driven based on the characteristics of HSI and HAD tasks.DM is intrinsically combined with the low-rank representation (LRR) model, and DM is utilized to obtain the key background dictionary tensor in the tensor LRR. Thus, anomalies can be accurately detected. According to the idea of RX algorithm in anomaly detection, the multivariate normal distribution approximating the HSI background is diffused to make it more suitable for suppressing the background. Also hyperspectral image anomaly detection is often faced with the lack of a priori information, essentially the challenge is to accurately reconstruct the background region while inferring the latent background of the anomalous region, due to which Chen et al. proposed a novel approach called Double Window Spectral Diffusion (DWSDiff) [56], which modeled accurate background estimation through an iterative spectral diffusion and reverse diffusion process. A dual-window strategy is also introduced to mitigate the effect of anomalous extended regions in the neighborhood on the background estimation. An anomaly generation strategy based on Principal Component Analysis (PCA) and Linear Spectral Mixture Model (LSMM) is devised to create realistic training data that overcomes the scarcity of labeled hyperspectral data. The same effect is improved.
Figure 12: Various diagrams under BSDM from left to right: anomaly diagram, 1D t-SN diagram, 2D t-SNE scatter diagram [133]
In conclusion, the road of diffusion model in hyperspectral image anomaly detection is still long and the research prospect is very great, more and more researchers will put their eyes here to promote the vigorous development of the field of anomaly detection.
IV-B7 Noise Suppression
The main challenge in hyperspectral imagery is to acquire images without losing important information. During the acquisition and/or further processing stages, the acquired image may be corrupted by noise or artifacts. Noise degrades the image quality and is especially noticeable when the objects are small and the contrast is relatively low. Due to the nature of the generative model, the diffusion model facilitates the solution of various noise reduction problems [135, 136, 137]. Among them, [136] is denoising in image restoration, where the robustness of denoising is enhanced by designing a multi-shape spatial rectangular self-attention module, a spectral latent diffusion enhancement module, and also by utilizing the diffusion model, which allows for a tighter alignment with the desired HSI.
In image processing, hyperspectral images are always accompanied by various noises due to the inherent limitations of imaging techniques and environmental conditions. In essence, the learning process of diffusion models is equivalent to the denoising process [19], which makes it possible to achieve excellent denoising performance in hyperspectral contexts. Different image modalities face different noise challenges. For example, hyperspectral image (HSI) recovery, which may be quite susceptible to distributional biases, is data-poor [138]. HSI needs to take into account the inhomogeneous distribution of noise in the spectral dimension [139] as well as the correlation of noise between different frequency bands [140, 141]. Fortunately, these noises can be efficiently removed by a diffusion model.Ceng and Cao et al. proposed Diff-Unmix [142] to address the challenge of recovering noise-degraded HSI through the fusion of spectral unmixing and conditional abundance generation, which employs a learnable block-based spectral unmixing strategy supplemented by a purely transformer-based backbone. We then introduce a self-supervised generative diffusion network to augment the abundance maps from the spectrally unmixed blocks. This network reconstructs noise-free decomposition probability distributions, effectively mitigating noise-induced degradation within these components. Finally, the reconstructed HSI is reconstructed by de-mixing reconstruction by mixing the diffusion-adjusted abundance maps with the spectral end-elements. he et al [140] proposed a truncated diffusion model that starts denoising from the intermediate step of the diffusion process instead of the pure noise in order to avoid destroying the inherently valid information in the HSI. In addition, Deng et al. proposed RGB image-enhanced hyperspectral Denoising Diffusion Probabilistic Models [143], which improved DM-based HSI denoising by combining RGB-DM pre-training on additional RGB images, and a fusion operator was introduced to integrate HSI-DM and RGB-DM, which improves the performance of DM-based HSI denoising.Kai et al. proposed a multi-scale spatial spectral for hyperspectral image frequency joint denoising network (MSFDN) [137], in order to describe and extract the noise more accurately and obtain better denoising effect, they constructed a subspace projection network to convert the hyperspectral image into a low-dimensional subspace, designed a noise separation network, a subspace inverse projection network and utilized the conditional diffusion model for better robustness.
Figure 13: Process diagram of noise suppression based on diffusion model
IV-B8 Data Recovery
Hyperspectral image (HSI) recovery aims to recover clean images from degraded observations [144], and the recovery of hyperspectral images plays a pivotal role in subsequent hyperspectral image applications. Hyperspectral data recovery faces challenges in effectively exploring the spatial nonlocal self-similarity and spectral low-rank properties inherent in HSI, while hyperspectral exists data scarcity. Existing model-based approaches have limitations in accurately modeling complex image features with manual a priori, and deep learning-based approaches have poor generalization capabilities, complex models, poor results, and parameter redundancy, Diffusion models, due to their impressive image recovery performance [145, 146, 147, 148, 149, 150, 151]. In particular, it has recently attracted attention in terms of noise robustness [92, 95, 140, 152, 115, 119]. Diffusion methods, however, are trained on large amounts of training data and perform very well within distributions, but may be quite susceptible to distributional bias.
To address the above problem, Miao et al. proposed a self-supervised diffusion model for HSI recovery, the Denoising Diffusion Probabilistic Models for Spatial Spectroscopy (DDS2M) [138], which works by inferring the parameters of the proposed variational spatial spectral module (VS2M) during reverse diffusion, using only the degraded HSI without any additional training data. A loss function based on variational inference is customized to enable untrained spatial and spectral networks to learn a posteriori distributions that serve as transitions in the sampling chain to help reverse the diffusion process. Chen et al [56] use a dual window, iterative diffusion model, and reverse reconstruction to mitigate the impact of background estimation. Li et al. propose an untrained diffusion model with a pre-trained diffusion model (HIR-DIFF) as a A supervised HSI recovery framework [141], which can recover clean HIS from the product of two low-rank components, is proposed based on singular value decomposition (SVD) and rank-revealed QR (RRQR) decomposition, and a novel exponential noise schedule is proposed to speed up the recovery process. Li et al. propose a spectral latent diffusion enhancement module [136], which introduces a latent diffusion-enhanced rectangular transformer for HSI recovery. A multi-shape spatial rectangular self-attentive module is proposed, which leads to a tighter alignment with the desired HSI.
Meanwhile, there are also data recovery problems like noise mixing characterized by signal correlated and signal uncorrelated components, visual perception is affected and discriminative image distortion. Si et al. proposed a new model of CASSIDiff to reconstruct CASSI metrics [153], which employs a feature fusion mechanism based on the Discrete Wavelet Transform (DWT) and a spatial-spectral attention mechanism, and has a high degree of reliability in the recovery task. reliability. Yi et al. proposed an efficient diffusion model (EDiffSR) [149] for RSI SR, developed an efficient activation network (EANet), and developed a practical conditional prior enhancement module (CPEM), which reduces the computational budget and retains more information. Meng et al. proposed a conditional diffusion model-based SRSISR method (FastDiffSR) [148], which reduces the number of sampling steps, employs residual images, and combines channel attention and spatial attention to obtain better visualization with significantly improved generation quality.
Figure 14: Diffusion model-based research progress in hyperspectral image processing and analysis the tasks are established in nine subfields: (1) data enhancement and synthesis, (2) data generation and complementation, (3) classification, (4) segmentation, (5) anomaly detection, (6) target detection, (7) noise suppression, and (8) data recovery. No abbreviated name, with two initial letters at the beginning and end
V Challenges and Limitations of the Diffusion Model
V-A Computational complexity and Time cost of the model
Diffusion Models achieve data generation and denoising through multi-step iterative Markov chains, whose computational complexity increases exponentially with the data dimensions. Taking the denoising diffusion probabilistic model (DDPM) [19] as an example, hundreds of forward and reverse diffusion computations are required to generate a single hyperspectral image, especially when processing hyperspectral data above 200 bands, and the GPU memory occupancy can be more than 20 GB (e.g., 1024 × 1024 resolution scenes). Although the improved DDIM [35] reduces the number of sampling steps to less than 50 through non-Markov chains, the spatial-spectral three-dimensional nature of hyperspectral data (see Section II.B) still leads to a single inference time as long as several minutes, which makes it difficult to satisfy real-time demanding tasks, such as agricultural monitoring and disaster warning. Although the Latent Diffusion Model (LDM) [18] reduces the computational volume by compressing the latent space, the information compression in the spectral dimension may lose key features (e.g., mineral absorption peaks), which affects the classification and anomaly detection accuracy (see Table II).
V-B Impact of high dimensionality of hyperspectral data on model training
The performance of diffusion models is highly dependent on large-scale labeled datasets, and the cost of labeling hyperspectral data far exceeds that of traditional RGB images. For example, field sampling is required to validate crop classes in agricultural applications, and the cost of labeling a single image can be hundreds of dollars. Although DDPM [19] generates synthetic samples through data augmentation, its fidelity is limited by the distribution of training data, and it is prone to Mode Collapse in scenarios with fewer samples (e.g., rare mineral detection). Experiments show that when the training samples are less than 1000, the classification accuracy of the diffusion model decreases by about 12% (F1-score from 0.85 to 0.75) compared to Variational Auto-Encoder (VAE) [9]. In addition, conditional diffusion models based on classifier guidance (e.g., ADM-G [11]) need to pre-train high-precision classifiers, but the class imbalance problem of hyperspectral data (e.g., less than 1% of anomaly samples in anomaly detection) leads to difficulties in classifier training, which restricts the model generalization ability.
V-C Adaptability issues to large-scale data
Real hyperspectral images often contain non-Gaussian noise (e.g., sensor thermal noise, atmospheric scattering noise), and most of the existing diffusion models are based on additive Gaussian noise assumptions (see Section III.A.1) The forward process design of the DDPM [19] is only applicable to isotropic Gaussian noise, and the denoising effect is significantly degraded for streak or PoissonGaussian hybrid noise (PSNR reduction of 8 15 dB). Although Score-SDE [21] extends the scope of noise modeling through stochastic differential equations, its inverse process relies on the accurate estimation of the score function, which is susceptible to error accumulation in bands where the signal-to-noise ratio (SNR) fluctuates drastically (e.g., near-infrared bands). For example, the Diffusion Models’ denoising performance degrades by about 20% compared to the conventional wavelet approach under extreme conditions where the noise standard deviation exceeds 30 dB. In addition, the mixed image element problem (see Section II.C) leads to a high degree of spatial coupling between the noise and the target signal, further exacerbating the modeling difficulty.
V-D Limitations of current models in practical applications
The black-box nature of diffusion models exposes them to credibility challenges in hyperspectral critical applications. For example, in anomaly detection tasks, the model may misclassify rare features as noise by overfitting the training data (literature [16]). Experiments have shown that LDM [18] based anomaly detectors have a 18% higher false alarm rate than the traditional RX algorithm (see Section II.C) in the detection of unknown mineral types. In addition, although the Latent Diffusion Model improves the computational efficiency, the physical meaning of the latent vectors is unclear, which makes it difficult for users to trace the generation logic of the detection results. In scenarios that require strict validation, such as military target identification, the interpretability disadvantage of diffusion models significantly constrains their practical application. Comparative studies have shown that traditional methods (e.g., SVM) [154] can provide a decision basis through feature importance analysis, whereas diffusion models can only output probability estimates and lack a transparent inference chain.
In summary, the practical deployment of diffusion models in hyperspectral image processing is still limited by core bottlenecks such as computational efficiency, data dependency, noise robustness and interpretability. Future research needs to combine hyperspectral data characteristics (e.g., spectral continuity, spatial-spectral correlation), develop lightweight architectures and adaptive noise modeling methods, and explore semi-supervised learning frameworks to reduce annotation costs, so as to promote its large-scale application in remote sensing, agriculture and other fields.
VI Comparative and Experimental analysis
In order to effectively illustrate the superiority of the diffusion model in processing hyperspectral images, this paper takes the experimental results of classification, target detection, anomaly detection, image recovery, and noise suppression as an example, and evaluates the performance of the diffusion model and other existing techniques through the visualization results and quantitative indexes.
TABLE II: Comparison of Diffusion Model for Hyperspectral Images with Traditional Deep Learning Models
Deep Learning Model Advantages Disadvantages Future Development Trends
Diffusion Model Generates high-quality image details. Specializes in denoising and data reconstruction. Suitable for dealing with high-dimensional features of hyperspectral images. High computational cost. Slow training and inference. Sensitive to hyperparameters. Optimizing computational efficiency. Combining physical models to improve reconstruction accuracy. Lightweight design to adapt to real-time processing needs.
Generative Adversarial Network (GAN) Generate realistic hyperspectral data to support data augmentation. Enhance robustness by adversarial training in classification tasks. Training is unstable (pattern crashes). Generating data may have artifacts. Improved training stability (e.g., Wasserstein GAN). Combined with attention mechanisms to enhance local feature generation.
Variational Auto-Encoder (VAE) Effective degradation and feature extraction. Generating latent space representations. Supporting unsupervised learning. Generated images are fuzzy and do not retain enough detail. limited ability to model complex distributions. Improved decoder structure for enhanced detail generation. Combined with diffusion model to improve generation quality.
Autoregressive Models (AR) Strong sequence modeling capability for pixel-by-pixel prediction. Excellent performance in temporal hyperspectral analysis. High computational complexity (cannot be parallelized). Difficult to handle large size hyperspectral images. Introducing sparse attention mechanism to reduce computation. Combining with Transformer architecture to improve the efficiency of long sequence modeling.
VI-A Experimental results and Performance assessment and Various metrics
VI-A1 Introduction to evaluation metrics and Datasets
The receiver operating characteristic curve (ROC) is a qualitative analysis of the model. The horizontal coordinate of the curve is the false positive rate (FPR) and the vertical coordinate is the true positive rate (TPR).
AUC (area under ROC curve) is the area under ROC curve, which is a quantitative analysis of the model. In practice, AUC is often used as a metrics for evaluating the accuracy of a method. The larger the AUC value, the better the detection performance of the algorithm, with mAUC representing the average AUC. where larger ${\rm AUC}_{(D,F)}$and ${\rm AUC}_{(D,\tau)}$ indicate better detection performance, and are combined together defined as ${\rm AUC}_{TD}$; ${\rm AUC}_{BS}$ represents the ability of the detector to minimize false alarms due to background interference; ${\rm AUC}_{ODP}$ denotes the overall probability of detection; ${\rm AUC}_{TD-BS}$ specifically evaluates the performance of the detector in recognizing a targets in background noise, taking into account both detectability and background suppression; ${\rm AUC}_{SNPR}$ reflects the detector’s ability to distinguish between signal and noise.
OA denotes the ratio of the number of correctly classified samples to the total number of samples in all test samples; AA denotes the average of the classification accuracy for each category; k × 100 is the Kappa coefficient is a measure of the consistency of the classifier, which takes into account stochastic consistency and provides a measure of the consistency between the classifier and the ground truth.
IoU is a measure of the degree of overlap between the predicted frame and the real frame; $\mathbf{P}_{d}$ is defined as the ratio of the number of correctly detected targets to the total number of targets; and $\mathbf{F}_{a}$ is defined as the ratio of the number of false alarms to the total number of pixels in the image.
PSNR is the Peak Signal to Noise Ratio, a measure of the difference in quality between a super-resolution processed image and the original high-resolution image.FID (Fréchet Inception Distance) is a metrics used to assess the quality of the generated images, measuring their similarity to each other. Patch Similarity evaluates the perceptual difference between the generated image and the original high-resolution image. parmas is the number of parameters for model training. entropy is usually used to measure the uncertainty and complexity of image information. niqe is the Natural Image Quality Evaluator, which evaluates the image quality metrics. SSIM (Structural Similarity) measures the similarity between two images, comparing images in terms of brightness, contrast and structure; FSIM (Feature Similarity) evaluates the image quality by calculating the phase coherence and gradient similarity in localized regions of the image, and is an improved version of SSIM; SAM (Sum of Absolute Errors) is another measure of the effectiveness of image denoising, which calculates the sum of the absolute value of the pixel differences between a denoised image and the original image counterparts.
TABLE III: Summary of datasets used for the following experiments
Datasets Sensor Spectral Band Size Resolution
Salinas AVIRIS 204 512$\times$217 3.7 m
Pavia University ROSIS 103 610$\times$340 1.3 m
Indian Pines AVIRIS 200 145$\times$145 20 m
SIRST FLIR LWIR 512$\times$512 0.5 m
NUDT-SIRST Infrared Sensor LWIR 640$\times$512 0.3 m
Airport-Beach-Urban AVIRIS 205/188/191 150$\times$150 20 m
Potsdam ISPRS RGB(3) 5000$\times$5000 5 cm
Toronto Toronto RGB(3) 1920$\times$1080 0.2 m
UC Merced USGS RGB(3) 256$\times$256 0.3 m
CAVE Hyperspectral Camera 31 512$\times$512 N/A
CAVE-Toy Hyperspectral Camera 31 512$\times$512 N/A
KAIST Thermal Camera LWIR 640$\times$480 0.5 m
VI-A2 Experimental results
In this section, several typical methods from the above methods will be selected and compared to study the performance of detection, where traditional methods:GAN, VAE, autoregressive models, other deep learning methods, are compared with diffusion models. Since the experimental dataset of each algorithm is different, a common dataset is chosen for comparison as much as possible, in order to verify the anomaly detection performance of different algorithms on the same dataset in a more reasonable way. Where, the bolded indicates the algorithm’s performance optimal value for this dataset.
Comparison of classification experimental results, as shown in Figure 15, the classification effect of the first three methods is very poor, different classification color adulteration, the latter three methods overall classification effect is better, among them, the diffusion model-based method SpectralDiff classification effect is the best, followed by DMVL, their most boundary delineation is better versus the other methods, with less color adulteration, and SpectralDiff, despite the tiny classification there are some problems, but the overall effect is significantly better than other mainstream methods:
Figure 15: Comparison of different datasets on classification methods for hyperspectral images: (a) Pseudo-color image (b) Real labeled image (c) SF [155] (d) miniGCN [156] (e) SSFTT [157] (f) DMVL [158] (g) SSGRN [159] (h) SpectralDiff [108]
TABLE IV: Comparison of Metrics for Different Classification Methods on Indian Pines Dataset
Methods Evaluation metrics (Indian Pines)
OA AA K×$\times$100 Time(s)
SF 67.782 80.639 63.820 29.81
miniGCN 63.916 69.091 59.439 0.95
SSFTT 91.566 95.661 90.392 31.58
DMVL 90.487 81.311 89.216
SSGRN 92.336 96.099 91.254 1.13
SpectralDiff 93.146 96.438 92.175 48.53
Table IV shows that the Spectral-Diff method is in the first place in Indian Pines dataset classification in terms of overall accuracy, average accuracy, and kappa coefficient, which is significantly improved over other methods and is the current optimal method.OA and AA are higher than the second place (SSGRN) by 0.8% and 0.34%, which indicates that the method is more robust in classification and more effective in capture features and reduce noise interference. However, it can also be seen that Spectral-Diff consumes much more time than other methods, which is its shortcoming.
TABLE V: Comparison of Metrics for Different Classification Methods on Pavia University Dataset
Methods Evaluation metrics (Pavia University)
OA AA K×$\times$100 Time(s)
SF 81.511 84.942 75.795 163.34
miniGCN 82.355 84.037 76.620 2.40
SSFTT 93.667 93.448 91.617 110.08
DMVL 89.768 86.116 86.580
SSGRN 93.850 94.436 91.919 4.99
SpectralDiff 94.775 93.843 93.061 256.23
As can be seen in Table V, the method has two metrics on the Pavia University dataset that are optimal 0.98% and 1.23% higher than the second (SSGRN), and the other is only 0.63% lower than the optimal, which is a relatively small gap, and it is still the optimal method among these methods, but the shortcoming of the use of samples is the long time consumed, which is more than twice as long as the other methods.
TABLE VI: Comparison of Metrics for Different Classification Methods on Salinas Dataset
Methods Evaluation metrics (Salinas)
OA AA K×$\times$100 Time(s)
SF 88.248 93.263 86.973 141.30
miniGCN 88.181 94.279 86.823 2.13
SSFTT 95.789 98.272 95.322 55.96
DMVL 97.005 95.853 96.668
SSGRN 96.539 96.354 96.144 2.98
SpectralDiff 98.971 99.465 98.854 242.16
Table VI shows the classification on Salinas dataset and the same conclusion as the previous two tables, combined with Tables IV, V, and VI, the Spectral-Diff method is optimal in terms of performance compared to the other methods, but it is more time consuming and slower. SSGRN is optimal if balancing accuracy and efficiency, and miniGCN is optimal if low latency and high tolerance accuracy.
Comparison of target detection experimental results, as shown in Figure 16, the overall effect of these mainstream methods used in target detection is good, some methods exist undetected certain points or misdetection, among them, the diffusion model-based method Diff-Mosaic detection effect is the best, for the detection of the point of the edge of the border details of the best description, and the real labeling map is basically consistent, basically there is no missing parts:
Figure 16: Comparison of different datasets on target detection methods for hyperspectral images: (a) Original image (b) Real labeled image (c) ALC-Net [160] (d) ACM [161] (e) UIU-Net [162] (f) DNA-Net [163] (g) Diff-Mosaic [118]
TABLE VII: Comparison of Metrics for Different Target Detection Methods on SIRST Dataset
Methods Evaluation metrics (SIRST)
IoU $P_{d}$ $F_{a}$
ALC-Net 70.00 71.71 23.91
ACM 76.17 86.31 16.07
UIU-Net 72.03 98.10 26.15
DNA-Net 76.97 95.41 3.54
Diff-Mosaic 79.44 99.99 3.19
Table VII shows that the Diff-Mosaic method has the optimal intersection ratio, detection probability, and false alarm rate for target detection on the SIRST dataset, which is 3.11% higher than the second place (DNA-Net), 1.89% higher than the second place (UIU-Net), and 87.8% lower than the second place (UIU-Net), respectively. Among them, the detection tends to 100% detection is very good with less error and is the best among these methods.
TABLE VIII: Comparison of Metrics for Different Target Detection Methods on NUDT-SIRST Dataset
Methods Evaluation metrics (NUDT-SIRST)
IoU $P_{d}$ $F_{a}$
ALC-Net 72.13 78.81 21.22
ACM 71.11 85.13 22.51
UIU-Net 89.00 98.73 6.02
DNA-Net 88.38 97.99 4.04
Diff-Mosaic 91.18 99.47 1.91
Table VIII shows the Diff-Mosaic in manual synthesis dataset NUDT-SIRST utilization, which is consistent with the metrics effect shown in Table 7, and moreover, it shows the superiority of this method among these methods with the best performance.
TABLE IX: Comparison of Parameters and Reasoning for Different Target Detection Methods
Methods Calculate the number of parameters Time (s)
ALC-Net 0.38 40.93
ACM 0.29 18.53
UIU-Net 4.7 43.42
DNA-Net 50.54 33.98
Diff-Mosaic 4.7 43.42
Table IX shows that Diff-Mosai is less in the number of parameters, but not the best, while the inference time is the slowest among them, showing the shortcomings of the method in applications, 57.3% higher than the fastest time (ACM). Meanwhile, combining Tables VII, VIII and IX, again Diff-Mosaic is the best if pursuing the best performance, DNA-Net is the best if balancing performance and efficiency, and ACM is the best if high tolerance and low latency.
Comparison of the results of the anomaly detection experiments, as shown in Figure 17, compared with the traditional mainstream methods, the diffusion model-based methods BSDM, DWSDiff in the anomaly detection of the best detection effect, high detection rate, low false detection rate, and the real labeling map is basically consistent with the other mainstream detection methods to detect the detection of the detection effect is poor, the detection rate is low (the existence of the detection of the anomalies do not exist), the false detection rate is high:
Figure 17: Comparison of different datasets on hyperspectral image anomaly detection methods: (a) Pseudo-color map (b) Real label map (c) GAED [164] (d) RGAE [165] (e) Auto-AD [166] (f) DFAN-HAD [167] (g) DirectNet [168] (h) BSDM [133] (i) TGFA [169] (j)DWSDiff [56]
TABLE X: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{(p,r)}$ $mAUC_{(p,r)}$ $mAUC_{(p,r)}$
GAED 0.90756 0.15556 0.02072
RGAE 0.87615 0.14519 0.02401
Auto-Net 0.95028 0.15303 0.00627
DFAN-HAD 0.96552 0.45211 0.09603
DirectNet 0.95193 0.12799 0.00484
BSDM 0.94486 0.17342 0.03055
TGFA 0.92655 0.14559 0.00629
DWSDiff 0.97427 0.55329 0.01635
TABLE XI: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{(p,r)}$ $mAUC_{gs}$ $mAUC_{opp}$
GAED 1.06312 0.8868 1.04240
RGAE 1.02135 0.8521 0.99734
Auto-Net 1.10331 0.9440 1.09704
DFAN-HAD 1.41762 0.8694 1.32160
DirectNet 1.07993 0.9470 1.07508
BSDM 1.11828 0.9143 1.08773
TGFA 1.07214 0.9292 1.06585
DWSDiff 1.52756 0.9579 1.51121
TABLE XII: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{rp-gs}$ $mAUC_{svps}$
GAED 0.13484 13.79983
RGAE 0.12119 14.38906
Auto-Net 0.14676 52.05284
DFAN-HAD 0.35608 5.27753
DirectNet 0.12315 49.52139
BSDM 0.14287 8.40576
TGFA 0.13830 32.42845
DWSDiff 0.53694 67.20618
As can be seen in Table X, this table shows the accuracy of the anomaly detection methods, in which BSDM and DWSDiff are based on diffusion model, DWSDiff is in the first place in all other items except for the integrated detection performance (${mAUC}_{TD}$) which is second only to DFAN, while the data of BSDM is not outstanding (third in the integrated detection performance), but it shows the effect is not is not bad.
TABLE XIII: Comparison of Reasoning Time for Each Anomaly Detection Method
Methods Reasoning time (s)
Airport Beach Urban Average Time (s)
GAED 0.0300 0.0496 0.0257 0.0289
RGAE 0.0135 0.0251 0.0137 0.0141
Auto-Net 0.0020 0.0032 0.0020 0.0022
DFAN-HAD 0.0232 0.0448 0.0269 0.0264
DirectNet 0.0145 0.02839 0.0153 0.0145
BSDM 1.2306 1.2740 1.2509 1.2421
TGFA 1.6190 1.5505 1.5236 1.5392
DWSDUR 0.8016 0.8169 0.8192 0.8094
Table XI shows that both diffusion models have longer and slower inference times for each medium dataset scenario, with average times 99.8% higher (BSDM) and 99.7% higher (DWSDiff) than the shortest time (Auto-Net), which is hundreds of times higher than the optimal time, and is very slow in comparison. Again seeking performance DWSDiff is optimal and BSDM is a good choice.
Comparison of the experimental results of image recovery, as shown in Fig. 18, the overall effect of image recovery are relatively modest, but the methods based on the diffusion model (DDPM, EDiffSR, FastDiffSR) are better handled in terms of details, and the degree of recovery is closer to the high-resolution map:
Figure 18: Comparison of different datasets on hyperspectral image recovery methods: (a) High resolution map (b) HESNet [170] (c) TransENet [171] (d) NDSRGAN [172] (e) HAT [173] (f) DDPM [19] (g) EDiffSR [149] (h) FastDiffSR [148]
TABLE XIV: Comparison of Metrics for Different Image Recovery Methods on Potsdam Dataset
Methods Evaluation metrics(Potsdam)
PSNR FID LPIPS Params
HESNet 32.932 44.192 0.3014 5.4330
TransENet 32.768 41.310 0.2878 37.459
NDSRGAN 30.529 51.209 0.2137 17.510
HAT 32.987 44.932 0.2962 25.821
DDPM 31.989 24.307 0.0997 26.166
EDuFSR 30.860 36.669 0.1104 26.790
FastDuFSR 33.130 22.832 0.1152 23.229
As can be seen in Table XII, the three methods based on diffusion model (DDPM, EDiffSR, FastDiffSR) show unique advantages in key indicators, among which, the FsatDiffSR method has the best overall performance in terms of PSNR and FID, which is far superior to the traditional methods (compared to the second: (HAT) 0.43%, (DDPM) 6.07%, and (DDPM) 13.5% difference), indicating that its recovered image reaches a higher level in both detail fidelity and distributional consistency, but in terms of the number of parameters, the diffusion model-based methods are all higher than HESNet: 79.4%, 79.7%, 76.6%, respectively, thus affecting the training time too long.
TABLE XV: Comparison of Metrics for Different Image Recovery Methods on Toronto Dataset
Methods Evaluation metrics (Toronto)
PSNR FID LPIPS Params
HESNet 25.413 118.791 0.3234 5.5070
TransENet 25.565 97.958 0.3112 37.533
NDSRGAN 23.237 97.922 0.2534 17.510
HAT 25.368 119.05 0.2957 25.821
DDPM 24.733 66.583 0.1759 26.166
EDdfSR 23.990 81.412 0.1626 26.790
EasDdfSR 25.527 55.761 0.1561 23.229
Similar to Table XII, it can be seen from Table XIII that the three methods based on the diffusion model have a good overall performance, especially FastDiffSR has the best overall performance, and combining the two tables, the three diffusion methods have a higher quality of generation and more semantic information retention than the traditional methods. However, the three methods still have the problem of larger number of parameters in comparison.
TABLE XVI: Comparison of Measurement metrics for Different Image Recovery Methods
Methods Measurement metrics
Entropy NIQE
HESNet 7.0826 19.913
TransENet 7.0853 20.319
NDSRGAN 7.0528 19.214
HAT 7.1177 20.385
DDPM 7.1204 16.532
EDiIfSR 7.1838 17.641
EastDiffSR 7.1254 16.342
The entropy and generation quality metrics in Table XIV show that the three diffusion methods retain more semantic information and have higher generation and recovery quality, corresponding to Tables XII and XIII.Also combining the three tables, overall FastDiffSR has a balanced performance in terms of recovery accuracy and efficiency, DDPM is more advantageous in terms of perceptual quality, and EDiffSR can be used as a practical choice for moderate performance.
Comparison of noise suppression experimental results, as shown in Figure 19, with the noise map as well as the real noiseless map compared to the overall performance of the Diff-Unmix method has the best effect, followed by DDRM, and finally DDS2M are based on the diffusion model method, they are more successful noise suppression; in contrast to the traditional mainstream method effect is worse, to LLxRGTV in which the best:
Figure 19: Comparison of different datasets on hyperspectral noise suppression methods: (a) Noise map (b) True noise-free map (c) LLxRGTV [174] (d) FGSLR [175] (e) LRTDCTV [176] (f) DIP [177] (g) LLnRPnP [178] (h) HLRTF [110] (i) DDRM [179] (j)DDS2M [138] (k)Diff-Unmix [142]
TABLE XVII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,3))
PSNR SSIM FSIM SAM Time (s)
LLxRGTV 27.640 0.679 0.868 0.218 38
FGSLR 25.561 0.480 0.718 0.470 499
LRTDCTV 24.593 0.533 0.739 0.431 42
DIP 20.063 0.405 0.798 0.538 74
LLnRPhP 25.102 0.592 0.768 0.425 289
HLRTF 30.340 0.689 0.874 0.329 25
DDRM 27.810 0.786 0.893 0.387 23
DDS2M 30.078 0.666 0.834 0.355 318
Diff-Unmix 31.408 0.902 0.958 0.282 42
TABLE XVIII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,2))
PSNR SSIM FSIM SAM Time (s)
LLaRGUX 31.152 0.802 0.917 0.205 38
FGSLR 30.126 0.737 0.878 0.262 249
LRTDCTV 25.952 0.658 0.816 0.406 43
DIP 24.181 0.608 0.825 0.475 72
LLnRPhP 28.664 0.748 0.861 0.379 240
HLRTF 33.011 0.808 0.925 0.275 23
DDRM 29.412 0.865 0.922 0.293 20
DDS2M 32.804 0.786 0.895 0.334 354
Diff-Unmix 33.059 0.964 0.940 0.116 37
The last three methods (DDRM/DDS2M/Diff-Unmix) in the experiment are based on the diffusion model, as can be seen from Tables XV and XVI, the first choice of the KAIST two different noise situations in the dataset, these three methods in the noise suppression performance have their own characteristics, and the Diff-Unmix in the key indexes is the most outstanding performance, which PSRN, SSIM, and FSIM are the highest values, compared with other methods to show its significant advantages in noise suppression, structure preservation and spectral fidelity, but the time processing of the method under the diffusion model is slightly higher (HLRTF is the fastest), and the overall efficiency is still in the acceptable range.
TABLE XIX: Comparison of Metrics for Different Noise Suppression Methods on CAVE Dataset
Methods Evaluation metrics (CAVE)
PSNR SSIM FSIM SAM Time (s)
LLxRGTX 28.743 0.712 0.829 0.121 40
FGSLR 24.474 0.346 0.554 0.511 1758
LRTDCTV 24.394 0.444 0.631 0.315 47
DIP 20.480 0.554 0.864 0.313 66
LLnRPhP 24.809 0.445 0.668 0.284 246
HLRTF 30.308 0.770 0.861 0.185 133
DDRM 30.521 0.754 0.873 0.196 16
DDS2M 30.308 0.724 0.863 0.251 319
Diff-Unmix 32.714 0.940 0.957 0.129 43
Table XVII shows that the methods under the diffusion model are overall better than the other methods, of which Diff-Unmix remains the most prominent, demonstrating that the diffusion model method is overall superior to the other methods, with the best performance and the noise suppression presenting the optimal effect.
TABLE XX: Comparison of Metrics for Different Noise Suppression Methods on CAVE-Toy Dataset
Methods Evaluation metrics (CAVE-Toy)
PSNR SSIM FSIM SAM Time (s)
LLARGUX 26.450 0.752 0.874 0.118 39
FGSLR 21.525 0.573 0.798 0.241 1149
LRTDCTV 25.214 0.665 0.783 0.228 45
DIP 22.092 0.594 0.895 0.191 67
LLnRPhP 23.088 0.541 0.717 0.207 240
HLRTF 27.308 0.730 0.941 0.161 41
DDRM 27.886 0.858 0.910 0.159 16
DDS2M 29.344 0.844 0.977 0.111 320
Diff-Unmix 28.046 0.945 0.993 0.156 43
Table XVIII is roughly in agreement with Tables XV, XVI and XVII diffusion models method noise suppression performance is generally the best, and in combination with the above tables, the overall effect is highly acceptable, with obvious denoising effect, close to noiseless, and with better detail retention, despite the fact that the temporal processing is not particularly fast.
VI-B Key discoveries and Summaries from the Experiments
Through the experimental comparisons in Section V.B.2, the diffusion model shows significant advantages in hyperspectral image processing tasks, but also exposes some limitations.
In the classification task, the OA (93.15%, 94.78%, 98.97%), AA (96.44%, 93.84%, 99.47%) and Kappa coefficients of the SpectralDiff method based on the diffusion model on the Indian Pines, Pavia University and Salinas datasets ( 92.18%, 93.06%, 98.85%) are better than traditional deep learning methods (e.g., GAN, VAE), indicating that it can effectively capture the complex spectral-spatial features of hyperspectral data. However, its inference time (48.53 s, 256.23 s, 242.16 s) is significantly higher than that of lightweight models such as miniGCN (0.95 s, 2.40 s, 2.13 s), revealing the computational efficiency bottleneck of the diffusion model in high-dimensional data processing.
In the target detection task, the diffusion-based Diff-Mosaic method achieves optimal IoU (79.44%, 91.18%), detection probability (Pd: 99.99%, 99.47%), and false alarm rate (Fa: 3.19%, 1.91%) on the SIRST and NUDT-SIRST datasets, and is especially good at complex background suppression and target localization accuracy. However, the number of parameters (4.7M) and the inference time (43.42 s) are still higher than that of the traditional method ACM (0.29M, 18.53 s), which indicates that the diffusion model needs to be further lightweighted to meet the real-time demand.
In the anomaly detection task, DWSDiff comprehensively leads in mAUC metrics, such as mAUC(D,F) (0.97427), mAUC(BS) (0.95791), and mAUC(SNPR) (67.20618), which verifies its ability to suppress complex backgrounds through the dual-window strategy and iterative diffusion. However, its inference time (0.8094 s) is still much higher than that of Auto-AD (0.0022 s), and the model is sensitive to hyperparameters (e.g., noise scheduling), and its generalization ability needs to be further optimized, suggesting that the practical deployment of diffusion models in anomaly detection needs to balance accuracy and efficiency.
In image recovery tasks, diffusion model methods (e.g., FastDiffSR, EDiffSR, and DDPM) significantly outperform conventional methods in key metrics (PSNR, FID, LPIPS). For example, on the Potsdam dataset, FastDiffSR achieves a PSNR of 33.13 dB and a FID of 22.83, indicating that it outperforms conventional methods in both detail fidelity and distribution consistency. In addition, FastDiffSR achieves a PSNR of 25.53 dB and a FID of 55.76 on the Toronto dataset, which further validates its advantage in semantic information retention. However, the high number of parameters of the diffusion model method (e.g., 23.23M for FastDiffSR) leads to a significant increase in training time. It shows that the diffusion model has strong generative capability, but its computational efficiency and model complexity still need to be optimized to meet the actual deployment requirements.
In noise suppression tasks, diffusion model-based methods (e.g., Diff-Unmix, DDRM, and DDS2M) show significant advantages in metrics such as PSNR, SSIM, and FSIM. For example, on the KAIST dataset, Diff-Unmix achieves a PSNR of 31.41 dB, an SSIM of 0.902, and an FSIM of 0.958, indicating significant advantages in noise suppression, structure preservation, and spectral fidelity. In addition, on the CAVE dataset, Diff-Unmix has a PSNR of 32.71 dB and an SSIM of 0.940, further validating its superiority. However, the long inference time of the diffusion model method (e.g., 42 s for Diff-Unmix) restricts its application in scenarios with high real-time requirements, suggesting that the diffusion model is able to remove the noise and retain the details efficiently, but its computational efficiency still needs to be further improved.
Some key points are also found.1. Accuracy advantage: diffusion models are significantly mainstream deep learning methods (e.g., GAN, VAE) in tasks such as classification, target detection, and anomaly detection, especially in complex background suppression, high-dimensional feature modeling, and noise robustness.2. Computational cost: the computational complexity and inference time of diffusion models are generally high, e.g., SpectralDiff The inference time on Salinas dataset reaches 242 seconds, which is far more than miniGCN (2.13 seconds), limiting its application in real-time scenarios (e.g., disaster monitoring).3. Model adaptation: diffusion models are sensitive to hyper-parameters (e.g., noise scheduling, number of diffusion steps), and need to be customized with optimization strategies for hyperspectral data characteristics. For example, DWSDiff mitigates the background estimation bias by a double-window strategy, but its generalization still needs to be verified.4. Complementarity with traditional methods: traditional methods (e.g., SVM) are still advantageous in terms of interpretability and lightness, while diffusion models are more suitable for high-precision and complex scene tasks. Future research can explore hybrid architectures to balance efficiency and performance.
VII Outlook
In recent years, diffusion models have made significant progress in tasks such as image generation, denoising and reconstruction by virtue of their powerful generative capabilities and modeling advantages over complex data distributions. The diffusion model shows unique potential in the multidimensional and high-information-density field of hyperspectral image processing. Hyperspectral images contain spectral information in hundreds of consecutive bands, but their high-dimensional characteristics and noise interference pose challenges for traditional methods. As shown in the previous chapter, diffusion models can accurately model spectral-spatial features through a step-by-step denoising process, thus achieving higher accuracy in the task. However, the fact is that diffusion models still have a long way to go in the hyperspectral domain, and there are many tasks that still need to be explored and further improved, and the current research is still limited by the consumption of computational resources, insufficient cross-modal fusion and lack of interpretability. In the future, technical innovations can be promoted in the following three aspects: innovation in fusion with other traditional deep learning models, optimization and deployment in real scenarios, and automated model design and hyperparameter tuning.
1, Fusion innovation with other deep learning models, so that the diffusion model and the mainstream architecture to form a complementary, such as CNN + diffusion model, the use of CNN to extract spatial-spectral features as the a priori conditions of the diffusion process, to improve the efficiency of the generation; Transformer + diffusion model, through the attention mechanism to model the long-range dependence, to optimize the super-resolution reconstruction of hyperspectral images; GAN / VAE + diffusion model, which first generates low-resolution samples with GAN or VAE, and then performs detail enhancement and denoising by diffusion model to balance the generation speed and quality; Physics-driven fusion: combining physical a priori such as radiative transfer model, to enhance the generalization ability of diffusion model in complex environments (e.g., atmospheric interference).
2. Optimization and deployment in real-world scenarios need to be achieved through model lightweighting (e.g., knowledge distillation, dynamic diffusion step size), hardware adaptation (parallelized FPGA/GPU-based design) and embedded deployment (low-precision inference on edge devices). For vertical domains such as agricultural disease monitoring and mineral exploration, task-oriented diffusion frameworks need to be developed to balance accuracy and real-time requirements.
3. Automated model design and hyperparameter tuning. The performance of diffusion models is highly dependent on noise scheduling, network architecture and training strategies. Automated techniques (e.g., neural architecture search, meta-learning) can adaptively optimize the model architecture; reinforcement learning-driven hyperparameter search can dynamically adjust the denoising process for different hyperspectral tasks. Meanwhile, the development of unified evaluation benchmarks (e.g., cross-dataset generalization metrics) will accelerate the application and iteration of automated techniques.
In conclusion, diffusion model provides a new paradigm for hyperspectral image processing, but its development still needs to break through the bottlenecks of computational efficiency, cross-modal fusion and domain adaptation. In the future, through the combination of algorithmic innovation, hardware synergy and automation technology, the diffusion model is expected to realize smarter and more efficient analysis systems in the fields of environmental monitoring, precision agriculture, etc., and to promote hyperspectral technology from the laboratory to large-scale applications.
VIII Conclusion
In summary, compared with other deep generative models, the diffusion model is robust to the inherent noise in hyperspectral images, better adapts to its variability and complexity, and provides a more stable training process. Meanwhile, diffusion model with progressive denoising mechanism with high-dimensional data modeling capability, the diffusion model has injected a revolutionary impetus into the field of hyperspectral image analysis. It effectively solves the limitations of traditional methods in high-dimensional data processing, noise interference and small sample scenarios through the Markov chain noise injection and inverse generation process. Therefore, it is an inevitable trend for diffusion model to be applied in the field of hyperspectral images. This paper firstly introduces the related knowledge of hyperspectral as well as the basic theory of diffusion model, and then, reviews the research of diffusion model in the field of hyperspectral images, including: data enhancement, generation and complementation, classification segmentation, anomaly detection, target detection, denoising and other tasks, moreover, this paper fully demonstrates the superiority of diffusion model in this field by comparing some visualization results and quantitative indexes on hyperspectral image processing between diffusion model and traditional model. The limitations of the existing diffusion models in the hyperspectral domain are also discussed, and it is emphasized that further exploratory studies can be carried out for extended applications and model deployment.

### v2

@@
-Diffusion Model in Hyperspectral Image Processing and Analysis: A Review
+
 Abstract
@@
 Traditional models have limitations in dealing with these complex data, and it is difficult to meet the increasing demand for analysis.
-In recent years, Diffusion Model, as an emerging generative model, has shown unique advantages in hyperspectral image processing.
-By simulating the diffusion process of data in time, the Diffusion Model can effectively process high-dimensional data, generate high-quality samples, and perform well in denoising and data enhancement.
+In recent years, Diffusion models, as a class of emerging generative approaches, have demonstrated promising capabilities in hyperspectral image (HSI) processing tasks.
+By simulating the diffusion process of data in time, the Diffusion Model are capable of modeling high-dimensional spectral structures, generate high-quality samples, and achieve competitive performance in spectral-spatial denoising tasks and data enhancement.
 In this paper, we review the recent research advances in diffusion modeling for hyperspectral image processing and analysis, and discuss its applications in tasks such as high-dimensional data processing, noise removal, classification, and anomaly detection.
@@
 It is shown that the diffusion model can significantly improve the accuracy and efficiency of hyperspectral image analysis, providing a new direction for future research.
-Index Terms:
+keywords:
 Hyperspectral image, Diffusion model, High dimensional data processing, Noise removal, Data enhancement, Anomaly detection.
-1School of Optical-Electrical and Computer Engineering, University of Shanghai for Science and Technology, No.
-516, Jungong Road, Shanghai 200093, China (e-mail: huxing@usst.edu.cn, 242250440@st.usst.edu.cn, dwzhang@usst.edu.cn)
-2School of Electronics and Electrical Engineering, Shanghai University of Engineering Science, Shanghai 201620, China (e-mail: dqq1019@163.com)
-3Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing 100094, China (e-mail: hongdf@aircas.ac.cn)
-4School of Electronic, Electrical and Communication Engineering, University of Chinese Academy of Sciences, Beijing 100049, China.
-I Introduction
-Hyperspectral image processing and analysis have important application values in remote sensing, agriculture, environmental monitoring and other fields.
-With the advancement of technology, the resolution and spectral range of hyperspectral images are continuously improved, which provides the possibility of fine feature identification and classification.
-However, the high dimensionality, data redundancy, and noise interference etc.
-of hyperspectral images also bring great challenges to the analysis.
-Traditional processing methods have certain limitations in dealing with these complex data, so the introduction of new models and methods has become the focus of current research.
-In recent years, Diffusion Model (DM), as an emerging generative model, has shown unique advantages in hyperspectral image processing and analysis.
-Hyperspectral imagery provides a wealth of spectral and spatial data by capturing spectral information of objects in multiple narrow bands.
-The high-dimensionality and complexity of this data make it widely used in agriculture, environmental monitoring, mineral exploration and other fields.
-However, hyperspectral image analysis also faces many challenges.
-First, there is a huge demand for high-dimensional data processing and storage with high computational complexity.
-Second, the high correlation between bands leads to data redundancy, which increases the difficulty of analysis.
-In addition, noise interference and mixed image element problems also affect the accuracy of the analysis.
-Modeling using neural networks has been a dominant force over the last decade or so, and has had a huge impact in a variety of domains since the emergence of generative models, ranging from images [1, 2] and audio [3, 4] to text [5] and point clouds [6].
-From a probabilistic modeling perspective, the key defining feature of a generative model is that it is trained in such a way so that its samples $\widetilde{{x}}\sim p_{\theta}(\widetilde{{x}})$ are from the same distribution as the training data distribution ${x}\sim p_{\theta}({x})$.
-Energy-based models (EBM) achieve this by defining non-normalized probability densities on the state space; however, these methods require Markov chain Monte Carlo (MCMC) sampling during both training and inference, which is a slow and iterative process [7] In the last few years, there has been a renewed interest in generative modeling due to the advancement of deep learning architectures in general, which have revealed improved visual fidelity and sampling speed.
-Specifically, Generative Adversarial Networks (GANs) [8], Variational Auto-Encoders (VAEs) [9] and Normalized Streams [10] among others, have emerged.
-In addition to this, generative models based on diffusion processes provide an alternative to existing VAEs, EBMs, GANs, and normalized streams, which do not require aligning a posteriori distributions, estimating tricky partition functions, introducing additional discriminator networks, or placing network constraints, respectively.
-To date, diffusion models have been found useful in a variety of domains, ranging from generative modeling tasks (e.g., image generation [11], image super-resolution [12] image restoration [13]) to discriminative tasks (e.g., image segmentation [14], classification [15], and anomaly detection [16]).
-Recently, the hyperspectral community has witnessed an exponential growth in the number of diffusion-based techniques, and a large number of studies have been dedicated to the application of diffusion models in various hyperspectral scenarios.
-Since diffusion modeling has recently received a great deal of attention from the research community, the literature has a large number of contributions in this direction.
-For this reason, this survey aims to provide a comprehensive overview of the latest advances and a comprehensive overview of such models in the field of hyperspectral spectroscopy.
-Different from this paper [17], firstly, it is aimed at generalized remote sensing images, covering multimodal data such as multispectral, SAR, and panchromatic images, with a wider scope of research but lack of in-depth analysis of hyperspectral properties, and focuses on sorting out the technological paths of traditional tasks such as image generation and super-resolution reconstruction, with the contribution of systematically establishing a methodological mapping of diffusion modeling in remote sensing for the first time.
-Secondly, it focuses more on the panoramic depiction of diffusion modeling applications in remote sensing, and although the conditional diffusion model [11] as well as the latent diffusion model [18] are described, the exploration of the emerging architectures stops at early applications before 2023, lacks a systematic analysis of practical engineering constraints, and does not provide an in-depth discussion of the key technological challenges specific to hyperspectral imagery.
-Instead, this paper focuses on the specific modality of hyperspectral imagery, and designs a multi-perspective classification of diffusion models in hyperspectral domain from the core characteristics of high dimensionality of data and spectral redundancy etc., which provides a systematic taxonomy for the study of diffusion models and their applications.
-Meanwhile, this paper classifies the existing diffusion models into three categories: denoising diffusion probabilistic models (DDPM) [19] and score-based generative models (SGM) [20] and stochastic differential equation-based models (SDE) [21], and analyzes the use of diffusion models in hyperspectral domain with specific classification.
-In addition, we will systematically sort out the progress of the application of diffusion models in vertical tasks, which is more comprehensive than the direction of [17], and categorize the tasks into eight categories: data enhancement and synthesis, data generation and complementation, classification, segmentation, image generation, anomaly detection and noise suppression, and data recovery.
-The significance of this paper lies in the following: first, it fills the gap of systematic review of diffusion model in hyperspectral field; second, it elaborates the methodological innovation of adaptation in hyperspectral field more comprehensively; third, it reveals the unique challenges and optimization paths of diffusion model in hyperspectral scenarios; fourth, it realizes a leap forward breakthrough in the three dimensions of theoretical depth, technological breadth, and application value, and lays down a key theoretical and practical foundation for hyperspectral image processing to move from laboratory research it lays a key theoretical foundation for hyperspectral image processing to move from laboratory research to operational applications.
-The overall structure of this paper is organized as follows: Section I will introduce the definition, characteristics, data properties, and main tasks in the analysis of hyperspectral images; Section II will introduce the theory, evolution, and variants of the Diffusion model, the analysis in comparison with the traditional model, and the application in image generation; Section III will introduce the use of Diffusion model in hyperspectral image analysis and its specific applications; Section IV will introduce the challenges and limitations of Diffusion model; Section V will introduce the comparison between Diffusion model and traditional methods, and conduct experiments and analysis; Section VI will provide a future outlook of Diffusion model in hyperspectral image; Section VII will summarize the content of this paper.
-II Basics of hyperspectral imagery
-II-A Definition and Characteristics of Hyperspectral Imagery
-Hyperspectral image is a spectral image with spectral rate in the range of ${10}^{-2}\lambda$ order of magnitude, covering the range from visible to infrared light range of multiple continuous bands, through the hyperspectral sensors (imaging spectrometer) mounted on different platforms, it can be imaged in the ultraviolet, visible, near-infrared and mid-infrared regions of the electromagnetic spectrum by tens to hundreds of consecutively subdivided spectral bands in the target region, and each pixel point includes information of numerous spectral bands, forming a continuous spectral curve, which can obtain spatial information as well as spectral information at the same time, truly combining spectra and images.
-Hyperspectral images benefit from the development and maturity of imaging spectroscopy technology, which is a sample that combines image information and spectral information, implying a wealth of features that are useful for feature identification and processing.
-Image information is containing external quality information such as size, shape, defects, etc., and spectral information is containing different internal physical structure and chemical composition.
-Both the remote sensing image with multiple segments and the continuous spectral information of each image element can be obtained, realizing the multi-dimensional expression of feature characteristics, and the combination of spatial and spectral information, which is convenient for accurate identification and classification.
-Hyperspectral images have a high spectral resolution, usually at the nanometer level, and are able to finely portray the reflectance spectra of features, differentiate the spectral characteristics of similar substances, and improve the ability to classify and identify features.
-At the same time, it has a large amount of information, redundancy of information, multi-segmented, data including dozens of hundreds of segments, huge amount of data, high requirements for storage and processing capacity, providing rich spectral information, able to capture subtle spectral differences, and correlate between them spatially and spectrally.
-Finally, hyperspectral images are also characterized by fine spectral feature expression, nonlinear characteristics, low signal-to-noise ratio and high dimensionality of data.
-Moreover, hyperspectral images are widely used in many fields, such as agriculture, environmental monitoring, mineral exploration, medicine and so on, and can recognize and classify different substances.
+1 Introduction
+Hyperspectral image processing and analysis are crucial in remote sensing, agriculture, and environmental monitoring, offering significant application value.
+Continuous advancements in technology have led to improved resolution and broader spectral ranges in hyperspectral images, enabling more precise feature identification and classification.
+However, the inherent challenges of high dimensionality, data redundancy, and noise interference in these images often limit traditional processing methods.
+Consequently, the development of novel models and methods has become a key research focus.
+Recently, Diffusion Models (DMs), as an emerging generative model, have demonstrated unique advantages in addressing these complexities within hyperspectral image processing and analysis.
+Hyperspectral imagery, with its ability to capture detailed spectral information across numerous narrow bands, offers a rich source of both spectral and spatial data.
+This high-dimensional and intricate data has found extensive applications in fields like agriculture, environmental monitoring, and mineral exploration.
+However, analyzing hyperspectral images presents several challenges.
+The sheer volume of high-dimensional data demands significant processing and storage capabilities, leading to high computational complexity.
+Moreover, the strong correlation among bands introduces data redundancy, complicating analysis.
+Issues such as noise interference and mixed pixel problems further impact the accuracy of the analysis.
+Neural networks have dominated the past decade, profoundly impacting various domains, including image [1, 2], audio [3, 4], text [5], and point cloud processing [6], particularly with the rise of generative models.
+From a probabilistic standpoint, a key feature of a generative model is its ability to produce samples ($\widetilde{{x}}\sim p_{\theta}(\widetilde{{x}})$) that match the distribution of the training data (${x}\sim p_{\theta}({x})$).
+While Energy-Based Models (EBMs) achieve this by defining non-normalized probability densities, they require slow, iterative Markov Chain Monte Carlo (MCMC) sampling during both training and inference [7].
+Recent advances in deep learning architectures have reignited interest in generative modeling, leading to improved visual fidelity and faster sampling speeds.
+This has seen the emergence of models such as Generative Adversarial Networks (GANs) [8], Variational Auto-Encoders (VAEs) [9], and Normalizing Flows [10].
+In addition to these, Diffusion Models (DMs) offer a compelling alternative to existing VAEs, EBMs, GANs, and Normalizing Flows, as they don’t require aligning posterior distributions, estimating complex partition functions, introducing extra discriminator networks, or imposing network constraints.
+To date, diffusion models have proven useful in a wide range of applications, from generative tasks like image generation [11], image super-resolution [12], and image restoration [13], to discriminative tasks such as image segmentation [14], classification [15], and anomaly detection [16].
+Recently, the hyperspectral community has seen an exponential increase in diffusion-based techniques, with numerous studies dedicated to applying diffusion models in various hyperspectral scenarios.
+Given the significant attention diffusion modeling has garnered from the research community, there’s a large body of literature on this topic.
+Consequently, this survey aims to provide a comprehensive overview of the latest advancements and applications of such models in the field of hyperspectral spectroscopy.
+This paper differentiates itself from [17] in several key aspects.
+First, while [17] provides a broader overview of diffusion models in generalized remote sensing, encompassing multispectral, SAR, and panchromatic data, it lacks an in-depth analysis of the specific properties and challenges of hyperspectral imagery.
+Its focus is more on traditional tasks like image generation and super-resolution, aiming to establish a foundational mapping of diffusion modeling in remote sensing.
+In contrast, this paper specifically targets the hyperspectral modality, offering a more focused and detailed examination.
+Second, although [17] touches upon conditional diffusion models [11] and latent diffusion models [18], its exploration of emerging architectures stops at early applications before 2023.
+It also doesn’t systematically analyze practical engineering constraints or deeply discuss the unique technological challenges inherent in hyperspectral imagery.
+Conversely, this paper provides a multi-perspective classification of diffusion models within the hyperspectral domain, specifically addressing core characteristics like high data dimensionality and spectral redundancy.
+This offers a systematic taxonomy for studying diffusion models and their applications in this field.
+Furthermore, we categorize existing diffusion models into Denoising Diffusion Probabilistic Models (DDPM) [19], Score-Based Generative Models (SGM) [20], and Stochastic Differential Equation-based Models (SDE) [21], analyzing their specific uses in the hyperspectral domain.
+We also systematically outline the progress of diffusion model applications across various tasks, offering a more comprehensive scope than [17].
+These tasks are categorized into eight areas: data enhancement and synthesis, data generation and complementation, classification, segmentation, image generation, anomaly detection, noise suppression, and data recovery.
+The significance of this paper is multifaceted: it fills a gap in the systematic review of diffusion models within the hyperspectral field; it provides a more comprehensive elaboration on the methodological innovations for adapting diffusion models to hyperspectral data; it uncovers unique challenges and optimization pathways for diffusion models in hyperspectral scenarios; and ultimately, it achieves a significant breakthrough across theoretical depth, technological breadth, and application value.
+This work lays a crucial theoretical and practical foundation for transitioning hyperspectral image processing from laboratory research to operational applications.
+The paper is structured as follows: Section I defines hyperspectral images, outlining their characteristics, data properties, and primary analysis tasks.
+Section II delves into the theory, evolution, and variants of Diffusion Models, comparing them with traditional models and exploring their application in image generation.
+Section III focuses on the specific use of Diffusion Models in hyperspectral image analysis, detailing various applications.
+Section IV addresses the challenges and limitations inherent in Diffusion Models.
+Section V presents a comparative analysis between Diffusion Models and traditional methods, supported by experimental results.
+Section VI offers a future outlook on the role of Diffusion Models in hyperspectral imaging, and Section VII concludes the paper with a summary of its content.
+2 Basics of hyperspectral imagery
+2.1 Definition and Characteristics of Hyperspectral Imagery
+A hyperspectral image is a spectral image characterized by a spectral resolution on the order of ${10}^{-2}\lambda$, encompassing tens to hundreds of continuously subdivided spectral bands across the visible to infrared light range.
+Captured by hyperspectral sensors (imaging spectrometers) mounted on various platforms, these images can cover the ultraviolet, visible, near-infrared, and mid-infrared regions of the electromagnetic spectrum within a target area.
+Crucially, each pixel in a hyperspectral image contains information from numerous spectral bands, forming a continuous spectral curve.
+This unique capability allows for the simultaneous acquisition of both spatial and spectral information, effectively integrating spectra with images.
+Thanks to the maturation of imaging spectroscopy technology, hyperspectral images offer a unique fusion of image and spectral information.
+This integration provides a rich array of features, proving highly beneficial for both feature identification and processing tasks.
+Image information captures external qualities like size, shape, and defects, while spectral information reveals internal physical structure and chemical composition.
+This dual capability allows for the acquisition of both multi-segment remote sensing imagery and continuous spectral data for each image element.
+This enables a multi-dimensional representation of feature characteristics, combining spatial and spectral information for more accurate identification and classification.
+Hyperspectral images boast high spectral resolution, typically at the nanometer level.
+This allows for a detailed portrayal of feature reflectance spectra, enabling the differentiation of spectrally similar substances and significantly enhancing the accuracy of feature classification and identification.
+Simultaneously, these images contain an enormous amount of data, often comprising dozens to hundreds of bands, leading to information redundancy and multi-segmentation.
+This vast data volume necessitates high demands for storage and processing capacity.
+Despite these challenges, hyperspectral images provide exceptionally rich spectral information, capable of capturing subtle spectral differences and revealing correlations across both spatial and spectral dimensions.
+Finally, hyperspectral images are characterized by their fine spectral feature expression, nonlinear characteristics, low signal-to-noise ratio, and high data dimensionality.
+These images are broadly applied across numerous fields, including agriculture, environmental monitoring, mineral exploration, and medicine, enabling the recognition and classification of diverse substances.
 Figure 1: Hyperspectral image (a) and Illustration of hyperspectral image (b)
-II-B High-dimensional characterization of hyperspectral image data
-The data of hyperspectral images are characterized by high dimensionality, which is mainly reflected in the multidimensional structure, data redundancy, computational complexity, and information richness.
-Hyperspectral image is a three-dimensional data cube, containing spatial, radiometric and spectral dimensions, with a large number of bands, consecutive bands forming high-dimensional vectors, and covering a range of bands with narrow band widths and high resolution at the nanometer level, which finely reflects the physicochemical properties, while adjacent bands are highly similar, resulting in redundancy of information and increasing the complexity of the data.
-Hyperspectral image is a combination of spatial and spectral, forming a “map in one”, which makes the dimension of the information itself is higher, when analyzing the data, we have to consider the spatial up and down and spectral characteristics, increasing the complexity of the model.
-High-dimensional data requires a large amount of computational resources when processing, especially when performing tasks such as feature extraction, classification, segmentation, etc., the computational complexity increases significantly.
-The storage requirements are high due to the huge amount of data, which is in high demand and costly.
-Traditional machine learning methods are slow to train in high-dimensional space and are susceptible to noise interference, which is especially obvious in high dimensions and increases the difficulty of processing; although it can be automatically extracted under deep learning, a large number of markers are required and the cost is extremely high.
-The high dimensional nature of the data creates excessive waste of resources for the processing, which can easily lead to dimensional disaster and inverted performance degradation.
-In this regard there is a great need for the use of dimensionality reduction methods, such as PCA, LDA, etc.
-to reduce the dimensionality, retain important information and reduce the complexity of the model.
-Under the complexity process, it is important to ensure that problems such as overfitting in training occur.
-II-C Main tasks in hyperspectral image analysis
-The main tasks of hyperspectral image analysis include classification, segmentation, and anomaly detection etc.
-These tasks have important application value in hyperspectral image processing and can help us to extract useful information from complex high-dimensional data.
-The following is a detailed description of these tasks:
-Classification: Hyperspectral image classification refers to the assignment of each pixel or region in an image to a specific class.
-These categories can be different types of features, such as vegetation, water bodies, buildings, and so on.
-It is often applied to crop classification and pest detection in agriculture; land use classification and forest cover monitoring in environmental monitoring; and mineral type identification in mineral exploration.
-Often exists high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of classification; data labeling: the classification task usually requires a large amount of labeled data, and the labeling cost of hyperspectral images is high; mixed image element problem: due to the limitation of spatial resolution, one pixel may contain multiple features, which increases the difficulty of classification and other problems.
-The traditional methods are statistical-based methods (e.g., maximum likelihood classification), machine learning-based methods (e.g., support vector machine SVM).
-Deep learning methods are convolutional neural networks (CNN), recurrent neural networks (RNN), etc.
-Segmentation: Hyperspectral image segmentation refers to dividing an image into multiple regions or objects, each with similar spectral characteristics.
-Segmentation can be pixel-level segmentation or object-level segmentation.
-It is often applied to the segmentation of tumor regions in medical image analysis, the segmentation of buildings and roads in urban planning, and the segmentation of crop regions in agriculture.
-Often exist high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of segmentation; noise interference: noise in hyperspectral images affects the accuracy of segmentation; mixed image element problem: a pixel may contain multiple features, which increases the difficulty of segmentation and other problems.
-The traditional methods are clustering based methods (e.g.
-Kmeans), edge detection based methods.
-Deep learning methods are Full Convolutional Networks (FCN), UNet, etc.
-Target Detection: Hyperspectral image target detection refers to locating and identifying specific targets or objects in an image.
-These targets can be military targets, mineral resources, vegetation types and so on.
-Often used in military target identification and tracking; mineral exploration in the detection of specific minerals; environmental monitoring in the detection of specific vegetation or pollution sources.
-Often exist high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of target detection; target diversity: the spectral characteristics of different targets may be similar, increasing the difficulty of detection; background interference: complex background will affect the accuracy of target detection and other problems.
-Traditional methods are based on spectral matching methods, statistical methods.
-Deep learning methods are Regional Convolutional Neural Networks (RCNN), YOLO (You Only Look Once) and so on.
-Anomaly Detection: Hyperspectral image anomaly detection refers to the detection of pixels or regions in an image that are significantly different from their surroundings.
-These anomalies may be unknown targets or features.
-It is often applied to the detection of unknown targets in the military, the detection of pollution sources in environmental monitoring, and the early detection of pests and diseases in agriculture.
-Often exist high-dimensional data processing: the high-dimensional nature of hyperspectral images increases the complexity of anomaly detection; diversity of anomalies: different types of anomalies may have different spectral characteristics, increasing the difficulty of detection; background interference: complex background can affect the accuracy of anomaly detection and other problems.
-The traditional methods are statistical based methods (e.g.
-RX algorithm), clustering based methods.
-Deep learning methods are Autoencoder, Generative Adversarial Networks (GAN), and so on.
-Noise suppression: Hyperspectral noise suppression refers to the reduction or elimination of noise in hyperspectral images by means of specific algorithms and techniques, in order to improve the quality of the image and the accuracy of the subsequent analysis.
-It is mainly used in agriculture, military, food industry, mining exploration and other fields.
-Commonly used methods are mainly frequency or transform domain frequency reduction, spatial frequency reduction, and deep learning frequency reduction, and so on.
-Data recovery: Hyperspectral data recovery refers to the removal of noise and interference in hyperspectral images through specific algorithms and technical means to restore their original spectral and spatial information to improve the quality and usability of data, which belongs to the further implementation of noise suppression.
-The main application areas are agriculture, environmental monitoring, biomedicine, geological exploration and so on.
-The commonly used methods are spectral reconstruction, spectral dimension recovery, low-rank representation modeling, and deep learning methods, and so on.
-Data Augmentation and Synthesis: Hyperspectral image data augmentation and synthesis refers to increasing the diversity and quantity of training data by generating new hyperspectral image samples, so as to improve the generalization ability of the model.
-It is often applied in model training to increase the diversity of training data and improve the generalization ability of the model; and in data completion to generate missing hyperspectral image data.
-Often exist high-dimensional data processing: the high-dimensional nature of hyperspectral images increases the complexity of data enhancement and synthesis; data quality: the generated samples need to be of high quality and diversity to ensure the validity of the model and other issues.
-The traditional methods are interpolation-based methods, statistics-based methods.
-Deep learning methods are Generative Adversarial Networks (GAN), Diffusion Model, and so on.
-Data generation and completion: Hyperspectral data generation and completion is a core task in data science, aiming to solve the problems of missing data, noise interference and data scarcity, and at the same time extend the application scenarios through synthetic data.
-Data generation generates synthetic data that conforms to the real data distribution through algorithms, which is used to enhance the data scale or simulate specific scenarios; data completion fills in the missing values or repairs incomplete data to maintain the consistency and completeness of the data distribution.
-It is usually used in remote sensing, medical imaging, agricultural monitoring and other fields.
-The traditional methods are based on statistics (interpolation), machine learning (K-nearest neighbor, decision tree) deep learning methods such as GAN, VAE, Diffusion model, etc.
-In summary, the main tasks of hyperspectral image analysis include classification, segmentation, target detection, anomaly detection and data enhancement and synthesis.
-These tasks have important applications in hyperspectral image processing and can help us extract useful information from complex high-dimensional data.
-Although these tasks face challenges such as high-dimensional data processing, noise interference, and mixed image element problems, the accuracy and efficiency of hyperspectral image analysis can be significantly improved by combining traditional and deep learning methods.
-III Diffusion Model
-Diffusion Model is an emerging generative model that has demonstrated powerful capabilities in the fields of image generation, denoising, and data enhancement in recent years.
-It gradually transforms noisy data into target data by modeling the diffusion process of data in time.
-The core idea of the diffusion model is to gradually transform the data into noise through the forward diffusion process, and then recover the original data from the noise through the reverse diffusion process.
-This approach performs well in high-dimensional data processing, image generation and denoising tasks, and especially has unique advantages in hyperspectral image processing.
-III-A Theoretical Foundations of the Diffusion Model
-Diffusion model is divided into two processes: forward and backward diffusion diffusion; and finally training objectives are performed to minimize the difference between forward and backward diffusion processes.
-The basic diffusion models are usually divided into (1) denoising diffusion probabilistic model (DDPM), (2) score-based generative model (SGM), and stochastic differential equation-based model (SDE/NCSN).
-The basic principles of these three diffusion models are described below.
-III-A1 Denoising Diffusion Probabilistic Model (DDPM)
-Denoising Diffusion Probabilistic Model (DDPM) [19, 22] utilize two Markov chains: a forward chain that perturbs the data into noise, and an inverse chain that converts the noise back into data.
-The former is usually manually designed to convert any data distribution into a simple prior distribution (e.g., a standard Gaussian distribution), while the latter Markov chain inverts the former by learning a transition kernel parameterized by a deep neural network.
-New data points are generated by first randomly sampling a vector from the previous distribution and ancestrally sampling it via an inverse Markov chain [23].
-Figure 2: Diffusion processes in diffusion models
-Formally, given a data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$,, a forward Markov process generates a sequence of random variables $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ replaced by a transition kernel $q(x_{t}\mid{x_{t-1}})$.
-Using the probability chain rule and the Markov property, the joint distribution of $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ is later decomposed, conditional on $\mathbf{x}_{0}$, and denoted as $q(x_{1},\ldots{,x_{T}}\mid{x_{0}})$, which becomes:
+2.2 High-dimensional characterization of hyperspectral image data
+Hyperspectral image data is inherently high-dimensional, primarily due to its multidimensional structure, significant data redundancy, high computational complexity, and exceptional information richness.
+A hyperspectral image functions as a three-dimensional data cube, encompassing spatial, radiometric, and spectral dimensions.
+It features a large number of contiguous bands, forming high-dimensional vectors.
+The narrow bandwidths and nanometer-level resolution of these bands precisely reflect physicochemical properties.
+However, the high similarity between adjacent bands leads to information redundancy, increasing data complexity.
+The "map in one" nature of hyperspectral images, combining spatial and spectral information, inherently elevates the data’s dimensionality.
+This necessitates considering both spatial and spectral characteristics during analysis, thereby increasing model complexity.
+Processing such high-dimensional data, especially for tasks like feature extraction, classification, and segmentation, demands substantial computational resources, leading to significantly increased computational complexity.
+Storage requirements are also high due to the immense data volume, which in turn incurs high costs.
+Traditional machine learning methods struggle with slow training in high-dimensional spaces and are susceptible to noise interference, a problem exacerbated in high dimensions, making processing more difficult.
+While deep learning can automate feature extraction, it often requires a large number of costly labeled samples.
+The high dimensionality of hyperspectral data can lead to excessive resource waste, often resulting in the "curse of dimensionality" and diminished performance.
+To counter this, dimensionality reduction techniques like PCA and LDA are crucial for preserving essential information while simplifying the model.
+During this complex process, it is vital to mitigate issues such as overfitting during training.
+2.3 Main tasks in hyperspectral image analysis
+The primary tasks in hyperspectral image analysis, including classification, segmentation, and anomaly detection, hold significant value in processing these complex, high-dimensional datasets.
+They are essential for extracting meaningful and useful information.
+Let’s delve into each of these tasks:
+Classification in hyperspectral imagery involves assigning each pixel or region to a predefined class, such as vegetation, water bodies, or buildings.
+This is widely applied in agriculture for crop classification and pest detection, in environmental monitoring for land use classification and forest cover monitoring, and in mineral exploration for identifying mineral types.
+However, several challenges persist: the high-dimensional nature of hyperspectral data increases classification complexity; data labeling is costly due to the extensive amount of labeled data required; and the mixed pixel problem, where a single pixel encompasses multiple features due to spatial resolution limitations, complicates accurate classification.
+Traditional approaches include statistical methods (e.g., maximum likelihood classification) and machine learning methods (e.g., Support Vector Machines, SVM).
+More recently, deep learning methods such as Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have also been employed.
+Segmentation in hyperspectral imagery involves partitioning an image into distinct regions or objects that share similar spectral characteristics.
+This can be performed at either the pixel or object level.
+Common applications include segmenting tumor regions in medical image analysis, identifying buildings and roads in urban planning, and delineating crop areas in agriculture.
+However, several challenges arise: the high dimensionality of hyperspectral data complicates the segmentation process; noise interference can compromise accuracy; and the mixed pixel problem, where a single pixel represents multiple features, makes precise segmentation difficult.
+Traditional approaches often rely on clustering-based methods (e.g., K-means) or edge detection techniques.
+Deep learning methods, such as Fully Convolutional Networks (FCNs) and UNet, are increasingly employed for this task.
+Target detection in hyperspectral imagery involves identifying and locating specific targets or objects, such as military assets, mineral resources, or vegetation types, within an image.
+Its applications are diverse, ranging from military target identification and tracking, to the detection of specific minerals in exploration, and the identification of particular vegetation or pollution sources in environmental monitoring.
+This task presents several challenges: the high dimensionality of hyperspectral data increases the complexity of detection; target diversity, where different targets may exhibit similar spectral characteristics, can hinder accurate detection; and background interference from complex environments can significantly impact detection accuracy.
+Traditional methods for target detection include spectral matching techniques and statistical methods.
+More recently, deep learning approaches like Region-based Convolutional Neural Networks (RCNN) and YOLO (You Only Look Once) have been widely adopted.
+Anomaly detection in hyperspectral imagery identifies pixels or regions that significantly deviate from their surrounding environment.
+These anomalies can represent unknown targets or unique features.
+This technique finds applications in military contexts for detecting unidentified targets, in environmental monitoring for pinpointing pollution sources, and in agriculture for the early detection of pests and diseases.
+However, challenges arise due to the high dimensionality of hyperspectral images, which increases the complexity of anomaly detection.
+Furthermore, the diversity of anomalies, where different types may exhibit varying spectral characteristics, complicates their detection.
+Lastly, background interference from complex environments can negatively impact detection accuracy.
+Traditional approaches include statistical methods (e.g., the RX algorithm) and clustering-based methods.
+More recently, deep learning methods such as Autoencoders and Generative Adversarial Networks (GANs) have been employed.
+Noise suppression in hyperspectral imagery aims to reduce or eliminate unwanted noise through specialized algorithms and techniques.
+This process is crucial for enhancing image quality and ensuring the accuracy of subsequent analyses.
+Its applications span diverse fields, including agriculture, military operations, the food industry, and mining exploration.
+Common methods for noise suppression primarily involve frequency or transform domain noise reduction, spatial noise reduction, and deep learning-based noise reduction.
+Hyperspectral data recovery is a more advanced form of noise suppression, focused on removing noise and interference from hyperspectral images to restore their original spectral and spatial information.
+This process significantly enhances data quality and usability.
+Key application areas include agriculture, environmental monitoring, biomedicine, and geological exploration.
+Common methods for data recovery include spectral reconstruction, spectral dimension recovery, low-rank representation modeling, and various deep learning approaches.
+Hyperspectral image data augmentation and synthesis involves generating new hyperspectral image samples to increase the diversity and quantity of training data, thereby enhancing a model’s generalization capability.
+This is frequently applied in model training to diversify data and improve model robustness, and in data completion to generate missing hyperspectral image data.Challenges in this area include: the high dimensionality of hyperspectral images, which complicates data augmentation and synthesis; and ensuring the quality and diversity of generated samples to maintain model validity.
+Traditional methods often rely on interpolation-based techniques and statistics-based methods.
+More recently, deep learning approaches like Generative Adversarial Networks (GANs) and Diffusion Models have shown promise.
+Hyperspectral data generation and completion is a fundamental task in data science that addresses issues like missing data, noise interference, and data scarcity, while also expanding application possibilities through synthetic data.
+Data generation involves creating synthetic data that aligns with real data distributions, often to augment dataset scale or simulate specific scenarios.
+Data completion, on the other hand, focuses on filling in missing values or repairing incomplete data to ensure consistency and completeness.
+These techniques are commonly employed in fields such as remote sensing, medical imaging, and agricultural monitoring.
+Traditional methods include statistics-based approaches (e.g., interpolation) and machine learning techniques (e.g., K-nearest neighbor, decision trees).
+More recently, deep learning methods like GANs, VAEs, and Diffusion Models have become prominent for these tasks.
+In summary, the core tasks of hyperspectral image analysis encompass classification, segmentation, target detection, anomaly detection, and data enhancement and synthesis.
+These tasks are vital in processing hyperspectral imagery, enabling the extraction of valuable information from complex, high-dimensional datasets.
+While these tasks contend with challenges such as high-dimensional data processing, noise interference, and mixed pixel issues, combining traditional and deep learning methodologies can significantly enhance the accuracy and efficiency of hyperspectral image analysis.
+3 Diffusion Model
+The Diffusion Model (DM) is an innovative generative model that has recently showcased remarkable capabilities in image generation, denoising, and data enhancement.
+Its core principle involves modeling a data diffusion process over time, progressively transforming noisy data into desired target data.
+Specifically, DMs operate by gradually converting original data into noise through a forward diffusion process, and then meticulously reconstructing the original data from this noise via a reverse diffusion process.
+This methodology proves highly effective for handling high-dimensional data, as well as for image generation and denoising tasks, demonstrating particular advantages in the realm of hyperspectral image processing.
+3.1 Theoretical Foundations of the Diffusion Model
+The Diffusion Model operates through two distinct processes: forward diffusion and backward diffusion.
+The training objective is to minimize the discrepancy between these two processes.
+Fundamentally, diffusion models are typically categorized into three main types: (1) Denoising Diffusion Probabilistic Models (DDPM), (2) Score-Based Generative Models (SGM), and (3) Stochastic Differential Equation-based Models (SDE/NCSN).
+The foundational principles of each of these three diffusion models are described below.
+3.1.1 Denoising Diffusion Probabilistic Model (DDPM)
+Denoising Diffusion Probabilistic Models (DDPMs) [19, 22] employ two Markov chains.
+A forward chain systematically introduces noise to transform the data into a simple prior distribution, typically a standard Gaussian.
+Conversely, an inverse chain learns to reverse this process, converting the noise back into data through a transition kernel parameterized by a deep neural network.
+New data points are generated by first randomly sampling a vector from the prior distribution, then using an inverse Markov chain to ancestrally sample it [23].
+Figure 2: Diffusion processes in diffusion models [19]
+Formally, given a data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$, a forward Markov process generates a sequence of random variables, $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}$, through a transition kernel $q(x_{t}\mid{x_{t-1}})$.
+By applying the probability chain rule and the Markov property, the joint distribution of these variables, conditional on $\mathbf{x}_{0}$, is then decomposed and denoted as $q(x_{1},\ldots,x_{T}\mid{x_{0}})$, which results in:
 $q(\mathbf{x}_{1},...,\mathbf{x}_{T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$ (1)
-In DDPM, the transition kernel $q(x_{t}\mid{x_{t-1}})$ is drawn manually to incrementally transform the data distribution $q(\mathbf{x}_{0})$ into a processable prior distribution.
-A typical design of a transition kernel is a Gaussian perturbation, and the most common choice of transition kernel is:
+In DDPMs, the transition kernel $q(x_{t}\mid{x_{t-1}})$ is specifically designed to gradually transform the data distribution $q(\mathbf{x}_{0})$ into a manageable prior distribution.
+A common and effective choice for this transition kernel is a Gaussian perturbation, typically defined as:
 $q(\mathrm{x}_{t}\mid\mathrm{x}_{t-1})=N(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\mathrm{I}).$ (2)
-where $\beta_{t}\in(0,1)$ is the hyperparameter chosen prior to model training.
-This kernel is applied to simplify the discussion, although other types of kernels apply in the same vein.
-As observed by Sohl-Dickstein et al.
-2015 [22] this Gaussian transition kernel allows us to marginalize the joint distribution in equation (1) to obtain an analytic form of $q(x_{t}\mid\mathbf{x}_{0})$ for all $t\in{0,1,\cdots,T}$.
-$\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}{:}=\prod_{s=0}^{t}\alpha_{s}$, in particular:
+where $\beta_{t}\in(0,1)$ is a hyperparameter set before model training.
+While other kernel types exist, this Gaussian transition kernel is used to simplify the discussion.
+As noted by Sohl-Dickstein et al.
+(2015) [22], this specific kernel allows for the marginalization of the joint distribution in equation (1), yielding an analytical form for $q(x_{t}\mid\mathbf{x}_{0})$ across all $t\in\{0,1,\cdots,T\}$.
+Specifically, with $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}{:}=\prod_{s=0}^{t}\alpha_{s}$:
 $q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right).$ (3)
-Intuitively, this forward process slowly injects noise into the data until all structure is lost.
-To generate new data samples, the DDPM first generates an unstructured noise vector from the prior distribution (which is usually easy to obtain), and then gradually removes the noise from it by running a learnable Markov chain in the opposite time direction.
-Specifically, the inverse Markov chain is parameterized by the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ and the learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$.
-We choose the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ because the forward process is constructed such that $q(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$.
+Intuitively, this forward process gradually introduces noise into the data until its original structure is completely lost.
+To generate new data samples, the DDPM first samples an unstructured noise vector from the prior distribution (which is typically straightforward to obtain).
+It then progressively removes this noise by running a learnable Markov chain in the reverse temporal direction.
+Specifically, this inverse Markov chain is parameterized by the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ and the learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$.
+We choose $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ for the prior because the forward process is designed such that $q(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$.
 The learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ employs:
 $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\right).$ (4)
-where $\theta$ denotes the model parameters and the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized by the deep neural network.
-The key to the success of the network training goal is to train the reverse process $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$ to match the actual time reversal of the forward process $q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$.
-This is usually achieved by minimizing the KL scatter (Kullback-Leibler) between the two:
+where $\theta$ represents the model’s parameters, while the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized by a deep neural network.
+The key to successfully training the network lies in ensuring the reverse process $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$ accurately matches the true time reversal of the forward process $q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$.
+This is typically achieved by minimizing the Kullback-Leibler (KL) divergence between the two:
 $\mathrm{KL}\left(q(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})|\||p_{\theta}(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})\right)$ (5)
@@
 $\overset{(iii)}{\operatorname*{\operatorname*{\geq}}}\mathbb{E}[-\log p_{\theta}\left(\mathbf{x}_{0}\right)]+\mathrm{const}$ (8)
-where “const” is a constant that does not depend on the model parameters $\theta$.
-The goal of DDPM training is to maximize the variational lower bound or equivalently minimize the negative variational lower bound, and thus it can be efficiently estimated by Monte Carlo sampling [24] and efficiently optimized by stochastic optimization [25].
-III-A2 Score-Based Generative Models (SGM/NCSN)
-At the heart of Score-Based Generative Models (SGM/NCSN) [26, 20, 27] is the notion of a (Stein) score (aka fraction or score function) [28].
-Given a probability density function p(x), the score function is defined as the gradient of the logarithmic probability density $\nabla_{\mathbf{x}}\log p(\mathbf{x})$.
-Unlike the Fisher score $\nabla_{\theta}\log p_{\theta}(\mathbf{x})$, which is commonly used in statistics, the Stein score considered here is a function of the data x and not of the model parameters $\theta$.
-It is a vector field pointing in the direction where the probability density function has the largest growth rate.
-The key idea of Score-Based Generative Models (SGM) [20] is to perturb the data with a series of augmented Gaussian noises and jointly estimate the score function for all noisy data distributions by training a deep neural network model (called Noise Conditioned Scoring Network, NCSN, in [20]) conditional on the noise level.
-The samples are generated by linking the score functions with decreasing noise levels by using score-based sampling methods including Ronzmann Monte Carlo [29, 30, 31, 20, 21], stochastic differential equations [32, 21]], ordinary differential equations [33, 34, 35, 36, 21], and various combinations thereof [21].
-In the formulation of the score-based generative model, training and sampling are completely decoupled, so that multiple sampling techniques can be used after estimating the score function.
+where "const" is a constant independent of the model parameters $\theta$.
+The objective of DDPM training is to maximize the variational lower bound, or equivalently, to minimize the negative variational lower bound.
+This can be efficiently estimated through Monte Carlo sampling [24] and optimized effectively via stochastic optimization [25].
+3.1.2 Score-Based Generative Models (SGM/NCSN)
+At the core of Score-Based Generative Models (SGM/NCSN) [26, 20, 27] lies the concept of a (Stein) score, also known as a fraction or score function [28].
+For a given probability density function $p(\mathbf{x})$, the score function is defined as the gradient of the logarithmic probability density: $\nabla_{\mathbf{x}}\log p(\mathbf{x})$.
+Unlike the Fisher score, $\nabla_{\theta}\log p_{\theta}(\mathbf{x})$, which is commonly used in statistics and depends on the model parameters $\theta$, the Stein score is a function of the data $\mathbf{x}$ itself.
+Essentially, it’s a vector field that points in the direction of the steepest ascent of the probability density function.
+The central idea behind Score-Based Generative Models (SGMs) [20] is to progressively add Gaussian noise to data and then collectively estimate the score function for all these noisy data distributions.
+This estimation is achieved by training a deep neural network, specifically known as a Noise Conditioned Scoring Network (NCSN) [20], which is conditioned on the noise level.
+Samples are then generated by linking these score functions across decreasing noise levels.
+This is done using various score-based sampling methods, including Ronzmann Monte Carlo [29, 30, 31, 20, 21], stochastic differential equations [32, 21], ordinary differential equations [33, 34, 35, 36, 21], and their combinations [21].
+A key advantage of the SGM formulation is the complete decoupling of training and sampling, which allows for the use of multiple sampling techniques once the score function has been estimated.
 […]
@@
 $\begin{split}(iii)\\ =&\mathbb{E}_{t\sim\mathcal{U}[[1,T]],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I})}[\lambda(t)\|\epsilon+\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\|^{2}]\\ +\mathrm{const}.\end{split}$ (13)
-where $\lambda(t)$ is a positive weighting function, $\mathbf{x_{t}}$ is computed from $\varepsilon$ and $\mathbf{x}_{0}$, $\mathcal{U}[[1,T]]$ is a uniform distribution on the set ${1,2,\cdots,T}$, and $\varepsilon$ is a deep neural network with parameter $\theta$ for predicting the given $\mathbf{x}_{t}$ and t noise vectors.
-Comparing the equations, the training objectives of the DDPM and SGM are equivalent once $\epsilon_{\theta}(\mathbf{x},t)=-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x},t)$ is set.
-III-A3 Diffusion model based on stochastic differential equations (Score-SDE)
-When DDPM and SGM can be further generalized to the case of infinite time step or noise level, where the perturbation and denoising process is the solution of a stochastic differential equation (SDE), a formulation known as Score SDE [21], which is utilized for the noise perturbation and sample generation, and the denoising process requires estimation of a score function of the distribution of noisy data.
-The SDE was used to score the data using a diffusion process controlled by the following stochastic differential equation (SDE) [21]:
+where $\lambda(t)$ represents a positive weighting function, while $\mathbf{x_{t}}$ is derived from $\varepsilon$ and $\mathbf{x}_{0}$.
+$\mathcal{U}[[1,T]]$ denotes a uniform distribution over the set $\{1,2,\ldots,T\}$, and $\varepsilon$ is a deep neural network with parameters $\theta$, tasked with predicting the noise vectors for a given $\mathbf{x}_{t}$ and $t$.
+Comparing the equations, the training objectives for DDPM and SGM become equivalent once $\epsilon_{\theta}(\mathbf{x},t)=-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x},t)$ is established.
+3.1.3 Diffusion model based on stochastic differential equations (Score-SDE)
+Both DDPM and SGM can be further generalized to an infinite time step or noise level.
+In this advanced formulation, known as Score SDE [21], the perturbation and denoising processes are governed by the solution of a stochastic differential equation (SDE).
+This approach is employed for both noise perturbation and sample generation, where the denoising process necessitates estimating a score function of the noisy data distribution.
+The Stochastic Differential Equation (SDE) was employed to score the data using a diffusion process, which is governed by the following SDE [21]:
 $\mathrm{dx}=\mathrm{f}(\mathrm{x},t)\mathrm{d}t+g(t)\mathrm{dw}.$ (14)
-where f(x,t) and g(t) are the diffusion and drift functions of the SDE, and w is a standard Wiener process (aka Brownian motion) The forward processes in both the DDPM and the SGM are discretizations of this SDE.
-2020 As demonstrated in the literature [21] by Song et al.
-the corresponding SDE for the DDPM is: $dx=-\frac{1}{2}\beta(t)xdt+\sqrt{\beta(t)}dw$.
-For the SGM the corresponding SDE is: $\mathrm{dx}=\sqrt{\frac{\mathrm{d}[\sigma(t)^{2}]}{\mathrm{d}t}}\mathrm{dw}$.
-where, $e\beta(\frac{t}{T})=T\beta_{t}$, $\sigma(\frac{t}{T})=\sigma_{t}$ when T tends to infinity.
-Here $q_{t}({x})$ is used to denote the distribution of $\mathbf{x}_{t}$ in the forward process.
-For any diffusion process of the form of equation (14), Anderson proposed [42] to show that it can be inverted by solving the following reverse time SDE:
+where $f(\mathbf{x},t)$ and $g(t)$ represent the drift and diffusion functions of the SDE, respectively, while $w$ denotes a standard Wiener process (also known as Brownian motion).
+The forward processes in both the DDPM and SGM are essentially discretizations of this SDE.
+As demonstrated by Song et al.
+in 2020 [21], the corresponding SDE for the DDPM is given by $dx=-\frac{1}{2}\beta(t)xdt+\sqrt{\beta(t)}dw$.
+For the SGM, the corresponding SDE is $\mathrm{dx}=\sqrt{\frac{\mathrm{d}[\sigma(t)^{2}]}{\mathrm{d}t}}\mathrm{dw}$.
+This holds true as $T$ approaches infinity, where $e\beta(\frac{t}{T})=\beta_{t}T$ and $\sigma(\frac{t}{T})=\sigma_{t}$.
+Here, $q_{t}({x})$ represents the distribution of $\mathbf{x}_{t}$ in the forward process.
+For any diffusion process structured like equation (14), Anderson [42] demonstrated that it can be inverted by solving the following reverse-time SDE:
 $\mathrm{d}\mathbf{x}=[\mathrm{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})]\mathrm{d}t+g(t)\mathrm{d}\overline{\mathbf{w}}.$ (15)
-where $\bar{\mathbf{W}}$ is the standard Wiener process when time flows backwards and dt denotes an infinitesimal negative time step.
-The solution trajectories of this reverse SDE have the same marginal density as those of the forward SDE, except that they evolve along opposite time directions [21].
-Intuitively, the solution of a reverse-time SDE is a diffusion process that gradually converts noise into data.
-In addition, the literature presented by Song et al.
-in 2020 [21] proved the existence of an ordinary differential equation (ODE), the probabilistic flow ODE, whose trajectory has the same margin as the reverse time SDE.
-The probabilistic flow ODE is given by the following equation:
+where $\bar{\mathbf{W}}$ represents the standard Wiener process as time flows backward, and $dt$ signifies an infinitesimal negative time step.
+The solution trajectories of this reverse SDE share the same marginal density as those of the forward SDE, but evolve in the opposite temporal direction [21].
+Conceptually, the solution of a reverse-time SDE describes a diffusion process that progressively transforms noise into data.
+Furthermore, Song et al.
+(2020) [21] demonstrated the existence of an ordinary differential equation (ODE), known as the probability flow ODE, whose trajectory exhibits the same marginals as the reverse-time SDE.
+The probability flow ODE is defined by the following equation:
 $\mathrm{d}\mathbf{x}=\left[\mathrm{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})\right]\mathrm{d}t.$ (16)
-Both the inverse time SDE and the probabilistic flow ODE allow sampling from the same data distribution because their trajectories have the same margins.
-If one knows the fractional function $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ for each time step t, one can derive its inverse-time equations, which can then be used to generate samples by numerical techniques such as annealed Ronzwan dynamics [20], numerical SDE solvers [32, 21], numerical ODE solvers [33, 34, 21, 36, 43]], and predictor-corrector methods (a combination of MCMC and numerical ODE/SDE solver combination) [21].
-As with the SGM, the transient fraction model $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is parameterized and the fraction function is estimated by generalizing the fraction matching objective in Equation (13) to continuous time, resulting in the following objective:
+Both the inverse time SDE and the probabilistic flow ODE allow for sampling from the same data distribution because their trajectories share identical marginals.
+If the fractional function $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ is known for each time step $t$, its inverse-time equations can be derived.
+These equations can then be used to generate samples via numerical techniques such as annealed Ronzwan dynamics [20], numerical SDE solvers [32, 21], numerical ODE solvers [33, 34, 21, 36, 43], and predictor-corrector methods (a combination of MCMC and numerical ODE/SDE solvers) [21].
+Similar to the SGM, the transient fractional model $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is parameterized, and the fractional function is estimated by generalizing the fractional matching objective in Equation (13) to continuous time, resulting in the following objective:
 $\begin{split}\mathbb{E}_{t\sim\mathcal{U}[0,T],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{0t}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)\|^{2}\right].\end{split}$ (17)
-where $\mathcal{U}[0,T]$ denotes the uniform distribution on $\mathbf{[}0,T]$.
-The above is an introduction to the principles and formulas of the basic diffusion model, through their theoretical support, which makes the diffusion model work in the modern image field, solving certain limitations of the traditional model, but also brings defects.
-The study of Denoising Diffusion Probabilistic Models (DDPM) is the main focus in targeting the hyperspectral field, and it is also the most applied diffusion model in this field, the other two types of applications are relatively less, and they will also be introduced in the following sections.
-III-B Evolution of the Diffusion model with different variants
-Diffusion model started in 2015, Sohl-Dickstein et al.
-first proposed the theoretical framework of diffusion model, which combines nonequilibrium thermodynamics with generative model for the first time [22].In 2019, Song Yang et al.
-proposed Score-Based Generative Models (SGM/NCSN) [20], combining Score-Based Generative Models with Score Matching,which is a score-based generative model that generates by estimating the logarithmic gradient of the data distribution.
-Immediately following the optimization of the model in 2020, the team proposed NCSN++ (Noise-Conditioned Score Networks++) [27], which improves the network architecture (introducing residual connectivity and adaptive normalization), optimizes the noise scheduling strategy, and proposes a more stable training objective function.
-In the same year, Ho et al.
-proposed DDPM (Denoising Diffusion Probabilistic Models) [19], which used fixed noise scheduling and reparameterization techniques and became the basis of modern diffusion models.
-Later, in order to address the deficiencies of DDPM in terms of generation quality, training efficiency, and log-likelihood optimization, and so on, Improve-DDPM (IDDPM) [44] was proposed by Nichol et al.
-in 2021, which introduced learnable variance, cosine noise scheduling, and hybrid loss function, and accelerated the sampling strategy by time-step-importance sampling, which made the DDPM effectively improved , but the computation is complicated, the dataset has limitations, the generalization ability is insufficient, and the generation quality and speed belong to this and that.Song et al.
-proposed DDIM (Denoising Diffusion Implicit Models) [35] in the same year, which accelerates the sampling through non-Markov chains.
-Meanwhile, Rombach et al.
-proposed LDM (Latent Diffusion Model) [18], which shifts the diffusion process to latent space instead of operating directly in pixel space, dramatically reducing the computational cost and ensuring the generation quality.
-The enhancement for NCSN++ was also introduced in the same year with Score-SDE (based on stochastic differential equations) [21] as well as ODE (probabilistic streaming) [21], which optimize the sampling process with stochastic and constant differentiation and optimize the sampling process, respectively.
-Also in 2021, Dhariwal and Nichol proposed Classifier Guidance (Conditional Diffusion Model/CDM) [11], is to significantly improve the quality of generation by incorporating a classifier to guide the generation.In 2022, Rombach et al.
-published Stable Diffusion [18], which is based on the LDM model improved to enhance text-to-image generation.
-There are important variants [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] that have been improved since then, but do not constitute a “landmark” model.
-Below is a timeline of the evolution of the diffusion model and its main variants:
+where $\mathcal{U}[0,T]$ denotes the uniform distribution over the interval $[0,T]$.
+The preceding discussion has introduced the fundamental principles and formulas of the basic diffusion model.
+This theoretical framework underpins the success of diffusion models in modern image processing, enabling them to overcome certain limitations of traditional models, albeit with their own inherent drawbacks.
+In the context of hyperspectral imaging, research primarily focuses on Denoising Diffusion Probabilistic Models (DDPMs), which are the most frequently applied diffusion models in this domain.
+While the other two types have seen comparatively less application, they will also be explored in subsequent sections.
+3.2 Evolution of the Diffusion model with different variants
+The genesis of Diffusion Models dates back to 2015, when Sohl-Dickstein et al.
+first introduced their theoretical framework, uniquely merging nonequilibrium thermodynamics with generative models [22].
+In 2019, Song Yang et al.
+advanced the field with Score-Based Generative Models (SGM/NCSN) [20], which leverage score matching to generate data by estimating the logarithmic gradient of the data distribution.
+This was further refined in 2020 with NCSN++ (Noise-Conditioned Score Networks++) [27], optimizing the network architecture (through residual connections and adaptive normalization), noise scheduling, and introducing a more stable training objective.
+The same year saw Ho et al.’s seminal work on DDPM (Denoising Diffusion Probabilistic Models) [19], which, by utilizing fixed noise scheduling and reparameterization, established the foundation for modern diffusion models.
+Addressing DDPM’s limitations in generation quality, training efficiency, and log-likelihood optimization, Nichol et al.
+introduced Improve-DDPM (IDDPM) [44] in 2021.
+This variant incorporated learnable variance, cosine noise scheduling, and a hybrid loss function, while also accelerating sampling through time-step-importance sampling.
+Despite these improvements, IDDPM still faced complexities in computation, dataset limitations, insufficient generalization, and trade-offs between generation quality and speed.
+Concurrently in 2021, Song et al.
+proposed DDIM (Denoising Diffusion Implicit Models) [35], accelerating sampling via non-Markovian chains.
+Significantly, Rombach et al.
+unveiled the Latent Diffusion Model (LDM) [18], which dramatically reduced computational cost and maintained generation quality by shifting the diffusion process to a latent space rather than operating directly on pixels.
+Further enhancements for NCSN++ emerged with Score-SDE (based on stochastic differential equations) [21] and ODE (probabilistic streaming) [21], both optimizing the sampling process through stochastic and constant differentiation, respectively.
+Also in 2021, Dhariwal and Nichol introduced Classifier Guidance (Conditional Diffusion Model/CDM) [11], significantly boosting generation quality by integrating a classifier to guide the process.In 2022, Rombach et al.
+released Stable Diffusion [18], an advancement based on the LDM model specifically enhanced for text-to-image generation.
+Since then, numerous important variants have been developed [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], though none have yet reached the "landmark" status of the aforementioned models.
+A timeline illustrating the evolution of the diffusion model and its principal variants is provided below.
 Figure 3: Evolution timeline of the Diffusion model
-III-C Traditional Deep Learning Models
-Generative Models (GM) is a class of machine learning models that can learn distributions from data and generate new samples.
-It includes data generation, data enhancement, image restoration, anomaly detection and so on.
-It is mainly valuable in computer vision, natural language processing and other fields.
-In this paper, we will focus on the traditional generative models: the GAN, VAE, autoregressive models.
-Generative adversarial network (GAN) is another important network framework for unsupervised anomaly detection techniques.GAN was originally conceptualized in the zero-sum game framework of game theory.
-It was proposed by Ian Goodfellow et al.
-in 2014 [8].
-The core idea is to generate data by adversarial training of two neural networks: a generator and a discriminator.
-The generator is responsible for generating data, and the discriminator is responsible for distinguishing the generated data from the real data.
-Through this adversarial process, the generator gradually learns to generate data similar to the distribution of the real data, which is fast to generate, but the training is unstable and prone to pattern collapse problem [58, 59, 60].
-Variational Auto-Encoder (VAE) [9] is a generative model that combines the advantages of Autoencoder and probabilistic graphical models to learn latent representations from data and generate new data samples.The core idea of VAE is to approximate the latent distributions of the data by Variational Inference (VII), and to generate the decoder by generating new data with lower generation quality but stable training [61, 62, 63].
-Autoregressive Models (ARMs) [64] are a class of generative models that model sequence data by assuming that the current data point is dependent on previous data points.
-The core idea of autoregressive models is to utilize the temporal dependence of sequence data to generate new data points step by step, e.g., RNN (Recurrent Neural Network) forms a recursive body in the process of processing the data; LSTM (Long Short-Term Memory Neural Network)) a kind of RNN with a special pattern.
-Autoregressive models generate the complete data by generating each part of the data step by step, which is slow in generating speed, but generating high quality [65, 66].
-III-D Successful Application of Diffusion Modeling in Image Generation
-Diffusion model has a wide range of applications, for images on the application began in 2020, mainly based on Denoising Diffusion Probabilistic Models (DDPM), first of all, the initial proposed DDPM [19], IDDM [44], and DDIM [35], etc., through the Markov chain as well as the addition of the noise and denoising process, and finally the prediction of the noise residuals to optimize the quality of the generation.
-However, the number of sampling steps is numerous and the computation speed is slow.
-Then ADM-G (Classifier-Guided-Diffusion) [11], introduced by OpenAI team in 2021, made the model surpass GAN in FID metrics for the first time by introducing classifier-guided generation.
-but dependent on classifiers and high computational cost.
-Later, in order to optimize text-image alignment and super-resolution generation, Imagen [46] was introduced by the Google Research team, which combines the T5-XXL text encoder with the cascade diffusion architecture, and has outstanding text alignment capability, but with great resource requirements.
-Later, in order to realize multimodality and efficient generation, Stabel-Diffusion [18] was proposed by Robin Rombach et al.
-in 2023, which compressed the diffusion process into the latent space, significantly reducing the computational complexity, but losing information in the latent space.
-Meanwhile, also by the OpenAI team, CLIP-Guded-Diffusion [49] was introduced, which combines the CLIP model to realize cross-modal generation, supports complex semantic control, and generates semantic consistency, but has large training requirements and slow generation speed.
-Subsequently, in order to unify the framework and extend the multi-tasking domain, OmniGen [53], the first unified framework supporting tasks such as text generation, image editing, and so on, without additional plug-ins, was introduced by the Wisdom Source Institute in 2024.
-Then, the Byte Jump team introduced D-Dit [55], which combines continuous and discrete diffusion techniques to achieve bidirectional text-image generation, support multimodal tasks, and improve parameter efficiency by about 20%.
-In 2025 by Yifan Zhou et al.
-proposed AF-LDM [57], which introduces shift isovariance loss, improves the UNet attention module, reduces potential spatial noise bias, improves video generation consistency, and enhances the model robustness, etc.
-At the same time, the diffusion model will continue to be innovated and developed in the hyperspectral field in the future, and these models are improved on the basis of the main diffusion model, aiming at solving the problem of image in classification, anomaly detection and recovery, and so on.
+3.3 Traditional Deep Learning Models
+Generative Models (GMs) are a class of machine learning models capable of learning data distributions and generating novel samples.
+Their applications are widespread in fields like computer vision and natural language processing, encompassing tasks such as data generation, data enhancement, image restoration, and anomaly detection.
+In this paper, we will primarily focus on traditional generative models, specifically Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and autoregressive models.
+The Generative Adversarial Network (GAN), proposed by Ian Goodfellow et al.
+in 2014 [8], is another significant framework for unsupervised anomaly detection, rooted in the concept of a zero-sum game.
+At its core, a GAN generates data through the adversarial training of two neural networks: a generator and a discriminator.
+The generator’s role is to produce data, while the discriminator’s task is to differentiate between this generated data and real data.
+This adversarial process drives the generator to progressively learn how to create data that closely resembles the distribution of real data.
+While GANs offer fast generation, they often suffer from unstable training and are prone to issues like mode collapse [58, 59, 60].
+The Variational Auto-Encoder (VAE) [9] is a generative model that blends the strengths of Autoencoders with probabilistic graphical models.
+Its core idea is to learn latent representations from data and subsequently generate new samples.
+VAEs achieve this by using Variational Inference (VI) to approximate the data’s latent distributions.
+While they offer stable training, the quality of the generated data can sometimes be lower [61, 62, 63].
+Autoregressive Models (ARMs) [64] are a class of generative models that operate by assuming the current data point in a sequence is dependent on preceding ones.
+Their core idea is to leverage this temporal dependence to generate new data points sequentially.
+Examples include Recurrent Neural Networks (RNNs), which employ a recursive structure for data processing, and Long Short-Term Memory (LSTM) Neural Networks, a specialized type of RNN.
+While ARMs are known for generating high-quality data, their step-by-step generation process can result in slower speeds [65, 66].
+3.4 Successful Application of Diffusion Modeling in Image Generation
+Since 2020, Diffusion Models have seen widespread application in image processing, primarily leveraging Denoising Diffusion Probabilistic Models (DDPMs).
+Early works like the initial DDPM [19], IDDM [44], and DDIM [35] optimized generation quality through Markov chains and refined noise addition/denoising processes, predicting noise residuals.
+However, these models were characterized by numerous sampling steps and slow computation.
+In 2021, the OpenAI team introduced ADM-G (Classifier-Guided-Diffusion) [11], which, by integrating classifier-guided generation, allowed the model to surpass GANs in FID metrics for the first time.
+Despite this breakthrough, it remained dependent on classifiers and incurred high computational costs.
+To enhance text-image alignment and super-resolution, the Google Research team later introduced Imagen [46], combining the T5-XXL text encoder with a cascaded diffusion architecture, resulting in outstanding text alignment capabilities but demanding substantial resources.
+Seeking multimodal and efficient generation, Robin Rombach et al.
+proposed Stable Diffusion in 2023 [18].
+This model significantly reduced computational complexity by compressing the diffusion process into a latent space, albeit with some information loss.
+Concurrently, the OpenAI team also released CLIP-Guided-Diffusion [49], which integrates the CLIP model to enable cross-modal generation and support complex semantic control, ensuring semantic consistency.
+However, this model also requires significant training resources and has a slower generation speed.
+More recently, in 2024, the Wisdom Source Institute introduced OmniGen [53], the first unified framework supporting tasks like text generation and image editing without additional plug-ins.
+Following this, the ByteDance team launched D-Dit [55], which combines continuous and discrete diffusion techniques to achieve bidirectional text-image generation, support multimodal tasks, and improve parameter efficiency by approximately 20%.
+Looking ahead to 2025, Yifan Zhou et al.
+are set to propose AF-LDM [57], aiming to enhance video generation consistency and model robustness by introducing shift isovariance loss and improving the UNet attention module to reduce potential spatial noise bias.
+Simultaneously, the diffusion model is poised for continued innovation and development within the hyperspectral domain.
+Future advancements in this area will build upon existing diffusion models, with a focus on resolving challenges in hyperspectral image classification, anomaly detection, and recovery.
 Figure 4: Improved division of the diffusion model according to different variants
-TABLE I: Basic Diffusion Models and Variants
+Table 1: Summary of basic diffusion models, variants and their characteristics
 Basic Diffusion Model Major Variants Reasons for Improvement Improvement Point Drawbacks Applications
@@
 Controlled generation, inverse problem solving.
-IV Applications in Hyperspectral Image Processing and Analysis
-Deep learning models are widely used in various fields and show powerful advantages, in recent years, some scholars have also applied them to the field of hyperspectral, while making many achievements.
-The application in hyperspectral image analysis is embodied in many aspects, mainly including data enhancement and synthesis, generation and complementation of hyperspectral data, image classification, segmentation, target detection, and anomaly detection, and many other aspects, in this section, the hyperspectral image processing is divided into two aspects: one is the image processing under the traditional deep learning model, and the other is the image processing under the diffusion-based model.
-The following is a description of these methods.
+4 Applications in Hyperspectral Image Processing and Analysis
+Deep learning models are widely adopted across various fields, demonstrating significant advantages.
+In recent years, scholars have successfully applied them to hyperspectral imaging, yielding numerous achievements.
+Their application in hyperspectral image analysis spans several areas, including data enhancement and synthesis, data generation and completion, image classification, segmentation, target detection, and anomaly detection.
+This section will categorize hyperspectral image processing into two main approaches: those using traditional deep learning models and those employing diffusion-based models.
+The following provides a detailed description of these methods.
 Figure 5: Deep learning model-based hyperspectral image processing
 Figure 6: Timeline of the birth of each model
-IV-A The main Diffusion model used in the field of hyperspectral imagery
-Diffusion Model based Hyperspectral Image Processing, which was first applied in the field of Hyperspectral in 2022, meanwhile, it was obtained through the research that the main use of Diffusion Modeling methods in this field is mainly Denoising Diffusion Probabilistic Models (DDPM), and this paper also presents the technical analysis of DDPM and its variants in the field in this section.
-The main diffusion models for hyperspectral image processing based on diffusion modeling are DDPM [19], IDDPM [44], DDIM [35], LDM [18], Classifier-Gided [11], Stable-Diffusion [18].
-First of all, the basic principles of DDPM have been introduced above, and the basic principles and formulas of IDDPM and DDIM are the same as those of DDPM, so we won’t repeat them here.LDM and Stable-Diffusion belong to the same category, in which the core ideas are the same, and it can be said that Stable-Diffusion is the specific implementation version of LDM, so the following will introduce the following LDM (Latent Diffusion Model), Classifier-Guided (Conditional Diffusion Model) two basic principles, formula.
-IV-A1 Conditional Diffusion Model (CDM)
-Conditional diffusion models are similar to the development of GANs in that diffusion models begin with unconditional generation, followed immediately by conditional generation [67].
-Unconditional generation is often used to explore the upper limits of the model’s capabilities, whereas conditional generation is more beneficial for applications as it allows the output to be controlled according to the human’s wishes.
-The first work to introduce this condition in diffusion models was [11], which guided the generation of diffusion models by adding classifiers to trained diffusion models, hence the term guided diffusion models.
-Although this approach has a low training cost, it increases the inference cost by using the classification results to guide the sampling process of the diffusion model and has poor control over the details to produce images with better results.
-Therefore, the Google team [68] decided to adopt a simple idea of controlling the generated results by conditionally retraining the DDPM and named it Classifier-Free Guidance.
-Formally, given the conditional information $\mathbf{c}$, the distribution of the combined DDPM changes from equations (2)(3), and accordingly, the optimization objective and sampling process are modified to the following form:
+4.1 The main Diffusion model used in the field of hyperspectral imagery
+Diffusion Model based Hyperspectral Image Processing was first applied to the hyperspectral field in 2022.
+Research indicates that Denoising Diffusion Probabilistic Models (DDPM) are predominantly used in this domain.
+This section will specifically analyze the technical aspects of DDPM and its variants within hyperspectral image processing.
+The primary diffusion models leveraged for this purpose include DDPM [19], IDDPM [44], DDIM [35], LDM [18], Classifier-Guided Diffusion [11], and Stable Diffusion [18].
+We’ve already covered the basic principles of DDPMs.
+Since the fundamental principles and formulas of IDDPM and DDIM are essentially identical to DDPM, we won’t reiterate them here.
+LDM (Latent Diffusion Model) and Stable-Diffusion fall under the same category, sharing core ideas, with Stable-Diffusion effectively being a specific implementation of LDM.
+Therefore, the following sections will introduce the basic principles and formulas of LDM and Classifier-Guided (Conditional Diffusion Model).
+4.1.1 Conditional Diffusion Model (CDM)
+Similar to the evolution of GANs, diffusion models initially focused on unconditional generation, swiftly transitioning to conditional generation [67].
+While unconditional generation is often used to push the boundaries of a model’s capabilities, conditional generation holds more practical value for applications, as it enables the output to be controlled based on human intent.
+The introduction of conditioning into diffusion models first appeared in [11].
+This pioneering work guided the generation of diffusion models by incorporating classifiers into pre-trained models, leading to the term "guided diffusion models." While this method offered low training costs, it increased inference costs by using classification results to steer the diffusion model’s sampling process, and it struggled to precisely control details for optimal image output.
+Consequently, the Google team [68] opted for a simpler approach: retraining the DDPM conditionally to control generated results, a technique they named Classifier-Free Guidance.
+Formally, when given conditional information $\mathbf{c}$, the combined DDPM distribution changes from equations (2) and (3).
+Consequently, both the optimization objective and the sampling process are modified to take the following form:
 $\mathcal{L}_{\mathrm{con}}(\theta)=\mathbb{E}_{x_{0},\epsilon,c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,c,t\right)\right\|^{2}\right].$ (18)
 $y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\tilde{\alpha}_{t}}}\epsilon_{\theta}(y_{t},c,t)\right)+\hat{\beta}_{t}z.$ (19)
-This approach is more widely used than guided diffusion models and is the basis for numerous models of important diffusion (e.g., DALL-E2 [45], Imagen [49], and stabilizing diffusion (SD) [18]).
-IV-A2 Latent Diffusion Model (LDM)
-The computational cost of direct manipulation of the diffusion model in the original image pixel space is high, Rombach et al [60] proposed to train and reason about the diffusion model in a low-dimensional latent space and named it Latent Diffusion Model (LDM).
-Figure 7: Architecture of LDM, where E and D denote the encoder and decoder, respectively
-As shown in Fig.
-7, the LDM consists of three parts.
-First, the specific diffusion process of the LDM, which shifts the workspace from pixel space to latent space using a pre-trained autoencoder, is shown in the leftmost gray box in Fig.
-7, where the noise is added to feature $\mathbf{z}$ to get the lower half of $\mathbf{z_{T}}$ as forward diffusion; the middle orange box is in latent space, where the U-Net with crosscutting attention supports the multimodal inputs in order to restore $\mathbf{z_{T}}$ to $\mathbf{z}$ as reverse diffusion.
-The right side box is the conditional encoder, which encodes various conditions (e.g., text and image) into feature vectors $\tau_{\theta}$ and feeds them into U-Net to generate an image.LDM is able to effectively reduce computational cost without degrading the quality of image synthesis while speeding up the inference speed, which is widely used in natural scenes and provides inspiration for the application of diffusion models to hyperspectral applications.
-IV-B Application of Diffusion model in image processing
-IV-B1 Data Enhancement and Synthesis
-In data enhancement and synthesis, usually super-resolution techniques have an important role and unique advantages in improving quality.
-Super-resolution aims at reconstructing high-resolution images from low-resolution images [69], but the images can be disturbed, due to imaging equipment, weather conditions or downsampling.
-The loss of detail in hyperspectral images can be significant when compared to natural images, making inverted reconstruction of high-resolution images more challenging.
-Before the advent of diffusion models, deep neural network based approaches, e.g., Convolutional Neural Networks (CNNs) and Transformers, were finding suitable mapping functions to integrate feature information into a single step.
-To achieve this goal, methods including self-supervised learning [70], adaptive marker selection [71] and frequency information assistance [72] were used.
-However, the diffusion model-based approach fuses low-resolution images and bootstrap images in each step of the diffusion process, which helps to better integrate different image information and, at the same time, has gained wide attention in the hyperspectral field.
-Although HSI has a high spectral resolution, its spatial resolution is relatively low [73], which may limit the performance of various HSI-based applications.
-To obtain high-resolution hyperspectral maps, there are two types of methods: panchromatic sharpening [74] and multispectral and HSI fusion [75].
-The former enhances HSI by injecting detail information in the panchromatic (PAN) image, while the latter utilizes multispectral images to help hyperspectral maps learn spatial details.
-For example, Shi et al [76] used multispectral and hyperspectral images in tandem as conditional inputs to a diffusion model, and then generated hyperspectral maps with high resolution from the information found in the two image modalities.Qu et al [77] devised a conditional recurrent diffusion framework that allows the spatial SR process and the spectral SR process to be guided by the multispectral image and the hyperspectral map, respectively, to generate high-resolution hyperspectral maps with complementary spatial and spectral information to generate high-resolution hyperspectral maps.
-More contemporary research focuses on the use of PAN image HSI super-resolution [78, 79, 80, 81, 82, 83, 84].Meng et al [79] proposed a modal calibration module to enhance and extract features in two images, where the enhanced features are used as conditional inputs to the diffusion model.Cao et al [78] argued that the different image modal unique information should not be mixed for processing and designed two conditional modulation modules to extract coarse-grained style information and fine-grained frequency information as conditional inputs.
+This approach has seen broader adoption than guided diffusion models and forms the foundation for many significant diffusion models, including DALL-E2 [45], Imagen [49], and Stable Diffusion (SD) [18].
+4.1.2 Latent Diffusion Model (LDM)
+Directly manipulating diffusion models in the original image pixel space incurs high computational costs.
+To address this, Rombach et al.
+[60] proposed training and performing inference with diffusion models within a low-dimensional latent space, terming this approach the Latent Diffusion Model (LDM).
+Figure 7: Architecture of LDM, where E and D denote the encoder and decoder, respectively [18]
+As depicted in Fig.
+7, the Latent Diffusion Model (LDM) comprises three distinct parts.
+Firstly, the leftmost gray box illustrates the LDM’s specific diffusion process: it shifts its operational domain from pixel space to a low-dimensional latent space by utilizing a pre-trained autoencoder.
+Within this space, noise is incrementally added to feature $\mathbf{z}$, yielding $\mathbf{z_{T}}$ as part of the forward diffusion.
+Secondly, the central orange box, also in latent space, houses a U-Net with cross-attention that supports multimodal inputs.
+This U-Net facilitates the reverse diffusion, restoring $\mathbf{z_{T}}$ back to $\mathbf{z}$.
+Lastly, the rightmost box contains the conditional encoder, which transforms various conditions (such as text and images) into feature vectors $\tau_{\theta}$.
+These vectors are then fed into the U-Net to guide image generation.
+The LDM’s ability to effectively reduce computational cost without sacrificing image synthesis quality, coupled with its accelerated inference speed, has led to its widespread adoption in natural scene generation and offers significant inspiration for applying diffusion models to hyperspectral applications.
+4.2 Application of Diffusion model in image processing
+4.2.1 Data Enhancement and Synthesis
+In data enhancement and synthesis, super-resolution techniques play a crucial role in improving image quality.
+Super-resolution aims to reconstruct high-resolution images from their low-resolution counterparts [69].
+However, these images can be degraded by factors like imaging equipment, weather conditions, or downsampling.
+This loss of detail is often more pronounced in hyperspectral images compared to natural images, making the inverse problem of reconstructing high-resolution versions significantly more challenging.
+Before the advent of diffusion models, approaches based on deep neural networks, such as Convolutional Neural Networks (CNNs) and Transformers, sought to integrate feature information in a single step by finding suitable mapping functions.
+Methods like self-supervised learning [70], adaptive marker selection [71], and frequency information assistance [72] were employed to achieve this.
+In contrast, diffusion model-based approaches fuse low-resolution images and bootstrap images at each step of the diffusion process.
+This iterative integration of different image information has proven highly effective and has garnered significant attention in the hyperspectral field.
+Although Hyperspectral Images (HSIs) boast high spectral resolution, their spatial resolution is often comparatively low [73], which can limit the effectiveness of various HSI-based applications.
+To overcome this, two primary methods are employed to achieve high-resolution hyperspectral maps: panchromatic sharpening [74] and multispectral and HSI fusion [75].
+Panchromatic sharpening enhances HSIs by injecting detailed spatial information from a high-resolution panchromatic (PAN) image.
+Conversely, multispectral and HSI fusion leverages multispectral images to help hyperspectral maps acquire spatial details.
+For instance, Shi et al.
+[76] utilized both multispectral and hyperspectral images as conditional inputs to a diffusion model, subsequently generating high-resolution hyperspectral maps by integrating information from these two image modalities.
+Similarly, Qu et al.
+[77] developed a conditional recurrent diffusion framework.
+This framework enables the spatial super-resolution process to be guided by the multispectral image and the spectral super-resolution process by the hyperspectral map, leading to the generation of high-resolution hyperspectral maps that boast complementary spatial and spectral information.
+More contemporary research is increasingly focused on utilizing Panchromatic (PAN) images for HSI super-resolution [78, 79, 80, 81, 82, 83, 84].
+For instance, Meng et al.
+[79] introduced a modal calibration module to enhance and extract features from both image types, subsequently using these enhanced features as conditional inputs for a diffusion model.
+Conversely, Cao et al.
+[78] posited that the unique information from different image modalities should not be mixed during processing.
+They instead designed two conditional modulation modules to separately extract coarse-grained style information and fine-grained frequency information, which are then used as conditional inputs.
 Figure 8: Converged network fusion mode [82]
-In order to take full advantage of the unique information of different modalities, Li et al [82] proposed a panchromatic sharpening network based on a biconditional diffusion model, where HSI and PAN are used as independent conditional inputs to learn the spectral features and spatial texture, respectively.
-The network structure is designed based on LDM and achieves excellent SR performance while significantly reducing the computational cost.Rui et al [84] also perform the sampling process in low-dimensional space.
-Based on the fact that the high-resolution hyperspectral map can be decomposed into the product of two low-rank tensors, one of the low-rank tensors is first computed using the low-resolution hyperspectral map.
-This tensor, together with the LRHSI and PAN, is then conditioned and fed into a pre-trained diffusion model [85] for the other low-rank tensor.
-Unlike the above methods, this method is a completely unsupervised deep learning method that does not require the involvement of high-resolution hyperspectral maps in any of the processes, providing feasibility for its application in practice.
-IV-B2 Data Generation and Completion
-Generation and Complementation in general can be categorized as text-to-image and image-to-image, but in the hyperspectral domain only image-to-image exists.
-For diffusion models, diffusion models have image bootstrapping to generate new images.
-Guided images can be in various forms such as maps [86], semantic layouts [87, 88, 89, 90].
-Even though these images contain only some specific information, they still show excellent performance in image generation.
-For example, Espinosa and Crowley [86] proposed to use maps and historical maps to train conditional diffusion models, which generated more realistic satellite images.Yuan et al.
-[87] generated high-quality hyperspectral images under the guidance of semantic masks, and also solved the intrinsic problem that diffusion models require a long time of training for model convergence.
-However, these things alone are not enough, so it is important to generate not only high-quality images, but also image-annotation pairs that are more practical.
-For example, Zhao et al [89] implemented a two-stage fine-tuning of the SD model to enable the generation from noise to annotations and then from annotations to images.Toker et al [90] extended the standard diffusion model to a joint probabilistic model for images and their corresponding labels, which enabled the simultaneous generation of labels and images.
-Hyperspectral images (HSI) can provide richer spectral spatial information.
-Meanwhile, the acquisition cost of HSI is much higher than that of multispectral images.
-Therefore, subsequent scientists then used easily available multispectral remote sensing images to generate hyperspectral images [91, 92].
-However, the hyperspectral maps generated by the diffusion model used need to match the input noise dimensions with the spectral bands, which leads to too much noise sampling space, thus making the model convergence difficult.
-To solve this problem, Zhang et al [91] proposed a spectral folding technique to convert the input hyperspectral map into a pseudo-color image before training the diffusion model.Inspired by the LDM, Liu et al [92] used Conditional Vector Quantization GAN (VQGAN) [93] to obtain the Latent Diffusion Model latententent space for the hyperspectral map, and performed the diffusion model training and sampling in this space process.
-Afterwards, in order to address the status quo of high cost and low spatial resolution, a DM-based hyperspectral spectral decomposition method, DiffUn [94], was proposed by Deng et al.
-The spectral prior distributions learned by unconditional DM and the likelihoods derived from the LMM were combined to facilitate the deblurring performance and synthesize the hyperspectral maps.A novel ISPDiff [95] model was proposed by Dong et al.
-A transparent unified scale-flexible model was designed to solve such problems using an iterative optimization algorithm with MSRN for scale-matching recovery and CMUN for cross-scale upsampling.
-Overall, image-to-image generation methods based on conditional diffusion models [67, 68], which use bootstrap images as conditional inputs to conditional DDPM, are essentially generating images from noise rather than directly.
-In order to realize a true image-to-image conversion, Yu et al.
-proposed UnmixDiff [96] to address the feasibility of depth generation model in HSI synthesis task, which synthesizes HSI from the perspective of material distribution and introduces a clear physical interpretation into the hyperspectral image synthesis task.
-By introducing an unmixing autoencoder in the diffusion generation model, the hyperspectral image synthesis is projected from the high-dimensional spectral domain to the low-dimensional abundance domain, which greatly reduces the computational complexity and effectively improves the spectral confidence of the generated hyperspectral images.
-By constructing a diffusion generation model in the abundance space, at the same time, hsi with real spatial details can be generated to maintain spectral consistency.
-Also at the same time this type of approach, which first employs self-supervised learning to extract the feature embeddings of the input patches, and then uses these embeddings as conditional inputs to guide the learning of the diffusion model.
-Notably, this preserves the spatial arrangement of each color patch in the original image so that the generated patches can be assembled into a large and coherent image.
+To fully leverage the unique information from different modalities, Li et al.
+[82] proposed a panchromatic sharpening network based on a biconditional diffusion model.
+In this architecture, both HSI and PAN images serve as independent conditional inputs, allowing the network to separately learn spectral features and spatial textures.
+Designed based on LDM, this network achieves excellent super-resolution (SR) performance while significantly reducing computational cost.
+Similarly, Rui et al.
+[84] also perform the sampling process in a low-dimensional space.
+Their method is rooted in the fact that a high-resolution hyperspectral map can be decomposed into the product of two low-rank tensors.
+First, one of these low-rank tensors is computed using the low-resolution hyperspectral map.
+This tensor, along with the LRHSI (low-resolution hyperspectral image) and PAN, is then conditioned and fed into a pre-trained diffusion model [85] to generate the other low-rank tensor.
+Unlike the aforementioned methods, this approach is a completely unsupervised deep learning method, meaning it doesn’t require high-resolution hyperspectral maps during any part of the process, making it highly feasible for practical applications.
+4.2.2 Data Generation and Completion
+In the realm of generation and complementation, while broad categories include text-to-image and image-to-image synthesis, the hyperspectral domain exclusively features image-to-image applications.
+Here, diffusion models leverage image bootstrapping to create new images.
+These models can be guided by various forms of input, such as maps [86] or semantic layouts [87, 88, 89, 90].
+Even with only specific information in these guiding images, excellent performance in image generation is still observed.
+For instance, Espinosa and Crowley [86] demonstrated that training conditional diffusion models with maps and historical maps could generate more realistic satellite images.
+Similarly, Yuan et al.
+[87] successfully produced high-quality hyperspectral images guided by semantic masks, simultaneously addressing the inherent challenge of lengthy training times for diffusion model convergence.
+However, generating high-quality images alone isn’t always sufficient; creating more practical image-annotation pairs is often crucial.
+To this end, Zhao et al.
+[89] implemented a two-stage fine-tuning of the Stable Diffusion (SD) model, enabling generation from noise to annotations and then from annotations to images.
+Furthermore, Toker et al.
+[90] extended the standard diffusion model into a joint probabilistic model for images and their corresponding labels, allowing for the simultaneous generation of both.
+Hyperspectral images (HSIs) offer rich spectral and spatial information, but their acquisition cost is significantly higher than that of multispectral images.
+Consequently, scientists have explored using readily available multispectral remote sensing images to generate HSIs [91, 92].
+However, a challenge arises because the diffusion models used for this task require the input noise dimensions to match the spectral bands of the hyperspectral map.
+This can lead to an excessively large noise sampling space, making model convergence difficult.
+To address this, Zhang et al.
+[91] introduced a spectral folding technique that converts the input hyperspectral map into a pseudo-color image before training the diffusion model.Inspired by the LDM, Liu et al.
+[92] utilized a Conditional Vector Quantization Generative Adversarial Network (VQGAN) [93] to derive a latent space for the hyperspectral map within the Latent Diffusion Model.
+Diffusion model training and sampling are then performed within this more manageable latent space.Furthermore, to tackle the issues of high cost and low spatial resolution, Deng et al.
+proposed DiffUn [94], a Diffusion Model-based hyperspectral spectral unmixing method.
+This approach combines spectral prior distributions learned by an unconditional Diffusion Model with likelihoods derived from the Linear Mixing Model (LMM) to enhance deblurring performance and synthesize hyperspectral maps.More recently, Dong et al.
+introduced a novel ISPDiff [95] model.
+This transparent, unified, and scale-flexible model addresses these problems using an iterative optimization algorithm.
+It incorporates a Multi-Scale Residual Network (MSRN) for scale-matching recovery and a Cross-Modal UNet (CMUN) for cross-scale upsampling.
+Overall, image-to-image generation methods based on conditional diffusion models [67, 68] fundamentally generate images from noise, rather than directly converting existing ones, by using bootstrap images as conditional inputs to conditional DDPMs.
+To achieve a true image-to-image conversion, Yu et al.
+proposed UnmixDiff [96].
+This model explores the feasibility of depth generative models in HSI (Hyperspectral Image) synthesis, synthesizing HSIs from the perspective of material distribution and introducing a clear physical interpretation into the task.
+By integrating an unmixing autoencoder within the diffusion generation model, hyperspectral image synthesis is projected from the high-dimensional spectral domain to a low-dimensional abundance domain.
+This significantly reduces computational complexity and effectively boosts the spectral fidelity of the generated HSIs.
+Constructing a diffusion generation model in the abundance space also allows for the generation of HSIs with realistic spatial details while maintaining spectral consistency.
+This type of approach notably employs self-supervised learning to extract feature embeddings from input patches.
+These embeddings then serve as conditional inputs, guiding the diffusion model’s learning process.
+Crucially, this method preserves the spatial arrangement of each color patch from the original image, enabling the generated patches to be seamlessly assembled into a large and coherent image.
 Figure 9: UnmixDiff generates a simple diagram of the HIS [96]
-IV-B3 Classification
-Among the applications in the hyperspectral field, classification task is very important in image analysis as it can accurately identify and characterize different structures and anomalies in an image.
-It helps professionals in their ability to interpret large amounts of complex data Despite this potential, image classification aims at obtaining more valuable information about detailed features, such as buildings or grass, by assigning each pixel to a specific class [97, 98].
-However, hyperspectral images contain many diverse and complex scenes and information, which increases the difficulty of accurate classification.
-Therefore, the current use of diffusion model to enhance classification is still a major challenge that needs to be further strengthened and solved.
-Meanwhile, because diffusion models can learn and simulate more complex data distributions than other deep learning models, more and more people are committed to applying diffusion models to hyperspectral image classification, which leads to better and better classification results [99, 100, 101, 102, 103, 104].
-For image classification, hyperspectral images exhibit more complex data distributions, posing greater challenges to the application of traditional deep learning models in classification.
-The good thing, however, is that diffusion models can better capture the spectral-spatial joint features of hyperspectral images [105, 106, 107], improve the classification accuracy [108, 109], and improve the generalization ability [110].
-Based on the diffusion model to realize the above properties, Zhou et al [111] used the temporal information of the diffusion model to build a time-stepping feature library, and also proposed a dynamic fusion module to combine spectral-spatial features with temporal features, and these improvements can make it possible to obtain sufficient image information before classification.Li et al [112] designed a two-branched diffusion model, which extracts features from HSI and LiDAR images, which achieves complementarity of information between different modal images, enhances the distinguishability between pixels, and also shows excellent performance in more complex hyperspectral tasks [37].Qu et al [113] go deeper by setting the encoder of the two-branch diffusion model to operate in parameter-sharing mode, which ensures that the shared features are extracted from the multimodal images.
-In addition, Chen et al [114] utilized the diffusion model to assist deep subspace construction and achieved excellent HSI classification performance in an unsupervised manner.
-Different from the above approaches, Qu et al [115] used a diffusion model to combine the SR (super-resolution) network and classification network to progressively reconstruct high-quality images by iteratively iterating the image super-resolution and classification tasks to achieve mutual guidance of the two tasks.
-The multi-scale classification results are utilized to guide the target-scale image to learn new knowledge and further reconstruct high-quality images.
-The super-resolution performance and classification performance are significantly improved.Jiang et al.
-proposed an image fusion [107] classification network based on Denoising Diffusion Probabilistic Models (DDPM), and after fusion, the classification accuracy is significantly improved.Zhu et al.
-proposed a Spatial-Spectral Diffusion Comparison Representation Network (DiffCRN) based on Denoising Diffusion Probabilistic Models (DDPM) [102], which is structured spatially self-attentive denoising module (SSAD) and spectral group self-attentive denoising module (SGSAD), and improves the efficiency of spectral spatial feature learning by designing new DDPM models loss and CL with log absolute error (LAE), as well as designing an adaptive weighted addition module (AWAM) and cross-time step spectral spatial fusion module (CTSSFM), which can effectively mitigate the interference due to spatial spectral heterogeneity and noise effects, etc., to effectively mitigate the interference due to spatial spectral heterogeneity and better classification.
+4.2.3 Classification
+Among applications in the hyperspectral field, the classification task is paramount for image analysis, enabling accurate identification and characterization of diverse structures and anomalies within an image.
+It significantly aids professionals in interpreting large volumes of complex data.
+Despite this potential, image classification aims to extract more valuable information about detailed features, such as buildings or grass, by assigning each pixel to a specific class [97, 98].
+However, hyperspectral images inherently contain highly diverse and complex scenes and information, which complicates accurate classification.
+Therefore, leveraging diffusion models to enhance classification remains a significant challenge that requires further strengthening and resolution.
+Nevertheless, because diffusion models can learn and simulate more intricate data distributions than other deep learning models, there’s a growing commitment to applying them to hyperspectral image classification, leading to increasingly promising results [99, 100, 101, 102, 103, 104].
+For image classification, the complex data distributions inherent in hyperspectral images present greater challenges for traditional deep learning models.
+Fortunately, diffusion models excel at capturing the joint spectral-spatial features of these images [105, 106, 107], leading to improved classification accuracy [108, 109] and enhanced generalization ability [110].
+Building on these strengths, Zhou et al.
+[111] utilized the temporal information of diffusion models to construct a time-stepping feature library.
+They also introduced a dynamic fusion module to combine spectral-spatial features with temporal features, allowing for sufficient image information before classification.
+Li et al.
+[112] designed a two-branched diffusion model that extracts features from both HSI and LiDAR images, achieving information complementarity across different modalities.
+This approach boosts pixel discriminability and demonstrates excellent performance in more complex hyperspectral tasks [37].
+Qu et al.
+[113] further deepened this by configuring the encoder of their two-branch diffusion model to operate in a parameter-sharing mode, ensuring shared features are extracted from multimodal images.
+Additionally, Chen et al.
+[114] leveraged diffusion models to assist in deep subspace construction, achieving impressive unsupervised HSI classification performance.
+Taking a different approach, Qu et al.
+[115] combined a diffusion model with both a super-resolution (SR) network and a classification network.
+This allowed for the progressive reconstruction of high-quality images by iteratively performing image super-resolution and classification tasks, creating a mutual guidance mechanism.
+By utilizing multi-scale classification results to guide the target-scale image in learning new knowledge and further reconstructing high-quality images, both super-resolution and classification performance were significantly improved.
+Jiang et al.
+proposed an image fusion classification network based on Denoising Diffusion Probabilistic Models (DDPMs) [107], which demonstrated a significant improvement in classification accuracy after fusion.
+Similarly, Zhu et al.
+introduced a Spatial-Spectral Diffusion Comparison Representation Network (DiffCRN), also based on DDPMs [102].
+This network incorporates a spatially self-attentive denoising module (SSAD) and a spectral group self-attentive denoising module (SGSAD).
+It enhances the efficiency of spectral-spatial feature learning by designing new DDPM loss functions and Contrastive Learning (CL) with Log Absolute Error (LAE).
+Furthermore, the inclusion of an adaptive weighted addition module (AWAM) and a cross-time step spectral spatial fusion module (CTSSFM) effectively mitigates interference from spatial spectral heterogeneity and noise effects, leading to better classification.
 Figure 10: Diffusion model based image classification process diagram
-IV-B4 Segmentation
-Image segmentation is an important task in computer vision, like semantic segmentation [116], image segmentation [14, 117], and in the field of deep learning, semantic segmentation is a subtask of image segmentation [112, 118].
-Hyperspectral image segmentation, which investigates simplifying the complexity of an image by decomposing it into multiple meaningful image segments.
-Specifically, it facilitates image analysis by providing useful information about hyperspectral related fields.
-However, deep learning models typically require a large number of different pixel-annotated training data to produce generalizable results.
-Nonetheless, the number of images and labels that can be used for hyperspectral image segmentation is limited due to the time, cost, and expertise required.
-For this reason, diffusion modeling has emerged as a promising approach in image segmentation research that synthesizes labeled data and eliminates the need for pixel annotation data [102, 104].
-Dong et al [95] are using Denoising Diffusion Probabilistic Models combined with Scale Matching Recovery Networks (P2MSRN) and cross-scale model-driven up-sampling networks.Chen et al [119] design image conditional mixing guides and combine them with Denoising Diffusion Probabilistic Models.Zhang et al [110] utilize Diffusion Probabilistic Models for labeling information.Zhu et al [120] utilize Variational Auto-Encoder as well as semi-supervised diffusion model for labeling annotation.Li et al.
-proposed a multidomain diffusion-driven feature learning network (MDFL) [112] to improve the pixel-level classification (i.e., segmentation) performance of high-dimensional remote sensing images through a multidomain diffusion model and feature reuse mechanism.Zhu et al.
-proposed a diffusion-based diffusion model coupled with contrast learning for diffusion contrast representation networks (DiffCRNs) [102] to extract the spatial-spectral features and finally generates pixel-level classification circles.
-Xu et al.
-proposed a data-enhanced semantic segmentation framework based on the diffusion model [104], which generates pseudo-labeled data through the diffusion model, and uses ControlNet to achieve cross-modal generation of “label→multispectral image”, and multi-stage training, which improves the segmentation accuracy.
-IV-B5 Target Detection
-Target detection differs from pixel-level ground classification in that target detection for hyperspectral images uses bounding boxes to localize instances like an airplane, a vehicle, or a ship in the image [121, 122].
-One of the most challenging problems in target detection is that the training data is ambiguous, disturbed by external conditions [123, 118], and lacks realism and diversity.
-This problem is due to the intrinsic nature of hyperspectral images, which usually results in objects that need to be detected being complex, difficult to separate, and distributed in different regions of the image.
-Therefore, augmenting the objects to be detected with diffusion models has become an effective solution.
-In order to solve all the above mentioned problems, Yu et al.
-proposed Diff-Mosaic (Enhanced Target Detection via Diffusion Prior), a diffusion model based approach [118], which introduces an enhancement network called Pixel-Prior, which coordinates pixels to generate highly coordinated and lifelike mosaic images, using a powerful diffusion model prior that resamples the pixels with samples to introduce real-world information.
-This resampling makes the generated samples realistic and diverse.
+4.2.4 Segmentation
+Image segmentation is a crucial task in computer vision, encompassing areas like semantic segmentation [116] and general image segmentation [14, 117].
+In deep learning, semantic segmentation is considered a subtask of image segmentation [112, 118].
+Hyperspectral image segmentation specifically aims to simplify image complexity by breaking it down into multiple meaningful segments.
+This process provides valuable information for hyperspectral-related fields, thereby facilitating image analysis.
+However, deep learning models typically demand a large volume of diverse pixel-annotated training data to yield generalizable results.
+Unfortunately, the number of images and labels available for hyperspectral image segmentation is limited due to the significant time, cost, and expertise required for annotation.
+Consequently, diffusion modeling has emerged as a promising avenue in image segmentation research, as it can synthesize labeled data, thereby eliminating the need for extensive pixel annotation [102, 104].
+Dong et al.
+[95] are utilizing Denoising Diffusion Probabilistic Models in conjunction with Scale Matching Recovery Networks (P2MSRN) and cross-scale model-driven up-sampling networks.
+Chen et al.
+[119] are designing image conditional mixing guides and integrating them with Denoising Diffusion Probabilistic Models.
+For labeling information, Zhang et al.
+[110] leverage Diffusion Probabilistic Models, while Zhu et al.
+[120] employ Variational Auto-Encoders alongside semi-supervised diffusion models for annotation.
+Li et al.
+proposed a multidomain diffusion-driven feature learning network (MDFL) [112] to enhance pixel-level classification (i.e., segmentation) performance of high-dimensional remote sensing images through a multidomain diffusion model and feature reuse mechanism.
+Zhu et al.
+also proposed DiffCRNs [102], a diffusion-based model coupled with contrastive learning, to extract spatial-spectral features and ultimately generate pixel-level classification circles.
+Furthermore, Xu et al.
+introduced a data-enhanced semantic segmentation framework based on the diffusion model [104].
+This framework generates pseudo-labeled data via the diffusion model, uses ControlNet for cross-modal "label→multispectral image" generation, and employs multi-stage training to improve segmentation accuracy.
+4.2.5 Target Detection
+In contrast to pixel-level ground classification, target detection for hyperspectral images focuses on localizing specific instances, such as an airplane, vehicle, or ship, within the image using bounding boxes [121, 122].
+A significant challenge in this area is the ambiguity and lack of realism and diversity in training data, often exacerbated by external disturbances [123, 118].
+This issue stems from the inherent nature of hyperspectral images, where objects needing detection are frequently complex, difficult to distinguish, and distributed across various image regions.
+Consequently, augmenting the objects to be detected using diffusion models has emerged as an effective solution.
+To address the aforementioned challenges, Yu et al.
+proposed Diff-Mosaic (Enhanced Target Detection via Diffusion Prior) [118].
+This diffusion model-based approach introduces an enhancement network called Pixel-Prior, which coordinates pixels to generate highly realistic and diverse mosaic images.
+It leverages a powerful diffusion model prior that resamples pixels with samples, thereby introducing real-world information and ensuring the generated samples are both realistic and varied.
 Similarly, Qi et al.
-proposed a novel joint framework for underwater hyperspectral image restoration and target detection based on conditional diffusion framework [123], which uses a variable spectral group extraction module and a joint underwater hyperspectral image restoration and target detection (JURTD) module to simulate different ranges of underwater target spectra, achieve high-quality restored images and superior detection performance, and effectively overcome the challenges of underwater hyperspectral imaging.
-spectral imaging challenges.Zhu et al.
-proposed ODGEN (Domain-specific Object Detection Data Generation with Diffusion Models), a novel approach [124] to generate high-quality images conditioned on bounding boxes, thereby facilitating data synthesis for object detection.The use of synthetic visual cues with spatial constraints and textual descriptions by object to control the diffusion model allows ODGEN to perform well at demonstrate robustness when dealing with complex scenes and specific domains.
-Meanwhile, in order to improve the precision and accuracy, Wang et al.
-proposed OrientedDiffDet (pairwise object-oriented diffusion model) [125], which combines oriented frame representation, feature alignment and accelerated diffusion, a process that detects objects directly from a set of random frames.
-In the training phase, the horizontal detection boxes are first converted to oriented detection boxes.
-The model then learns to reverse this conversion process by diffusing from ground truth oriented boxes to a random distribution.
-In the inference phase, the model progressively refines a set of randomly generated boxes to generate the final output.
+presented a novel joint framework for underwater hyperspectral image restoration and target detection, based on a conditional diffusion framework [123].
+This framework employs a variable spectral group extraction module and a joint underwater hyperspectral image restoration and target detection (JURTD) module to simulate different ranges of underwater target spectra.
+This achieves high-quality restored images and superior detection performance, effectively overcoming the inherent challenges of underwater hyperspectral imaging.
+Zhu et al.
+introduced ODGEN (Domain-specific Object Detection Data Generation with Diffusion Models) [124], a novel approach to generate high-quality, bounding-box-conditioned images.
+This facilitates data synthesis specifically for object detection.
+By utilizing synthetic visual cues with spatial constraints and textual descriptions per object to control the diffusion model, ODGEN demonstrates robust performance when handling complex scenes and specific domains.
+Meanwhile, to enhance precision and accuracy, Wang et al.
+proposed OrientedDiffDet (pairwise object-oriented diffusion model) [125].
+This approach combines oriented frame representation, feature alignment, and accelerated diffusion, allowing for direct object detection from a set of random frames.
+During the training phase, horizontal detection boxes are initially converted to oriented detection boxes.
+The model then learns to reverse this conversion by diffusing from ground truth oriented boxes to a random distribution.
+In the inference phase, the model progressively refines a set of randomly generated boxes to produce the final output.
 Figure 11: OrientedDiffDet target detection diagram with targets boxed in different colors in the detection diagram [125]
-IV-B6 Anomaly Detection
-Image anomaly detection is an important topic in computer vision, aiming to highlight anomalous regions of an image [115, 126, 127, 128].
-In recent years, generative modeling has greatly influenced the queries for anomaly detection and shown promising results.
-Therefore, it will be explored in this section how the diffusion model, as the main generative model in anomaly detection, can be applied in hyperspectral collar anomaly detection.
-In fact, there are relatively few studies on hyperspectral image anomaly detection based on diffusion modeling, however, there are relatively numerous applications in fields such as medical impact areas [129, 130], industrial images [131, 132], etc., so with the development of diffusion modeling in this hyperspectral image anomaly detection will be increasingly used.
-Meanwhile, in the field of anomaly detection, it is often faced that on the one hand, the complexity of the background of the input hyperspectral image leads to the confusion of anomalies in the background.
-On the other hand, the lack of labeled samples for hyperspectral images leads to poor generalization of existing anomaly detection methods.
-To solve this problem, there Ma et al.
-proposed a novel solution for anomaly detection, BSDM (Background Suppression Diffusion Model) [133], which simultaneously learns the Latent Background distribution and generalizes it to different datasets to suppress the complex background.
-It is characterized by three aspects:(1) Pseudo background noise is designed for the complex background of HSIs and Diffusion Model (DM) is used to learn the latent background distribution therein.
-(2) For the generalizability problem, a statistical bias module is applied to adapt the BSDM to datasets with different domains without labeling the samples.
-(3) To achieve background suppression, the inference process of DM was innovatively improved by feeding the original hyperspectral image into a denoising network, which removes the background as noise.
-They were also the first to apply diffusion modeling to hyperspectral image anomaly detection.
+4.2.6 Anomaly Detection
+Image anomaly detection, a critical area in computer vision, focuses on highlighting unusual regions within an image [115, 126, 127, 128].
+In recent years, generative modeling has significantly influenced anomaly detection research, showing promising results.
+This section will explore how diffusion models, as a prominent generative model, can be applied to hyperspectral image anomaly detection.
+While studies specifically on diffusion model-based hyperspectral image anomaly detection are relatively few, there are numerous applications in other fields, such as medical imaging [129, 130] and industrial inspection [131, 132].
+Given the ongoing advancements in diffusion modeling, its use in hyperspectral image anomaly detection is expected to grow significantly in the future.
+In the field of anomaly detection, a persistent challenge is the complexity of background clutter in hyperspectral images, which often obscures anomalies.
+Additionally, the scarcity of labeled hyperspectral samples leads to poor generalization in existing anomaly detection methods.
+To tackle these issues, Ma et al.
+proposedBSDM (Background Suppression Diffusion Model) [133], a novel anomaly detection solution.
+BSDM simultaneously learns the latent background distribution and generalizes it across different datasets to suppress complex backgrounds.
+This approach is characterized by three key aspects: 1.
+Pseudo background noise is designed for the complex backgrounds of HSIs, and a Diffusion Model (DM) is employed to learn the underlying latent background distribution.
+2.
+To address generalizability, a statistical bias module adapts BSDM to datasets from various domains without requiring labeled samples.
+3.
+For effective background suppression, the DM’s inference process was innovatively improved by feeding the original hyperspectral image into a denoising network, which removes the background as if it were noise.
+This work also marked the first application of diffusion modeling to hyperspectral image anomaly detection.
 Wu et al.
-proposed the diffusion background dictionary method (DBD) [134], which combines the advantages of model-driven and data-driven based on the characteristics of HSI and HAD tasks.DM is intrinsically combined with the low-rank representation (LRR) model, and DM is utilized to obtain the key background dictionary tensor in the tensor LRR.
-Thus, anomalies can be accurately detected.
-According to the idea of RX algorithm in anomaly detection, the multivariate normal distribution approximating the HSI background is diffused to make it more suitable for suppressing the background.
-Also hyperspectral image anomaly detection is often faced with the lack of a priori information, essentially the challenge is to accurately reconstruct the background region while inferring the latent background of the anomalous region, due to which Chen et al.
-proposed a novel approach called Double Window Spectral Diffusion (DWSDiff) [56], which modeled accurate background estimation through an iterative spectral diffusion and reverse diffusion process.
-A dual-window strategy is also introduced to mitigate the effect of anomalous extended regions in the neighborhood on the background estimation.
-An anomaly generation strategy based on Principal Component Analysis (PCA) and Linear Spectral Mixture Model (LSMM) is devised to create realistic training data that overcomes the scarcity of labeled hyperspectral data.
-The same effect is improved.
+introduced the diffusion background dictionary method (DBD) [134], combining the strengths of model-driven and data-driven approaches based on HSI and HAD (Hyperspectral Anomaly Detection) task characteristics.
+DBD intrinsically integrates the DM with the low-rank representation (LRR) model, utilizing the DM to obtain the crucial background dictionary tensor within the tensor LRR, thus enabling accurate anomaly detection.
+Following the principles of the RX algorithm in anomaly detection, the multivariate normal distribution approximating the HSI background is diffused to better suit background suppression.
+Furthermore, hyperspectral image anomaly detection often grapples with a lack of prior information, primarily the challenge of accurately reconstructing background regions while inferring the latent background of anomalous areas.
+To this end, Chen et al.
+proposed Double Window Spectral Diffusion (DWSDiff) [56].
+This novel approach models accurate background estimation through an iterative spectral diffusion and reverse diffusion process.
+A dual-window strategy is also introduced to mitigate the impact of extended anomalous regions in the neighborhood on background estimation.
+Finally, an anomaly generation strategy based on Principal Component Analysis (PCA) and the Linear Spectral Mixture Model (LSMM) is devised to create realistic training data, overcoming the scarcity of labeled hyperspectral data and leading to improved performance.
 Figure 12: Various diagrams under BSDM from left to right: anomaly diagram, 1D t-SN diagram, 2D t-SNE scatter diagram [133]
-In conclusion, the road of diffusion model in hyperspectral image anomaly detection is still long and the research prospect is very great, more and more researchers will put their eyes here to promote the vigorous development of the field of anomaly detection.
-IV-B7 Noise Suppression
-The main challenge in hyperspectral imagery is to acquire images without losing important information.
-During the acquisition and/or further processing stages, the acquired image may be corrupted by noise or artifacts.
-Noise degrades the image quality and is especially noticeable when the objects are small and the contrast is relatively low.
-Due to the nature of the generative model, the diffusion model facilitates the solution of various noise reduction problems [135, 136, 137].
-Among them, [136] is denoising in image restoration, where the robustness of denoising is enhanced by designing a multi-shape spatial rectangular self-attention module, a spectral latent diffusion enhancement module, and also by utilizing the diffusion model, which allows for a tighter alignment with the desired HSI.
-In image processing, hyperspectral images are always accompanied by various noises due to the inherent limitations of imaging techniques and environmental conditions.
-In essence, the learning process of diffusion models is equivalent to the denoising process [19], which makes it possible to achieve excellent denoising performance in hyperspectral contexts.
-Different image modalities face different noise challenges.
-For example, hyperspectral image (HSI) recovery, which may be quite susceptible to distributional biases, is data-poor [138].
-HSI needs to take into account the inhomogeneous distribution of noise in the spectral dimension [139] as well as the correlation of noise between different frequency bands [140, 141].
-Fortunately, these noises can be efficiently removed by a diffusion model.Ceng and Cao et al.
-proposed Diff-Unmix [142] to address the challenge of recovering noise-degraded HSI through the fusion of spectral unmixing and conditional abundance generation, which employs a learnable block-based spectral unmixing strategy supplemented by a purely transformer-based backbone.
-We then introduce a self-supervised generative diffusion network to augment the abundance maps from the spectrally unmixed blocks.
-This network reconstructs noise-free decomposition probability distributions, effectively mitigating noise-induced degradation within these components.
-Finally, the reconstructed HSI is reconstructed by de-mixing reconstruction by mixing the diffusion-adjusted abundance maps with the spectral end-elements.
-he et al [140] proposed a truncated diffusion model that starts denoising from the intermediate step of the diffusion process instead of the pure noise in order to avoid destroying the inherently valid information in the HSI.
-In addition, Deng et al.
-proposed RGB image-enhanced hyperspectral Denoising Diffusion Probabilistic Models [143], which improved DM-based HSI denoising by combining RGB-DM pre-training on additional RGB images, and a fusion operator was introduced to integrate HSI-DM and RGB-DM, which improves the performance of DM-based HSI denoising.Kai et al.
-proposed a multi-scale spatial spectral for hyperspectral image frequency joint denoising network (MSFDN) [137], in order to describe and extract the noise more accurately and obtain better denoising effect, they constructed a subspace projection network to convert the hyperspectral image into a low-dimensional subspace, designed a noise separation network, a subspace inverse projection network and utilized the conditional diffusion model for better robustness.
+In conclusion, the application of diffusion models in hyperspectral image anomaly detection remains a rapidly evolving field with significant research potential.
+We anticipate a growing number of researchers will focus on this area, driving substantial advancements in anomaly detection.
+4.2.7 Noise Suppression
+A primary challenge in hyperspectral imaging is acquiring images without losing crucial information.
+During acquisition or subsequent processing, images can be corrupted by noise or artifacts, significantly degrading quality, especially with small objects or low contrast.
+Due to their generative nature, diffusion models are well-suited to addressing various noise reduction problems [135, 136, 137].
+For instance, [136] focuses on denoising in image restoration by designing a multi-shape spatial rectangular self-attention module and a spectral latent diffusion enhancement module.
+Leveraging the diffusion model allows for a tighter alignment with the desired HSI, enhancing denoising robustness.
+In image processing, hyperspectral images (HSIs) are invariably affected by various types of noise due to the inherent limitations of imaging techniques and environmental factors.
+Fundamentally, the learning process of diffusion models is akin to a denoising process [19], making them exceptionally effective for noise reduction in hyperspectral contexts.
+Different image modalities face unique noise challenges.
+For instance, HSI recovery can be particularly susceptible to distributional biases and often suffers from data scarcity [138].
+HSIs also contend with the inhomogeneous distribution of noise along the spectral dimension [139] and the correlation of noise across different frequency bands [140, 141].
+Fortunately, diffusion models can efficiently address these noise issues.
+Ceng and Cao et al.
+proposed Diff-Unmix [142] to tackle the challenge of recovering noise-degraded HSIs by fusing spectral unmixing with conditional abundance generation.
+This method employs a learnable block-based spectral unmixing strategy, supported by a purely transformer-based backbone.
+They then introduce a self-supervised generative diffusion network to augment the abundance maps from the spectrally unmixed blocks, reconstructing noise-free decomposition probability distributions and effectively mitigating noise-induced degradation within these components.
+Finally, the HSI is reconstructed through de-mixing by combining the diffusion-adjusted abundance maps with the spectral end-members.
+He et al.
+[140] introduced a truncated diffusion model that initiates denoising from an intermediate step of the diffusion process, rather than from pure noise.
+This approach aims to preserve the inherently valid information within the HSI.
+Additionally, Deng et al.
+proposed RGB image-enhanced hyperspectral Denoising Diffusion Probabilistic Models [143], improving DM-based HSI denoising by integrating RGB-DM pre-training on supplementary RGB images.
+They also introduced a fusion operator to combine HSI-DM and RGB-DM, further enhancing the performance of DM-based HSI denoising.
+Kai et al.
+proposed a multi-scale spatial spectral joint denoising network for hyperspectral images (MSFDN) [137].
+To more accurately describe and extract noise for superior denoising, they constructed a subspace projection network to convert HSIs into a low-dimensional subspace.
+They also designed a noise separation network and a subspace inverse projection network, leveraging a conditional diffusion model for enhanced robustness.
 Figure 13: Process diagram of noise suppression based on diffusion model
-IV-B8 Data Recovery
-Hyperspectral image (HSI) recovery aims to recover clean images from degraded observations [144], and the recovery of hyperspectral images plays a pivotal role in subsequent hyperspectral image applications.
-Hyperspectral data recovery faces challenges in effectively exploring the spatial nonlocal self-similarity and spectral low-rank properties inherent in HSI, while hyperspectral exists data scarcity.
-Existing model-based approaches have limitations in accurately modeling complex image features with manual a priori, and deep learning-based approaches have poor generalization capabilities, complex models, poor results, and parameter redundancy, Diffusion models, due to their impressive image recovery performance [145, 146, 147, 148, 149, 150, 151].
-In particular, it has recently attracted attention in terms of noise robustness [92, 95, 140, 152, 115, 119].
-Diffusion methods, however, are trained on large amounts of training data and perform very well within distributions, but may be quite susceptible to distributional bias.
-To address the above problem, Miao et al.
-proposed a self-supervised diffusion model for HSI recovery, the Denoising Diffusion Probabilistic Models for Spatial Spectroscopy (DDS2M) [138], which works by inferring the parameters of the proposed variational spatial spectral module (VS2M) during reverse diffusion, using only the degraded HSI without any additional training data.
-A loss function based on variational inference is customized to enable untrained spatial and spectral networks to learn a posteriori distributions that serve as transitions in the sampling chain to help reverse the diffusion process.
-Chen et al [56] use a dual window, iterative diffusion model, and reverse reconstruction to mitigate the impact of background estimation.
+4.2.8 Data Recovery
+Hyperspectral Image (HSI) recovery aims to reconstruct clean images from degraded observations [144], playing a pivotal role in subsequent HSI applications.
+This recovery process faces significant hurdles in effectively leveraging the intrinsic spatial nonlocal self-similarity and spectral low-rank properties of HSIs, further complicated by existing data scarcity.
+Traditional model-based approaches struggle to accurately model complex image features with manual prior assumptions, while deep learning methods often suffer from poor generalization, model complexity, suboptimal results, and parameter redundancy.
+However, diffusion models have recently garnered considerable attention for their impressive image recovery performance [145, 146, 147, 148, 149, 150, 151], particularly for their robustness to noise [92, 95, 140, 152, 115, 119].
+It’s worth noting, though, that diffusion methods are typically trained on large datasets and perform exceptionally well within those distributions, but they can be quite susceptible to distributional bias.
+To tackle these issues, Miao et al.
+introduced DDS2M (Denoising Diffusion Probabilistic Models for Spatial Spectroscopy) [138], a self-supervised diffusion model for HSI recovery.
+This model operates by inferring the parameters of its proposed variational spatial spectral module (VS2M) during reverse diffusion, using only the degraded HSI without needing any additional training data.
+A custom loss function based on variational inference enables untrained spatial and spectral networks to learn a posteriori distributions, which serve as transitions in the sampling chain to aid the reverse diffusion process.
+Chen et al.
+[56] employ a dual window, iterative diffusion model, and reverse reconstruction to mitigate the impact of background estimation.
 Li et al.
-propose an untrained diffusion model with a pre-trained diffusion model (HIR-DIFF) as a A supervised HSI recovery framework [141], which can recover clean HIS from the product of two low-rank components, is proposed based on singular value decomposition (SVD) and rank-revealed QR (RRQR) decomposition, and a novel exponential noise schedule is proposed to speed up the recovery process.
-Li et al.
-propose a spectral latent diffusion enhancement module [136], which introduces a latent diffusion-enhanced rectangular transformer for HSI recovery.
-A multi-shape spatial rectangular self-attentive module is proposed, which leads to a tighter alignment with the desired HSI.
-Meanwhile, there are also data recovery problems like noise mixing characterized by signal correlated and signal uncorrelated components, visual perception is affected and discriminative image distortion.
-Si et al.
-proposed a new model of CASSIDiff to reconstruct CASSI metrics [153], which employs a feature fusion mechanism based on the Discrete Wavelet Transform (DWT) and a spatial-spectral attention mechanism, and has a high degree of reliability in the recovery task.
-reliability.
+proposed HIR-DIFF (HSI Recovery via Diffusion Model with Untrained Prior) [141], a supervised HSI recovery framework.
+This framework, based on singular value decomposition (SVD) and rank-revealed QR (RRQR) decomposition, recovers clean HSIs from the product of two low-rank components.
+They also introduced a novel exponential noise schedule to accelerate the recovery process.
+Furthermore, Li et al.
+presented a spectral latent diffusion enhancement module [136], which incorporates a latent diffusion-enhanced rectangular transformer for HSI recovery.
+Their proposed multi-shape spatial rectangular self-attentive module leads to a tighter alignment with the desired HSI.
+Meanwhile, in the realm of data recovery, problems often arise from noise mixing, characterized by both signal-correlated and signal-uncorrelated components, which can affect visual perception and cause discriminative image distortion.
+To address this, Si et al.
+proposed CASSIDiff [153] for reconstructing CASSI metrics.
+This model employs a feature fusion mechanism based on the Discrete Wavelet Transform (DWT) and a spatial-spectral attention mechanism, demonstrating high reliability in recovery tasks.
 Yi et al.
-proposed an efficient diffusion model (EDiffSR) [149] for RSI SR, developed an efficient activation network (EANet), and developed a practical conditional prior enhancement module (CPEM), which reduces the computational budget and retains more information.
+introduced an efficient diffusion model (EDiffSR) [149] for RSI (Remote Sensing Image) Super-Resolution.
+They developed an efficient activation network (EANet) and a practical conditional prior enhancement module (CPEM), which collectively reduce the computational budget while retaining more information.
 Meng et al.
-proposed a conditional diffusion model-based SRSISR method (FastDiffSR) [148], which reduces the number of sampling steps, employs residual images, and combines channel attention and spatial attention to obtain better visualization with significantly improved generation quality.
+proposed FastDiffSR [148], a conditional diffusion model-based SRSISR (Single-Image Super-Resolution) method.
+FastDiffSR significantly improves generation quality and visual clarity by reducing the number of sampling steps, utilizing residual images, and combining channel and spatial attention mechanisms.
 Figure 14: Diffusion model-based research progress in hyperspectral image processing and analysis the tasks are established in nine subfields: (1) data enhancement and synthesis, (2) data generation and complementation, (3) classification, (4) segmentation, (5) anomaly detection, (6) target detection, (7) noise suppression, and (8) data recovery.
 No abbreviated name, with two initial letters at the beginning and end
-V Challenges and Limitations of the Diffusion Model
-V-A Computational complexity and Time cost of the model
+5 Challenges and Limitations of the Diffusion Model
+5.1 Computational complexity and Time cost of the model
 Diffusion Models achieve data generation and denoising through multi-step iterative Markov chains, whose computational complexity increases exponentially with the data dimensions.
 Taking the denoising diffusion probabilistic model (DDPM) [19] as an example, hundreds of forward and reverse diffusion computations are required to generate a single hyperspectral image, especially when processing hyperspectral data above 200 bands, and the GPU memory occupancy can be more than 20 GB (e.g., 1024 × 1024 resolution scenes).
-Although the improved DDIM [35] reduces the number of sampling steps to less than 50 through non-Markov chains, the spatial-spectral three-dimensional nature of hyperspectral data (see Section II.B) still leads to a single inference time as long as several minutes, which makes it difficult to satisfy real-time demanding tasks, such as agricultural monitoring and disaster warning.
-Although the Latent Diffusion Model (LDM) [18] reduces the computational volume by compressing the latent space, the information compression in the spectral dimension may lose key features (e.g., mineral absorption peaks), which affects the classification and anomaly detection accuracy (see Table II).
-V-B Impact of high dimensionality of hyperspectral data on model training
+Although the improved DDIM [35] reduces the number of sampling steps to less than 50 through non-Markov chains, the spatial-spectral three-dimensional nature of hyperspectral data (see Section 2.2) still leads to a single inference time as long as several minutes, which makes it difficult to satisfy real-time demanding tasks, such as agricultural monitoring and disaster warning.
+Although the Latent Diffusion Model (LDM) [18] reduces the computational volume by compressing the latent space, the information compression in the spectral dimension may lose key features (e.g., mineral absorption peaks), which affects the classification and anomaly detection accuracy (see Table 2).
+5.2 Impact of high dimensionality of hyperspectral data on model training
 The performance of diffusion models is highly dependent on large-scale labeled datasets, and the cost of labeling hyperspectral data far exceeds that of traditional RGB images.
@@
 In addition, conditional diffusion models based on classifier guidance (e.g., ADM-G [11]) need to pre-train high-precision classifiers, but the class imbalance problem of hyperspectral data (e.g., less than 1% of anomaly samples in anomaly detection) leads to difficulties in classifier training, which restricts the model generalization ability.
-V-C Adaptability issues to large-scale data
-Real hyperspectral images often contain non-Gaussian noise (e.g., sensor thermal noise, atmospheric scattering noise), and most of the existing diffusion models are based on additive Gaussian noise assumptions (see Section III.A.1) The forward process design of the DDPM [19] is only applicable to isotropic Gaussian noise, and the denoising effect is significantly degraded for streak or PoissonGaussian hybrid noise (PSNR reduction of 8 15 dB).
+5.3 Adaptability issues to large-scale data
+Real hyperspectral images often contain non-Gaussian noise (e.g., sensor thermal noise, atmospheric scattering noise), and most of the existing diffusion models are based on additive Gaussian noise assumptions (see Section 3.1.1) The forward process design of the DDPM [19] is only applicable to isotropic Gaussian noise, and the denoising effect is significantly degraded for streak or PoissonGaussian hybrid noise (PSNR reduction of 8 15 dB).
 Although Score-SDE [21] extends the scope of noise modeling through stochastic differential equations, its inverse process relies on the accurate estimation of the score function, which is susceptible to error accumulation in bands where the signal-to-noise ratio (SNR) fluctuates drastically (e.g., near-infrared bands).
 For example, the Diffusion Models’ denoising performance degrades by about 20% compared to the conventional wavelet approach under extreme conditions where the noise standard deviation exceeds 30 dB.
-In addition, the mixed image element problem (see Section II.C) leads to a high degree of spatial coupling between the noise and the target signal, further exacerbating the modeling difficulty.
-V-D Limitations of current models in practical applications
+In addition, the mixed image element problem (see Section 2.3) leads to a high degree of spatial coupling between the noise and the target signal, further exacerbating the modeling difficulty.
+5.4 Limitations of current models in practical applications
 The black-box nature of diffusion models exposes them to credibility challenges in hyperspectral critical applications.
 For example, in anomaly detection tasks, the model may misclassify rare features as noise by overfitting the training data (literature [16]).
-Experiments have shown that LDM [18] based anomaly detectors have a 18% higher false alarm rate than the traditional RX algorithm (see Section II.C) in the detection of unknown mineral types.
+Experiments have shown that LDM [18] based anomaly detectors have a 18% higher false alarm rate than the traditional RX algorithm (see Section 2.3) in the detection of unknown mineral types.
 In addition, although the Latent Diffusion Model improves the computational efficiency, the physical meaning of the latent vectors is unclear, which makes it difficult for users to trace the generation logic of the detection results.
@@
 Future research needs to combine hyperspectral data characteristics (e.g., spectral continuity, spatial-spectral correlation), develop lightweight architectures and adaptive noise modeling methods, and explore semi-supervised learning frameworks to reduce annotation costs, so as to promote its large-scale application in remote sensing, agriculture and other fields.
-VI Comparative and Experimental analysis
+6 Comparative and Experimental analysis
 In order to effectively illustrate the superiority of the diffusion model in processing hyperspectral images, this paper takes the experimental results of classification, target detection, anomaly detection, image recovery, and noise suppression as an example, and evaluates the performance of the diffusion model and other existing techniques through the visualization results and quantitative indexes.
-TABLE II: Comparison of Diffusion Model for Hyperspectral Images with Traditional Deep Learning Models
+Table 2: Comparison of diffusion and traditional deep learning models for HSI
 Deep Learning Model Advantages Disadvantages Future Development Trends
@@
 Combining with Transformer architecture to improve the efficiency of long sequence modeling.
-VI-A Experimental results and Performance assessment and Various metrics
-VI-A1 Introduction to evaluation metrics and Datasets
+6.1 Experimental results and Performance assessment and Various metrics
+6.1.1 Introduction to evaluation metrics and Datasets
 The receiver operating characteristic curve (ROC) is a qualitative analysis of the model.
@@
 SSIM (Structural Similarity) measures the similarity between two images, comparing images in terms of brightness, contrast and structure; FSIM (Feature Similarity) evaluates the image quality by calculating the phase coherence and gradient similarity in localized regions of the image, and is an improved version of SSIM; SAM (Sum of Absolute Errors) is another measure of the effectiveness of image denoising, which calculates the sum of the absolute value of the pixel differences between a denoised image and the original image counterparts.
-TABLE III: Summary of datasets used for the following experiments
+Table 3: Description of benchmark datasets used for evaluation
 Datasets Sensor Spectral Band Size Resolution
@@
 KAIST Thermal Camera LWIR 640$\times$480 0.5 m
-VI-A2 Experimental results
-In this section, several typical methods from the above methods will be selected and compared to study the performance of detection, where traditional methods:GAN, VAE, autoregressive models, other deep learning methods, are compared with diffusion models.
-Since the experimental dataset of each algorithm is different, a common dataset is chosen for comparison as much as possible, in order to verify the anomaly detection performance of different algorithms on the same dataset in a more reasonable way.
-Where, the bolded indicates the algorithm’s performance optimal value for this dataset.
-Comparison of classification experimental results, as shown in Figure 15, the classification effect of the first three methods is very poor, different classification color adulteration, the latter three methods overall classification effect is better, among them, the diffusion model-based method SpectralDiff classification effect is the best, followed by DMVL, their most boundary delineation is better versus the other methods, with less color adulteration, and SpectralDiff, despite the tiny classification there are some problems, but the overall effect is significantly better than other mainstream methods:
-Figure 15: Comparison of different datasets on classification methods for hyperspectral images: (a) Pseudo-color image (b) Real labeled image (c) SF [155] (d) miniGCN [156] (e) SSFTT [157] (f) DMVL [158] (g) SSGRN [159] (h) SpectralDiff [108]
-TABLE IV: Comparison of Metrics for Different Classification Methods on Indian Pines Dataset
+6.1.2 Experimental results
+In this section, a selection of typical methods from those previously discussed will be chosen for a comparative study of their detection performance.
+This comparison will encompass traditional methods such as GANs, VAEs, autoregressive models, other deep learning approaches, and diffusion models.
+To ensure a more reasonable assessment of anomaly detection performance across different algorithms, a common dataset will be used for comparison whenever possible.
+In the results, the bolded values will indicate the optimal performance achieved by an algorithm on that specific dataset.
+As illustrated in Figure 15, the classification performance of the first three methods is notably poor, exhibiting significant color adulteration across different classes.
+In contrast, the latter three methods demonstrate superior overall classification effects.
+Among these, the diffusion model-based method, SpectralDiff, achieves the best classification results, closely followed by DMVL.
+Both SpectralDiff and DMVL show improved boundary delineation compared to other methods, with minimal color adulteration.
+Although SpectralDiff has some minor classification issues in tiny areas, its overall performance is significantly better than other mainstream approaches.
+Figure 15: Classification results on multiple datasets using different methods: (a) Pseudo-color image (b) Real labeled image (c) SF [155] (d) miniGCN [156] (e) SSFTT [157] (f) DMVL [158] (g) SSGRN [159] (h) SpectralDiff [108]
+Table 4: Comparison of Metrics for Different Classification Methods on Indian Pines Dataset
 Methods Evaluation metrics (Indian Pines)
@@
 SpectralDiff 93.146 96.438 92.175 48.53
-Table IV shows that the Spectral-Diff method is in the first place in Indian Pines dataset classification in terms of overall accuracy, average accuracy, and kappa coefficient, which is significantly improved over other methods and is the current optimal method.OA and AA are higher than the second place (SSGRN) by 0.8% and 0.34%, which indicates that the method is more robust in classification and more effective in capture features and reduce noise interference.
-However, it can also be seen that Spectral-Diff consumes much more time than other methods, which is its shortcoming.
-TABLE V: Comparison of Metrics for Different Classification Methods on Pavia University Dataset
+Table 4 reveals that the Spectral-Diff method achieves the top position in the Indian Pines dataset classification across overall accuracy (OA), average accuracy (AA), and kappa coefficient.
+Its performance shows a significant improvement over other methods, marking it as the current optimal approach.
+Notably, its OA and AA are 0.8% and 0.34% higher than the second-place method (SSGRN), indicating superior robustness in classification, more effective feature capture, and reduced noise interference.
+However, a drawback of Spectral-Diff is its considerably higher time consumption compared to other methods.
+Table 5: Comparison of Metrics for Different Classification Methods on Pavia University Dataset
 Methods Evaluation metrics (Pavia University)
@@
 SpectralDiff 94.775 93.843 93.061 256.23
-As can be seen in Table V, the method has two metrics on the Pavia University dataset that are optimal 0.98% and 1.23% higher than the second (SSGRN), and the other is only 0.63% lower than the optimal, which is a relatively small gap, and it is still the optimal method among these methods, but the shortcoming of the use of samples is the long time consumed, which is more than twice as long as the other methods.
-TABLE VI: Comparison of Metrics for Different Classification Methods on Salinas Dataset
+As Table 5 shows, this method achieves optimal results on the Pavia University dataset for two metrics, surpassing the second-place method (SSGRN) by 0.98% and 1.23%.
+While another metric is only 0.63% lower than the optimum, which is a relatively small difference, it remains the best among these methods.
+However, a significant drawback is its time consumption, which is more than double that of other approaches.
+Table 6: Comparison of Metrics for Different Classification Methods on Salinas Dataset
 Methods Evaluation metrics (Salinas)
@@
 SpectralDiff 98.971 99.465 98.854 242.16
-Table VI shows the classification on Salinas dataset and the same conclusion as the previous two tables, combined with Tables IV, V, and VI, the Spectral-Diff method is optimal in terms of performance compared to the other methods, but it is more time consuming and slower.
-SSGRN is optimal if balancing accuracy and efficiency, and miniGCN is optimal if low latency and high tolerance accuracy.
-Comparison of target detection experimental results, as shown in Figure 16, the overall effect of these mainstream methods used in target detection is good, some methods exist undetected certain points or misdetection, among them, the diffusion model-based method Diff-Mosaic detection effect is the best, for the detection of the point of the edge of the border details of the best description, and the real labeling map is basically consistent, basically there is no missing parts:
-Figure 16: Comparison of different datasets on target detection methods for hyperspectral images: (a) Original image (b) Real labeled image (c) ALC-Net [160] (d) ACM [161] (e) UIU-Net [162] (f) DNA-Net [163] (g) Diff-Mosaic [118]
-TABLE VII: Comparison of Metrics for Different Target Detection Methods on SIRST Dataset
+Table 6 presents the classification results on the Salinas dataset, leading to the same conclusion as Tables 4 and 5.
+Across all three datasets (Indian Pines, Pavia University, and Salinas), the Spectral-Diff method consistently demonstrates optimal performance when compared to other approaches.
+However, this superior accuracy comes at the cost of higher computational time and slower processing speeds.
+If the goal is to balance accuracy with efficiency, SSGRN emerges as the optimal choice.
+Conversely, for applications demanding low latency and high tolerance for accuracy, miniGCN is the most suitable option.
+As seen in Figure 16, the overall performance of these mainstream target detection methods is good.
+However, some methods exhibit instances of undetected points or false positives.
+Among them, the diffusion model-based method, Diff-Mosaic, demonstrates the best detection effect, providing the most accurate and detailed boundary descriptions for detected points.
+It is largely consistent with the true labeling map, with virtually no missing components.
+Figure 16: Target detection results on various datasets with competing methods: (a) Original image (b) Real labeled image (c) ALC-Net [160] (d) ACM [161] (e) UIU-Net [162] (f) DNA-Net [163] (g) Diff-Mosaic [118]
+Table 7: Comparison of Metrics for Different Target Detection Methods on SIRST Dataset
 Methods Evaluation metrics (SIRST)
@@
 Diff-Mosaic 79.44 99.99 3.19
-Table VII shows that the Diff-Mosaic method has the optimal intersection ratio, detection probability, and false alarm rate for target detection on the SIRST dataset, which is 3.11% higher than the second place (DNA-Net), 1.89% higher than the second place (UIU-Net), and 87.8% lower than the second place (UIU-Net), respectively.
-Among them, the detection tends to 100% detection is very good with less error and is the best among these methods.
-TABLE VIII: Comparison of Metrics for Different Target Detection Methods on NUDT-SIRST Dataset
+Table 7 demonstrates that the Diff-Mosaic method achieves optimal performance for target detection on the SIRST dataset.
+It surpasses the second-place methods (DNA-Net and UIU-Net) with a 3.11% higher intersection ratio and a 1.89% higher detection probability.
+Furthermore, its false alarm rate is 87.8% lower than UIU-Net.
+This indicates that Diff-Mosaic’s detection rate approaches 100% with minimal errors, making it the superior method among those compared.
+Table 8: Comparison of Metrics for Different Target Detection Methods on NUDT-SIRST Dataset
 Methods Evaluation metrics (NUDT-SIRST)
@@
 Diff-Mosaic 91.18 99.47 1.91
-Table VIII shows the Diff-Mosaic in manual synthesis dataset NUDT-SIRST utilization, which is consistent with the metrics effect shown in Table 7, and moreover, it shows the superiority of this method among these methods with the best performance.
-TABLE IX: Comparison of Parameters and Reasoning for Different Target Detection Methods
+Table 8 further corroborates the findings from Table 7, showing that Diff-Mosaic consistently delivers optimal performance on the manually synthesized NUDT-SIRST dataset.
+This reinforces the method’s superior effectiveness among the compared approaches.
+Table 9: Comparison of Parameters and Reasoning for Different Target Detection Methods
 Methods Calculate the number of parameters Time (s)
@@
 Diff-Mosaic 4.7 43.42
-Table IX shows that Diff-Mosai is less in the number of parameters, but not the best, while the inference time is the slowest among them, showing the shortcomings of the method in applications, 57.3% higher than the fastest time (ACM).
-Meanwhile, combining Tables VII, VIII and IX, again Diff-Mosaic is the best if pursuing the best performance, DNA-Net is the best if balancing performance and efficiency, and ACM is the best if high tolerance and low latency.
-Comparison of the results of the anomaly detection experiments, as shown in Figure 17, compared with the traditional mainstream methods, the diffusion model-based methods BSDM, DWSDiff in the anomaly detection of the best detection effect, high detection rate, low false detection rate, and the real labeling map is basically consistent with the other mainstream detection methods to detect the detection of the detection effect is poor, the detection rate is low (the existence of the detection of the anomalies do not exist), the false detection rate is high:
-Figure 17: Comparison of different datasets on hyperspectral image anomaly detection methods: (a) Pseudo-color map (b) Real label map (c) GAED [164] (d) RGAE [165] (e) Auto-AD [166] (f) DFAN-HAD [167] (g) DirectNet [168] (h) BSDM [133] (i) TGFA [169] (j)DWSDiff [56]
-TABLE X: Comparison of metrics for each anomaly detection methods
+Table 9 indicates that while Diff-Mosaic has a comparatively lower number of parameters, it isn’t the most parameter-efficient method.
+More significantly, its inference time is the slowest among those compared, being 57.3% higher than the fastest method (ACM).
+This highlights a key limitation of Diff-Mosaic in practical applications where speed is critical.
+When considering the collective results from Tables 7, 8, and 9, the optimal choice for target detection depends on specific priorities.
+If the goal is to achieve the absolute best performance, Diff-Mosaic stands out.
+However, if balancing strong performance with efficiency is paramount, DNA-Net proves to be the most suitable option.
+Conversely, for applications demanding high tolerance for minor accuracy variations and low latency, ACM is the ideal choice.
+As presented in Figure 17, when comparing the results of anomaly detection experiments, the diffusion model-based methods, BSDM and DWSDiff, demonstrate superior detection performance compared to traditional mainstream approaches.
+They achieve a high detection rate and a low false detection rate, largely aligning with the true labeling map.
+In contrast, other mainstream detection methods exhibit poorer detection effects, characterized by lower detection rates (missing existing anomalies) and higher false detection rates.
+Figure 17: Comparative anomaly detection performance across hyperspectral datasets: (a) Pseudo-color map (b) Real label map (c) GAED [164] (d) RGAE [165] (e) Auto-AD [166] (f) DFAN-HAD [167] (g) DirectNet [168] (h) BSDM [133] (i) TGFA [169] (j)DWSDiff [56]
+Table 10: Comparison of metrics for each anomaly detection methods
 Methods Evaluation metrics
@@
 DWSDiff 0.97427 0.55329 0.01635
-TABLE XI: Comparison of metrics for each anomaly detection methods
+Table 11: Comparison of metrics for each anomaly detection methods
 Methods Evaluation metrics
@@
 DWSDiff 1.52756 0.9579 1.51121
-TABLE XII: Comparison of metrics for each anomaly detection methods
+Table 12: Comparison of metrics for each anomaly detection methods
 Methods Evaluation metrics
@@
 DWSDiff 0.53694 67.20618
-As can be seen in Table X, this table shows the accuracy of the anomaly detection methods, in which BSDM and DWSDiff are based on diffusion model, DWSDiff is in the first place in all other items except for the integrated detection performance (${mAUC}_{TD}$) which is second only to DFAN, while the data of BSDM is not outstanding (third in the integrated detection performance), but it shows the effect is not is not bad.
-TABLE XIII: Comparison of Reasoning Time for Each Anomaly Detection Method
+As Table 10, 11, 12 illustrates, which presents the accuracy of anomaly detection methods, BSDM and DWSDiff are diffusion model-based approaches.
+Among these, DWSDiff ranks first in all metrics except for integrated detection performance (${mAUC}_{TD}$), where it is second only to DFAN.
+While BSDM’s data isn’t as outstanding (ranking third in integrated detection performance), its effectiveness is still commendable.
+Table 13: Comparison of Reasoning Time for Each Anomaly Detection Method
 Methods Reasoning time (s)
@@
 DWSDUR 0.8016 0.8169 0.8192 0.8094
-Table XI shows that both diffusion models have longer and slower inference times for each medium dataset scenario, with average times 99.8% higher (BSDM) and 99.7% higher (DWSDiff) than the shortest time (Auto-Net), which is hundreds of times higher than the optimal time, and is very slow in comparison.
-Again seeking performance DWSDiff is optimal and BSDM is a good choice.
-Comparison of the experimental results of image recovery, as shown in Fig.
-18, the overall effect of image recovery are relatively modest, but the methods based on the diffusion model (DDPM, EDiffSR, FastDiffSR) are better handled in terms of details, and the degree of recovery is closer to the high-resolution map:
-Figure 18: Comparison of different datasets on hyperspectral image recovery methods: (a) High resolution map (b) HESNet [170] (c) TransENet [171] (d) NDSRGAN [172] (e) HAT [173] (f) DDPM [19] (g) EDiffSR [149] (h) FastDiffSR [148]
-TABLE XIV: Comparison of Metrics for Different Image Recovery Methods on Potsdam Dataset
+Table 13 indicates that both diffusion models, BSDM and DWSDiff, exhibit significantly longer and slower inference times across various medium dataset scenarios.
+Their average times are 99.8% (BSDM) and 99.7% (DWSDiff) higher than the shortest time achieved by Auto-Net, translating to hundreds of times slower than the optimal.
+This highlights a considerable challenge in terms of computational efficiency.
+Nonetheless, if the primary goal is to maximize anomaly detection performance, DWSDiff remains the optimal choice, with BSDM also being a strong contender.
+As depicted in Figure 18, the overall image recovery effects are relatively modest across all methods.
+However, diffusion model-based approaches (DDPM, EDiffSR, FastDiffSR) demonstrate superior handling of details, resulting in a recovery quality that is notably closer to the high-resolution ground truth.
+Figure 18: Visual comparison of hyperspectral image recovery methods: (a) High resolution map (b) HESNet [170] (c) TransENet [171] (d) NDSRGAN [172] (e) HAT [173] (f) DDPM [19] (g) EDiffSR [149] (h) FastDiffSR [148]
+Table 14: Comparison of Metrics for Different Image Recovery Methods on Potsdam Dataset
 Methods Evaluation metrics(Potsdam)
@@
 FastDuFSR 33.130 22.832 0.1152 23.229
-As can be seen in Table XII, the three methods based on diffusion model (DDPM, EDiffSR, FastDiffSR) show unique advantages in key indicators, among which, the FsatDiffSR method has the best overall performance in terms of PSNR and FID, which is far superior to the traditional methods (compared to the second: (HAT) 0.43%, (DDPM) 6.07%, and (DDPM) 13.5% difference), indicating that its recovered image reaches a higher level in both detail fidelity and distributional consistency, but in terms of the number of parameters, the diffusion model-based methods are all higher than HESNet: 79.4%, 79.7%, 76.6%, respectively, thus affecting the training time too long.
-TABLE XV: Comparison of Metrics for Different Image Recovery Methods on Toronto Dataset
+As evidenced in Table 14, the three methods based on diffusion models—DDPM, EDiffSR, and FastDiffSR—demonstrate unique advantages in key image recovery indicators.
+Notably, FastDiffSR exhibits the best overall performance in terms of PSNR and FID, significantly outperforming traditional methods.
+Compared to the second-best methods (HAT for PSNR and DDPM for FID), FastDiffSR shows substantial improvements: 0.43% for PSNR over HAT, and 6.07% and 13.5% for FID over DDPM.
+This indicates that FastDiffSR recovers images with superior detail fidelity and distributional consistency.
+However, a drawback for all diffusion model-based methods is their higher number of parameters compared to HESNet, specifically 79.4% higher for DDPM, 79.7% for EDiffSR, and 76.6% for FastDiffSR.
+This increased parameter count contributes to longer training times.
+Table 15: Comparison of Metrics for Different Image Recovery Methods on Toronto Dataset
 Methods Evaluation metrics (Toronto)
@@
 EasDdfSR 25.527 55.761 0.1561 23.229
-Similar to Table XII, it can be seen from Table XIII that the three methods based on the diffusion model have a good overall performance, especially FastDiffSR has the best overall performance, and combining the two tables, the three diffusion methods have a higher quality of generation and more semantic information retention than the traditional methods.
-However, the three methods still have the problem of larger number of parameters in comparison.
-TABLE XVI: Comparison of Measurement metrics for Different Image Recovery Methods
+Similar to Table 14, Table 15 also demonstrates that the three methods based on the diffusion model exhibit strong overall performance, with FastDiffSR again standing out as the best.
+When combining insights from both tables, it’s clear that these three diffusion methods produce higher quality generations and retain more semantic information compared to traditional approaches.
+However, a consistent issue across all three diffusion methods remains their relatively larger number of parameters.
+Table 16: Comparison of Measurement metrics for Different Image Recovery Methods
 Methods Measurement metrics
@@
 EastDiffSR 7.1254 16.342
-The entropy and generation quality metrics in Table XIV show that the three diffusion methods retain more semantic information and have higher generation and recovery quality, corresponding to Tables XII and XIII.Also combining the three tables, overall FastDiffSR has a balanced performance in terms of recovery accuracy and efficiency, DDPM is more advantageous in terms of perceptual quality, and EDiffSR can be used as a practical choice for moderate performance.
-Comparison of noise suppression experimental results, as shown in Figure 19, with the noise map as well as the real noiseless map compared to the overall performance of the Diff-Unmix method has the best effect, followed by DDRM, and finally DDS2M are based on the diffusion model method, they are more successful noise suppression; in contrast to the traditional mainstream method effect is worse, to LLxRGTV in which the best:
-Figure 19: Comparison of different datasets on hyperspectral noise suppression methods: (a) Noise map (b) True noise-free map (c) LLxRGTV [174] (d) FGSLR [175] (e) LRTDCTV [176] (f) DIP [177] (g) LLnRPnP [178] (h) HLRTF [110] (i) DDRM [179] (j)DDS2M [138] (k)Diff-Unmix [142]
-TABLE XVII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
-Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,3))
+Table 16, which presents entropy and generation quality metrics, reinforces the findings from Tables 14 and 15.
+It demonstrates that the three diffusion methods—FastDiffSR, DDPM, and EDiffSR—retain more semantic information and achieve higher generation and recovery quality.
+Overall, FastDiffSR offers a balanced performance in terms of both recovery accuracy and efficiency.
+DDPM holds an advantage in perceptual quality, while EDiffSR can be considered a practical choice for moderate performance needs.
+As illustrated in Figure 19, when comparing the noise map and the true noiseless map, the Diff-Unmix method demonstrates the best overall performance in noise suppression, followed by DDRM and then DDS2M.
+All of these are diffusion model-based approaches, and they achieve notably successful noise suppression.
+In contrast, traditional mainstream methods exhibit poorer effects, with LLxRGTV being the best among them.
+Figure 19: Performance comparison of denoising methods on hyperspectral images: (a) Noise map (b) True noise-free map (c) LLxRGTV [174] (d) FGSLR [175] (e) LRTDCTV [176] (f) DIP [177] (g) LLnRPnP [178] (h) HLRTF [110] (i) DDRM [179] (j)DDS2M [138] (k)Diff-Unmix [142]
+Table 17: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
+Methods Evaluation metrics(KAIST Noise: N(0,0,3))
 PSNR SSIM FSIM SAM Time (s)
@@
 Diff-Unmix 31.408 0.902 0.958 0.282 42
-TABLE XVIII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
-Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,2))
+Table 18: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
+Methods Evaluation metrics(KAIST Noise: N(0,0,2))
 PSNR SSIM FSIM SAM Time (s)
@@
 Diff-Unmix 33.059 0.964 0.940 0.116 37
-The last three methods (DDRM/DDS2M/Diff-Unmix) in the experiment are based on the diffusion model, as can be seen from Tables XV and XVI, the first choice of the KAIST two different noise situations in the dataset, these three methods in the noise suppression performance have their own characteristics, and the Diff-Unmix in the key indexes is the most outstanding performance, which PSRN, SSIM, and FSIM are the highest values, compared with other methods to show its significant advantages in noise suppression, structure preservation and spectral fidelity, but the time processing of the method under the diffusion model is slightly higher (HLRTF is the fastest), and the overall efficiency is still in the acceptable range.
-TABLE XIX: Comparison of Metrics for Different Noise Suppression Methods on CAVE Dataset
+In the experiments detailed in Tables 17 and 18, the last three methods—DDRM, DDS2M, and Diff-Unmix—are all based on the diffusion model.
+When evaluated on the KAIST dataset under two different noise conditions, these three methods demonstrate distinct strengths in noise suppression.
+Diff-Unmix consistently shows the most outstanding performance across key indices, achieving the highest PSNR, SSIM, and FSIM values.
+This highlights its significant advantages in noise suppression, structure preservation, and spectral fidelity compared to other methods.
+While the processing time for diffusion model-based methods like Diff-Unmix is slightly higher (with HLRTF being the fastest), their overall efficiency remains within an acceptable range.
+Table 19: Comparison of Metrics for Different Noise Suppression Methods on CAVE Dataset
 Methods Evaluation metrics (CAVE)
@@
 Diff-Unmix 32.714 0.940 0.957 0.129 43
-Table XVII shows that the methods under the diffusion model are overall better than the other methods, of which Diff-Unmix remains the most prominent, demonstrating that the diffusion model method is overall superior to the other methods, with the best performance and the noise suppression presenting the optimal effect.
-TABLE XX: Comparison of Metrics for Different Noise Suppression Methods on CAVE-Toy Dataset
+Table 19 indicates that the diffusion model-based methods generally outperform other approaches, with Diff-Unmix remaining the most prominent.
+This consistently demonstrates the overall superiority of diffusion models in noise suppression, delivering optimal performance and the best noise suppression effects.
+Table 20: Comparison of Metrics for Different Noise Suppression Methods on CAVE-Toy Dataset
 Methods Evaluation metrics (CAVE-Toy)
@@
 Diff-Unmix 28.046 0.945 0.993 0.156 43
-Table XVIII is roughly in agreement with Tables XV, XVI and XVII diffusion models method noise suppression performance is generally the best, and in combination with the above tables, the overall effect is highly acceptable, with obvious denoising effect, close to noiseless, and with better detail retention, despite the fact that the temporal processing is not particularly fast.
-VI-B Key discoveries and Summaries from the Experiments
-Through the experimental comparisons in Section V.B.2, the diffusion model shows significant advantages in hyperspectral image processing tasks, but also exposes some limitations.
-In the classification task, the OA (93.15%, 94.78%, 98.97%), AA (96.44%, 93.84%, 99.47%) and Kappa coefficients of the SpectralDiff method based on the diffusion model on the Indian Pines, Pavia University and Salinas datasets ( 92.18%, 93.06%, 98.85%) are better than traditional deep learning methods (e.g., GAN, VAE), indicating that it can effectively capture the complex spectral-spatial features of hyperspectral data.
-However, its inference time (48.53 s, 256.23 s, 242.16 s) is significantly higher than that of lightweight models such as miniGCN (0.95 s, 2.40 s, 2.13 s), revealing the computational efficiency bottleneck of the diffusion model in high-dimensional data processing.
-In the target detection task, the diffusion-based Diff-Mosaic method achieves optimal IoU (79.44%, 91.18%), detection probability (Pd: 99.99%, 99.47%), and false alarm rate (Fa: 3.19%, 1.91%) on the SIRST and NUDT-SIRST datasets, and is especially good at complex background suppression and target localization accuracy.
-However, the number of parameters (4.7M) and the inference time (43.42 s) are still higher than that of the traditional method ACM (0.29M, 18.53 s), which indicates that the diffusion model needs to be further lightweighted to meet the real-time demand.
-In the anomaly detection task, DWSDiff comprehensively leads in mAUC metrics, such as mAUC(D,F) (0.97427), mAUC(BS) (0.95791), and mAUC(SNPR) (67.20618), which verifies its ability to suppress complex backgrounds through the dual-window strategy and iterative diffusion.
-However, its inference time (0.8094 s) is still much higher than that of Auto-AD (0.0022 s), and the model is sensitive to hyperparameters (e.g., noise scheduling), and its generalization ability needs to be further optimized, suggesting that the practical deployment of diffusion models in anomaly detection needs to balance accuracy and efficiency.
-In image recovery tasks, diffusion model methods (e.g., FastDiffSR, EDiffSR, and DDPM) significantly outperform conventional methods in key metrics (PSNR, FID, LPIPS).
-For example, on the Potsdam dataset, FastDiffSR achieves a PSNR of 33.13 dB and a FID of 22.83, indicating that it outperforms conventional methods in both detail fidelity and distribution consistency.
-In addition, FastDiffSR achieves a PSNR of 25.53 dB and a FID of 55.76 on the Toronto dataset, which further validates its advantage in semantic information retention.
-However, the high number of parameters of the diffusion model method (e.g., 23.23M for FastDiffSR) leads to a significant increase in training time.
-It shows that the diffusion model has strong generative capability, but its computational efficiency and model complexity still need to be optimized to meet the actual deployment requirements.
-In noise suppression tasks, diffusion model-based methods (e.g., Diff-Unmix, DDRM, and DDS2M) show significant advantages in metrics such as PSNR, SSIM, and FSIM.
-For example, on the KAIST dataset, Diff-Unmix achieves a PSNR of 31.41 dB, an SSIM of 0.902, and an FSIM of 0.958, indicating significant advantages in noise suppression, structure preservation, and spectral fidelity.
-In addition, on the CAVE dataset, Diff-Unmix has a PSNR of 32.71 dB and an SSIM of 0.940, further validating its superiority.
-However, the long inference time of the diffusion model method (e.g., 42 s for Diff-Unmix) restricts its application in scenarios with high real-time requirements, suggesting that the diffusion model is able to remove the noise and retain the details efficiently, but its computational efficiency still needs to be further improved.
+Table 20 largely corroborates the findings from Tables 17, 18, and 19, confirming that diffusion model-based methods generally offer the best noise suppression performance.
+When considering all these tables together, the overall effect of diffusion models is highly acceptable, demonstrating a clear denoising effect that closely approaches noiseless images while retaining superior detail.
+This is true despite their processing time not being exceptionally fast.
+6.2 Key discoveries and Summaries from the Experiments
+As evidenced by the experimental comparisons in Section 5.2.2, diffusion models exhibit significant advantages in hyperspectral image processing tasks, while also revealing certain limitations.
+In classification tasks, the SpectralDiff method, based on diffusion models, demonstrates superior accuracy metrics (OA, AA, and Kappa coefficients) across the Indian Pines, Pavia University, and Salinas datasets.
+For instance, its OA values were 93.15%, 94.78%, and 98.97% respectively, outperforming traditional deep learning methods like GANs and VAEs.
+This indicates its effectiveness in capturing complex spectral-spatial features.
+However, a significant drawback is its inference time, which is considerably higher (e.g., 48.53 s on Indian Pines) compared to lightweight models such as miniGCN (0.95 s), exposing a computational efficiency bottleneck for high-dimensional data processing.
+For target detection, the diffusion-based Diff-Mosaic method achieves optimal performance on the SIRST and NUDT-SIRST datasets, with excellent IoU (79.44% on SIRST), high detection probability (99.99% on SIRST), and a low false alarm rate (3.19% on SIRST).
+It particularly excels at complex background suppression and precise target localization.
+Despite this, its parameter count (4.7M) and inference time (43.42 s) remain higher than traditional methods like ACM (0.29M parameters, 18.53 s inference time).
+This suggests that further lightweighting of diffusion models is needed to meet real-time demands in target detection.
+In anomaly detection, DWSDiff comprehensively leads in mAUC metrics (e.g., ${mAUC}_{TD}$ of 0.97427), validating its ability to suppress complex backgrounds through its dual-window strategy and iterative diffusion.
+While BSDM’s performance is also respectable, DWSDiff stands out.
+Nevertheless, DWSDiff’s inference time (0.8094 s) is still substantially higher than highly efficient methods like Auto-AD (0.0022 s).
+Additionally, the model’s sensitivity to hyperparameters (e.g., noise scheduling) and its need for further optimization in generalization ability indicate that practical deployment of diffusion models in anomaly detection requires a careful balance between accuracy and efficiency.
+For image recovery, diffusion model methods such as FastDiffSR, EDiffSR, and DDPM significantly outperform conventional methods in key metrics like PSNR and FID.
+For example, FastDiffSR achieved a PSNR of 33.13 dB and an FID of 22.83 on the Potsdam dataset, demonstrating superior detail fidelity and distributional consistency.
+Similarly, on the Toronto dataset, FastDiffSR maintained its advantage with a PSNR of 25.53 dB and an FID of 55.76, confirming its strength in semantic information retention.
+However, the high number of parameters in these diffusion models (e.g., FastDiffSR with 23.23M parameters) leads to a substantial increase in training time.
+This highlights that while diffusion models possess strong generative capabilities, their computational efficiency and model complexity require further optimization for practical deployment.
+In noise suppression tasks, diffusion model-based methods, including Diff-Unmix, DDRM, and DDS2M, show significant advantages across metrics like PSNR, SSIM, and FSIM.
+On the KAIST dataset, Diff-Unmix achieved a PSNR of 31.41 dB, an SSIM of 0.902, and an FSIM of 0.958, underscoring its benefits in noise suppression, structure preservation, and spectral fidelity.
+Its superiority was further validated on the CAVE dataset with a PSNR of 32.71 dB and an SSIM of 0.940.
+Despite their effectiveness in removing noise and retaining details, the long inference time of diffusion models (e.g., 42 s for Diff-Unmix) restricts their application in scenarios demanding high real-time performance, suggesting a need for improved computational efficiency.
 Some key points are also found.1.
@@
 Future research can explore hybrid architectures to balance efficiency and performance.
-VII Outlook
+7 Outlook
 In recent years, diffusion models have made significant progress in tasks such as image generation, denoising and reconstruction by virtue of their powerful generative capabilities and modeling advantages over complex data distributions.
@@
 In the future, through the combination of algorithmic innovation, hardware synergy and automation technology, the diffusion model is expected to realize smarter and more efficient analysis systems in the fields of environmental monitoring, precision agriculture, etc., and to promote hyperspectral technology from the laboratory to large-scale applications.
-VIII Conclusion
-In summary, compared with other deep generative models, the diffusion model is robust to the inherent noise in hyperspectral images, better adapts to its variability and complexity, and provides a more stable training process.
-Meanwhile, diffusion model with progressive denoising mechanism with high-dimensional data modeling capability, the diffusion model has injected a revolutionary impetus into the field of hyperspectral image analysis.
-It effectively solves the limitations of traditional methods in high-dimensional data processing, noise interference and small sample scenarios through the Markov chain noise injection and inverse generation process.
-Therefore, it is an inevitable trend for diffusion model to be applied in the field of hyperspectral images.
-This paper firstly introduces the related knowledge of hyperspectral as well as the basic theory of diffusion model, and then, reviews the research of diffusion model in the field of hyperspectral images, including: data enhancement, generation and complementation, classification segmentation, anomaly detection, target detection, denoising and other tasks, moreover, this paper fully demonstrates the superiority of diffusion model in this field by comparing some visualization results and quantitative indexes on hyperspectral image processing between diffusion model and traditional model.
-The limitations of the existing diffusion models in the hyperspectral domain are also discussed, and it is emphasized that further exploratory studies can be carried out for extended applications and model deployment.
+8 Conclusion
+In summary, diffusion models offer significant advantages over other deep generative models for hyperspectral image analysis.
+They demonstrate robustness to the inherent noise in hyperspectral images, adapt better to their variability and complexity, and provide a more stable training process.
+With their progressive denoising mechanism and high-dimensional data modeling capabilities, diffusion models have injected a revolutionary impetus into the field of hyperspectral image analysis.
+They effectively overcome the limitations of traditional methods in handling high-dimensional data, noise interference, and small sample scenarios through their Markov chain noise injection and inverse generation process.
+Therefore, the application of diffusion models in hyperspectral imaging is an inevitable trend.
+This paper first introduced relevant knowledge about hyperspectral images and the basic theory of diffusion models.
+Subsequently, it reviewed research on diffusion models within the hyperspectral domain, covering tasks such as data enhancement, generation and complementation, classification segmentation, anomaly detection, target detection, and denoising.
+Moreover, by comparing various visualization results and quantitative indices between diffusion models and traditional models in hyperspectral image processing, this paper fully demonstrated the superiority of diffusion models in this field.
+Finally, the limitations of existing diffusion models in the hyperspectral domain were discussed, emphasizing the need for further exploratory studies concerning extended applications and model deployment.